A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.02

Title word cross-reference

2 × 2 [55, 227]. ^{(}R) [649]. H [619]. κ [77]. R × C [776].

-Likelihood [619].

13 [689]. 1916 [387].

2005 [757]. 2007 [750]. 20th [321]. 2B [312]. 2nd [374, 186, 301, 493, 565, 618, 304, 519, 312, 555, 617, 490, 587, 197].

3rd [516, 763, 302, 589].

5th [377].

80th [464].

A.Baddeley [554]. Abdominal [389]. Abney [832]. Abuse [96]. Academic [140]. Academy [700]. Accidental [154]. Accidents [12]. Accounting [802]. Accurate [287]. Achievement [603, 631, 822, 140]. Acquired [773]. Acquisitions [632]. across [123, 236]. Active [609]. Actively [473]. Activity [544, 629]. Actuarial [812]. Acute [353, 419, 775, 808, 293, 354]. Adapting [609]. Adaptive [7, 644, 688]. Additive [567]. Adequacy [223]. Adjust [478]. Adjusted [235]. Adjustment [575]. Adjustments [180]. Administrative [419]. Admissions [141]. Adult [825]. Adulthood [505]. Adults [95]. Advanced [178, 165, 312, 758]. Advances [821, 367]. aetiological [45]. Affecting [576]. Age [417, 327, 746]. Agencies [809]. Ages [139]. Aggregate [673, 675, 69, 51]. Agreement [417, 313]. Agresti [641]. Agriculture [587]. Air [27, 445]. Aitken [657]. Aitkin [490]. al [232, 237, 57, 44, 49, 52]. Alcohol [31]. Alfassi [460]. Algebra [459]. Algorithm [209, 637]. Algorithms [493]. Allenby [528]. Alleviating [674]. Allison [642]. Allowing [23]. Alternative [630, 538, 416, 206]. Ambulance [781]. American [91]. among [17, 503, 70, 125]. Analyses [736, 480, 165, 10, 476, 98, 686]. Analysing [259, 805, 759, 437, 469, 342]. Analysis [864, 610, 82, 523, 222, 221, 17, 54, 96, 65, 520, 310, 1, 336, 624, 56, 64, 466, 517, 230, 596, 546, 81, 168, 454, 319, 467, 780, 665, 586, 790, 783, 292, 848, 200, 725, 190, 687, 591, 646, 179, 477, 539, 642, 327, 348, 793, 547, 692, 649, 726, 849, 619, 320, 375, 20, 715, 524, 71, 325, 104, 166, 590, 69, 779, 361, 672, 720, 827, 792, 862, 301, 556, 463, 865, 491, 621, 840, 833, 439, 641, 730, 856, 529, 592, 367, 346]. Analysts [643]. Analytical [460]. Analyzing [374]. Ancient [179]. Anders [751]. Andrew [383]. Aneurysm [389]. Angle [286]. Ankerst [519]. Anonymized [76, 29]. Antenatal [255]. Anthropometry [99]. Aortic [389]. Application [82, 782, 503, 165, 482, 450, 258, 605, 773, 210, 826, 746, 805, 361, 45]. Applications [72, 678, 821, 824, 417, 454, 447, 720, 493, 623, 527, 618, 435, 762, 760, 594, 425, 754, 756, 332, 335]. Applied [775, 836, 221, 372, 556, 865, 555, 710, 194]. Appliquée [645]. Applying [700, 549, 218]. Approach [869, 65, 453, 701, 21, 165, 497, 635, 160, 785, 649, 611, 777, 69, 186, 793]. Approaches [413, 315, 326]. Approaching [715]. Approximation [493]. Approximations [756]. Ardilly [581]. Area [572, 141, 823, 213, 633, 174, 282, 446, 573]. Areas [635]. Arising [155]. Aspects [125, 387]. Assess [808]. Assessing [348, 632, 95]. Assessment [572, 76, 845, 320, 549, 223]. Assessor [160]. Asset [658]. Assets [721]. Assisted [451]. Association [386, 476, 672, 459]. Associations [724]. Assortative [723]. Assumptions [50, 776, 183]. Assurance [808]. Asymmetric [624]. Asymptotics [710]. Atkinson [709, 183]. Atlas [766]. Atmospheric [306]. Attainment [137, 611]. Attempts [478]. Attitude [34, 361]. Attitudes [670, 503, 9]. Attitudinal [95]. Attract [162]. Attributes [160]. Attrition [471]. Auditing [72]. Augmentation [482]. aureus [775]. Austin [393]. Australia [78]. Author [486]. Authority [42]. Authors [129, 169, 277, 35, 84, 408, 534, 661, 768, 872]. Automatic [212, 205, 303]. Available [482]. Averaging [724, 802, 858]. Avoiding [636]. Award [700]. Aylin [105].

B [864, 811, 592, 736, 593, 757, 766, 554, 461, 586, 813, 494, 646, 642, 689, 694, 620, 870, 836, 490, 648, 765, 737, 460, 856]. Back [14, 24, 37, 59, 73, 86, 102, 116, 131, 142, 156, 171, 184, 214, 238, 262, 296, 328, 364, 394, 423, 457, 487, 514, 552, 579, 615, 639, 695, 716, 740, 769, 796, 819, 842, 873]. bacteraemia} [775]. Bad [281, 725]. Baesens [864]. Bailer [437]. Bailey [234, 829, 459]. Balanced [416]. Balding [145]. Ballas [757]. Barker [461]. Barnett [834]. Baron [684]. Baseball [258]. Based [678, 719, 571, 675, 22, 63, 622, 785, 800, 793, 743, 779, 803, 720]. Basic [864, 618, 713, 813]. Basics [560]. Bayes [691]. Bayesian [372, 312, 838, 634, 165, 543, 64, 211, 724, 719, 472, 327, 826, 418, 748, 447, 777, 257, 315, 301, 341, 528, 647, 649, 462, 856, 657]. BBC [718]. Be [40]. Beasley [642]. Beaumont [582]. Become [452, 476]. Bed [8]. before [512]. Behavioral [732, 733]. Behaviour [254, 176, 574, 8]. Beijing [353, 354, 293]. Being [610, 358, 420]. Beirlant [425]. Belgium [704]. beneath [290]. Beneficial [20]. Benefit [284]. Benzene [54]. Berk [855]. Berman [643]. Best [545, 162, 57, 105, 49]. Best-Qualified [162]. Beta [527]. Better [755]. between [538, 670, 453, 176, 417, 724, 285, 706, 675, 77, 476, 11]. Beveridge [500]. beyond’ [207, 527]. Bhattacharya [754]. Bias [723, 54, 254, 629, 19, 674, 319, 389, 677, 501, 545, 476, 529, 180, 774]. Biase [677]. Biased [551, 7, 779]. Biases [510, 53, 826]. Biedermann [657]. Billion [844]. Binary [388, 773, 68, 9, 257]. Binomial [734, 693]. Bioequivalence [760]. Bioinformatics [525, 430, 835, 739, 186, 339, 198]. Biology [566, 763, 620]. Biomedical [316]. Biosecurity [318]. Biospheric [671]. Biostatistics [713, 369, 814, 441, 351, 316]. Bird [200]. Birth [17, 287, 826, 679]. Birthday [464]. Births [125]. Bjerve [261]. Boger [460]. Boland [812]. Bollen [685]. Book [864, 526, 811, 185, 215, 239, 263, 351, 297, 329, 365, 395, 523, 458, 424, 440, 488, 515, 553, 580, 616, 640, 732, 682, 707, 727, 752, 788, 810, 831, 854, 224, 222, 221, 374, 562, 592, 656, 792, 862, 276, 314, 464, 557, 251, 869, 736, 835, 186, 250, 270, 376, 315, 377, 871, 266, 275, 428, 432, 518, 530, 593, 525, 834, 839, 402, 301, 339, 429, 430, 520, 493, 560, 818, 264, 187, 310, 735, 380, 378, 379, 468, 492, 686, 405, 242, 336, 371, 761, 127, 623, 643, 757, 766]. Book [868, 527, 341, 624, 563, 565, 316, 372, 426, 427, 466, 517, 556, 581, 618, 352, 463, 554, 313, 271, 218, 219, 332, 338, 381, 337, 407, 434, 435, 461, 497, 564, 584, 596, 597, 598, 795, 685, 712, 832, 858, 650, 651, 400, 366, 367, 713, 658, 566, 866, 467, 273, 267, 245, 330, 304, 331, 528, 865, 755, 521, 533, 586, 762, 790, 243, 399, 404, 867, 240, 189, 298, 299, 491, 759, 813, 369, 431, 558, 559, 621, 731, 191, 193, 202, 217, 216, 246, 691, 690, 199, 200, 814, 190, 306, 516]. Book [519, 660, 739, 860, 645, 684, 738, 398, 494, 659, 687, 220, 591, 655, 840, 346, 583, 833, 406, 646, 433, 312, 201, 403, 83, 485, 760, 622, 642, 647, 709, 859, 758, 58, 268, 333, 334, 373, 397, 340, 348, 734, 794, 555, 244, 585, 714, 763, 269, 274, 303, 307, 436, 437, 469, 532, 793, 617, 654, 689, 692, 396, 302, 459, 582, 589, 649, 812, 863, 653, 694, 837, 841, 529, 265, 305, 438, 439, 496, 561, 619, 620, 710, 861, 247, 344, 309, 342, 345, 711, 248, 195, 196, 347, 375, 300, 272, 715, 441, 465]. Book [524, 594, 733, 764, 815, 816, 855, 857, 241, 198, 311, 425, 368, 223, 192, 641, 870, 194, 495, 595, 401, 308, 343, 489, 531, 588, 590, 652, 693, 754, 756, 836, 182, 335, 203, 462, 370, 730, 490, 587, 648, 765, 522, 567, 644, 688, 737, 729, 460, 791, 817, 856, 249, 197, 188, 349, 350, 657, 838]. Bootstrap [431, 620]. Borenstein [529]. Both [323]. Boundaries [326, 299]. Bounded [527]. Bounds [776]. Box [686, 449]. Bradley [22]. Brain [726]. Brams [755]. Bratti [232]. Braun [793]. Brazzale [710]. Breaking [452, 326]. Breast [165, 450, 385, 743]. Breeding [416]. Briggs [596]. Bristol [104]. Britain [757, 766, 506, 804, 721, 88, 576, 718, 18, 12, 292, 70, 16]. British [17, 81, 168, 359, 284, 479, 162, 175]. Brown [555]. BSc [455]. Buccianti [687]. Buck [654]. Bucklew [427]. Building [2, 115]. Built [772]. Buncher [516]. Business [860]. Butler [756]. By-Products [826].

C [464, 520, 818, 735, 787, 712, 650, 621, 516, 840, 583, 709, 654, 649, 439, 561, 710, 711, 441, 465, 489, 644, 688, 729, 657]. C. [787]. Calculating [742]. Calculations [729]. Calculus [836, 191]. Calls [781]. Can [725]. Cancer [136, 178, 165, 56, 450, 20, 43, 385, 743, 803]. Capital [593, 576, 864]. Capitation [62]. Capture [504]. Capture-Recapture [504].

Captures [612]. Carbon [671, 550]. Cards [419]. Care [315, 255, 627, 283]. Careers [547, 95]. Carlin [856]. Carlo [620, 824, 414]. Carmona [517]. Carroll [711]. Case [221, 178, 783, 289, 141, 705, 360, 710, 68, 114, 282, 743]. Case-Study [178, 289, 141, 68, 114, 743]. Casella [857]. cases [45]. Categorical [469, 418, 641, 462]. Catholics [30]. Causal [453, 372, 106, 706, 668, 737]. Celebrating [464]. Cell [575]. Censoring [545]. Census [572, 511, 576, 326, 210, 569, 800, 628, 206, 124, 575, 207, 75]. Census-Based [800]. Censuses [122]. Century [321]. Cervical [20]. Chain [670, 824, 414]. Challenge [744]. Chambers [57]. Chance [712]. Chang [644, 688]. Change [254, 28, 777, 672]. Changepoint [680]. Changes [254]. Changing [761]. Chappelle [811]. Characteristic [759]. Characteristics [10, 825, 391]. Characterization [608]. Charalambides [489]. Charts [235, 775]. Chaudhuri [617]. Checking [50]. Checks [307]. Chemistry [345]. Child [357, 597, 598, 475]. Childbearing [506]. Childhood [54, 177, 359, 825, 476]. Children [845, 705, 43]. China [353, 179, 354, 293]. Choice [851, 700, 445]. Choosing [538]. Chow [644, 688, 729]. Chronologies [299]. Cities [576, 27]. City [114]. Cizek [518]. Claeskens [858]. Class [96, 503, 807, 680, 725, 179, 30, 746, 175]. Classification [538, 496, 573]. Classified [160, 140]. Claude [859]. Cleavages [161]. Clerk [384]. Clients [504]. Climate [718, 777]. Clinical [237, 113, 419, 315, 178, 4, 369, 519, 7, 583, 79, 714, 342, 764, 644, 688, 729, 183, 869, 565]. Close [308]. clues [45]. Cluster [808, 42]. clustered [45]. Clustering [508, 571, 47, 637]. Clusters [40, 39, 41, 48]. Co [33, 89]. Co-Operation [33]. Co-Ordinating [89]. Coastline [43].

Codes [604]. Coding [630, 591]. Coefficient [289, 63]. Cohabitation [17, 18, 785, 16]. Cohabiting [505]. Coherence [675]. Cohort [17]. Cointegrated [592]. Cold [718]. Coleman [860]. College [290, 454]. Colman [689]. Colorectal [178]. Combinatorial [489]. Combinatorics [459, 712]. Combine [826]. Combined [547]. Combining [27, 635, 673, 809, 210, 628]. Commandeur [730]. Comments [232, 237, 234, 57, 44, 49, 52]. Commissioned [104]. Committees [5]. Communicating [149, 253, 155]. Communication [152, 146, 147, 153]. Communities [70]. Community [479, 391, 483, 421]. Commuting [575]. Companion [758]. Comparative [610, 649]. Comparing [419, 123]. Comparison [353, 113, 374, 121, 480, 807, 606, 19, 255, 285, 29, 78, 800, 575, 293, 354]. Comparisons [138]. Compensation [742]. Competing [17]. Competition [158]. Complex [564, 606, 290, 507]. Complexity [712, 822]. Compliance [93, 605]. Component [47, 335]. Components [864, 189]. Composite [320]. Composition [324]. Compositional [687]. Comprehensive [141]. Computational [832, 794, 194, 520, 492, 867, 791]. Computer [521, 684, 451, 808]. Computer-Assisted [451]. Computing [526]. Concepts [864, 713]. Concise [314]. Conditional [137, 546]. Conditions [11]. Conducting [203]. Confidence [287]. Confidential [809]. Conflict [747]. Conflicting [748]. Congdon [462]. Conglomerates [593]. Conscious [850]. Consent [502]. Consequences [850]. Consistent [848]. Constrained [495]. Constructing [299, 578]. Construction [848]. Consultations [570]. Consumer [338]. Consumption [31]. Contact [801, 478]. Contemporary [10]. Contents [279, 663, 410]. Context [161]. Contingency [807]. Continuation [285]. Continuing [95]. Continuous [527, 285, 283]. Contraception [166]. Contraceptive [82, 166]. Contract [509]. Control [630, 775, 550, 345, 377]. Controlled [124]. Controlling [473]. Conversion [474]. Coronary [149]. Correct [482]. Correcting [19]. Correction [101]. Correlated [166]. Correlation [849]. Correlations [209, 776]. Correspondence [128, 486, 484, 183, 591]. Corrigendum [353, 767, 13, 168, 115, 354]. Cost [165, 596]. Cost-Effectiveness [165, 596]. Costs [473]. Council [638, 260, 392, 750, 852, 513]. Countries [653]. Country [724]. Couper [380]. Couples [323, 452]. Course [314, 191, 823, 689, 140, 432]. Covariate [113]. Covariates [67, 34, 746]. Coverage [607]. Cox [464, 557, 463, 449, 690]. Cradle [766]. Cradle-to-Grave [766]. Crainiceanu [711]. Crash [22, 689]. Crawley [428]. Creating [573]. Credit [864]. Cricket [259, 512]. Crime [176, 635, 11, 153]. Crimes [11, 446]. Criminal [544, 176]. Criteria [538, 183]. Cross [121, 122, 160, 140]. Cross-Classified [160, 140]. Cross-National [121, 122]. Crossing [299]. Crowley [519]. Cumulative [702, 450]. Curran [685]. Current [413]. Curve [685]. Curves [846, 759]. custodies} [72]. custodiet [72]. Cycle [9].

D [557, 525, 757, 766, 463, 650, 566, 621, 691, 690, 519, 860, 583, 760, 642, 709, 654, 561, 711, 441, 733, 531, 836, 765, 737, 838]. Dagnelie [645]. Daily [27]. Dalgaard [861]. Daneshkhah [654]. Darby [57]. Data [636, 419, 221, 374, 862, 209, 54, 608, 572, 301, 178, 480, 336, 761, 33, 56, 543, 64, 372, 517, 463, 482, 388, 81, 168, 89, 674, 454, 719, 319, 285, 284, 467, 331, 865, 665, 673, 699, 450, 292, 200, 809, 659, 687, 591, 605, 646, 179, 773, 31, 213, 210, 626, 826, 109, 208, 418, 748, 692, 507, 649, 628, 679, 439, 342, 20, 764, 104, 641, 666, 575, 69, 144, 856, 361, 803, 45, 523, 596, 437, 469, 524, 462, 460, 839, 624, 193, 583, 793]. Data-Inclusive [608]. David [786]. Davidian [862]. Davison [464, 710]. Day [259, 420]. Death [508]. Debt [721]. Decayed [3, 13]. December [750]. Decision [413, 21, 802, 188]. Decision-Making [413]. Decisions [506, 512, 188]. Decline [70]. Deficiency [773]. Degree [140]. Demands [61]. DeMets [583]. Democracy [755]. Demographic [112, 114]. Demography [500]. Density [220]. Dental [3, 13]. Deonier [520]. Dependence [508, 511, 1, 23]. Dependent [746]. Deprivation [108]. Derivatives [322]. Design [222, 93, 435, 674, 473, 7, 840, 642, 705, 644, 688, 720, 183, 461, 857, 521, 790, 587]. Design-Based [720]. Designed [459]. Designing [755, 374]. Designs [113, 709, 808, 790]. Detailed [125, 628]. Detecting [544, 47, 67, 679, 257]. Detection [209, 212]. Determinants [112, 358]. Determination [543, 447]. Developing [653]. Development [869, 871, 357, 33, 598, 475, 760, 48, 597, 817]. Diagnosed [719]. Diagnostic [2, 115, 307]. Diagnostics [63]. Dialling [110]. Diaries [31]. Dice [794]. Dictionary [219]. Dietary [546]. Differences [631, 29, 822]. Different [781, 501]. Differentials [722]. Difficulties [122]. Digit [110]. Digital [794]. Dilution [19]. Dioxide [550]. Direct [805]. Directions [91]. Disciplinary [551, 299, 705]. Disclosure [630, 76, 549]. Discontinuation [166]. Discontinuity [93]. Discrete [523, 824, 327, 632, 700, 524, 489]. Discriminant [375]. Discussion [145, 207, 228, 107, 742, 105, 718, 719, 119, 134, 743, 227]. Disease [701, 40, 724, 673, 46, 39, 41, 45]. Diseases [47]. Disinfection [826]. Disorder [477]. Display [336]. Disposition [604]. Disruption [610]. Dissuade [148]. Distribution [850, 656, 482, 745, 500]. Distributions [735, 527, 762, 258, 489, 765]. Divergence [622]. Diversity [574]. Division [755]. Divisive [637]. Divorcing [452]. DNA [167, 23, 144, 840, 642]. Do [452, 647, 626, 159]. Dobson [834]. Dodge [464]. Dominic [712]. Donald [295]. Donev [709]. Donor [124]. Don’t [420]. Dorling [757, 766]. Dorofeev [584]. Dorp [527]. Dorset [42]. Doubly [480]. Dr [383]. Draper [232, 686]. Drinking [746]. Driver [22]. Drop [181]. Drop-out [181]. Dropping [71]. Drug [96, 760, 817]. Dudzic [758]. Dugard [715]. Duration [749, 81, 168, 785, 481]. Durations [82]. Dutch [720]. Dybowski [525]. Dynamic [659]. Dynamical [754]. Dynamics [658, 667].

Early [505]. Earnings [357, 742, 607]. Eastern [387]. Ecological [54, 55, 56, 606, 50, 674, 776, 53, 108, 501, 43, 39, 51, 227, 433]. Economic [864, 610, 749, 33, 90, 112, 673, 802, 11, 358, 593]. Economics [724]. Economy [848]. Ecosystem [550]. ed [526]. Edgington [731]. Edit [205]. Editing [210]. Editorial [499, 226, 537, 318, 821, 88, 205, 133, 4, 772, 173, 89, 697, 665, 356, 844, 253, 471, 569, 61, 626, 16, 118, 158, 75, 26, 159, 39, 600, 412]. Edits [208]. edn [493, 565, 618, 516, 519, 555, 763, 617, 589, 490, 587, 374, 186, 377, 301, 304, 312, 302, 197]. eds [518, 525, 516, 519, 529]. Educated [844]. Education [232, 357, 231, 453, 230, 290, 287, 141, 162, 67, 570, 95]. Educational [603, 631, 822]. Edwards [642]. Effect [357, 631, 823, 785, 849, 20, 611, 575, 95]. Effective [577]. Effectiveness [474, 165, 596]. Effects [636, 93, 781, 122, 136, 749, 323, 290, 285, 780, 160, 292, 822, 141, 483, 668, 619, 805, 166, 114, 420, 446, 737]. Efficiency [767, 824, 703, 356, 325, 120]. Ehrenberg [99]. Eiser [654]. Elderly [628, 283]. Electoral [324, 9]. Elementary [738]. Elements [196, 656]. Elicitation [414, 677]. Eliciting [654, 45]. Elliott [44, 693]. Embedded [720]. Emergence [387]. Emergency [781]. Empirical [608, 291, 574]. Empirically [779]. Employed [95]. Employment [506, 231, 509, 547, 91]. End [390, 6, 178, 79]. Engineering [435]. Engineers [735]. England [386, 56, 112, 671, 778, 628, 20, 358, 421]. English [551]. Enhancing [570]. Enterprise [813]. Environmental [54, 437, 841, 114, 332, 217]. Epidemic [209, 387]. Epidemiological [126, 309]. Epidemiology [413, 3, 13, 572, 316, 713, 305]. Equation [685, 611]. Equations [773]. Error [96, 212, 497, 540, 126, 711]. Errors [473, 607]. Essential [643, 643]. Essentials [533]. Estimate [451, 748, 628]. Estimated [548]. Estimates [254, 719, 675, 633, 174]. Estimating [742, 28, 22, 179, 77, 773, 548, 747, 476, 504, 80, 324, 868]. Estimation [824, 480, 139, 220, 201, 213, 309, 68, 594, 140, 870, 220]. Estimations [308]. Estimator [388, 612]. Estimators [479]. Ethics [5]. Ethnic [29, 70]. Ethnicity [722]. Eubank [585]. EUREDIT [205]. Europe [288]. European [749, 479, 61, 391, 483]. Evaluate [846, 784]. Evaluating [357, 380, 205]. Evaluation [96, 88, 167, 90, 359, 780, 725, 501, 94, 550, 496, 180, 385, 676, 315]. Evaluations [91]. Evans [105]. Event [427, 327, 69]. Everitt [586, 736, 813, 646]. Every [745]. Evidence [413, 804, 721, 167, 27, 801, 502, 391, 483, 748, 20, 774, 779, 551]. Evolution [258, 468]. Evolutionary [340, 144]. Ewens [430].

Exact [523, 524]. Examination [98]. Examining [414]. Example [106, 476, 793]. Example-Based [793]. Examples [385, 304]. Excess [744]. Exercises [581]. Exercising [643]. Exhibit [362]. Exit [718]. Expectancies [360]. Expectancy [742, 139, 675]. Expected [743]. Expenditure [448]. Experience [718, 505]. Experiences [572, 604]. Experimental [461, 790, 840, 709, 210, 222, 818]. Experimentalists [706]. Experiments [374, 435, 790, 539, 459, 91, 42, 587, 720, 521, 642]. Expert [677, 654]. Explanatory [166]. Exploration [724]. Exploratory [467]. Exploring [288, 358, 669, 331]. Exponential [735]. Exposure [54, 22]. Exposures [476]. Expression [647]. Extended [22]. Extending [449, 91, 522]. Extent [153]. Extremes [425].

F [523, 871, 463, 497, 491, 659, 591, 620, 465, 524, 816, 730, 657]. Factor [361]. Factors [673]. Fahmy [757]. Failed [478]. Failure [680]. Fair [755]. Fair-Division [755]. Fall [70]. Fallacy [108]. Families [527]. Family [506, 122, 359, 284, 450, 784, 637, 547]. Fang [521]. Faraway [522]. Farewell [237]. Fatal [12]. Fatalities [154]. Fatality [353, 293, 354]. Faulty [183]. Fayers [107]. FBCS [455]. FCIS [455]. Fenn [614]. Field [341]. Fields [623]. Fight [653]. Figueras [687]. Filled [3, 13]. Filter [585]. Finance [518, 310, 194, 590, 693, 836, 347]. Financial [864, 593, 721, 517, 191, 660, 632, 350, 300]. Finding [545]. Findings [603, 114]. Finkenstädt [648]. Finland [604, 360]. Firm [772]. Firms [723, 593, 289]. First [17, 65, 137, 327, 601]. First-in-Man [601]. First-Partnership [17]. FIS [455]. Fisheries [605]. Fission [491]. Fitting [604]. Fitzmaurice [862, 556]. Flexibility [776]. Flexible [699]. Flows [10, 628]. Flux [671]. Fluxes [550]. Foot [724]. Foot-and-Mouth [724]. Football [551].

Force [538, 670, 10, 29, 213, 633, 720].

Forecast [306]. Forecasting [782, 697, 778]. Forensic [23, 549, 144, 657]. Foresight [217]. Foreword [229, 147]. Formation [17, 505]. Forward [331]. Foundations [772]. Founded [289]. Fractional [2, 115]. Framework [627, 174, 51]. Francis [490]. Fraud [451]. Frederick [614]. Free [141]. Frequency [201]. Friedman [583]. Friendly [659]. Friendship [804]. Front [15, 25, 38, 60, 74, 87, 103, 117, 132, 143, 157, 172, 204, 225, 252, 280, 317, 355, 382, 411, 444, 470, 498, 536, 568, 599, 625, 664, 696, 717, 741, 770, 797, 820, 843, 874]. Functional [221]. Functions [450]. Fund [421]. Fundamental [218]. Funding [62]. Furberg [583]. Future [742, 500, 569, 206].

G [862, 834, 430, 493, 492, 686, 556, 712, 858, 528, 533, 813, 621, 687, 646, 433, 642, 582, 561, 764, 857]. Galbraith [491]. Gałecki [694]. Gao [863]. Gap [188]. Garbolino [657]. Garthwaite [654]. Gaussian [308, 546, 623]. Gearing [455]. Geisser [294]. Gelman [52]. Gender [670]. Gene [647]. General [502, 10]. Generalized [233, 834, 290, 773, 828, 619, 567]. Generation [303]. Generations [97]. Genome [520]. Genomics [642]. Gentleman [835]. Geographical [804, 54, 326, 46, 387]. geographically [45]. Geohazards [155]. Geometric [379]. George [786, 455]. Geosciences [687]. Geostatistics [841]. Geriatric [8]. Germ [42]. Germany [721]. Gestel [864]. Ghettos [574]. Giants [744]. GIS [761]. Gittoes [232]. Givens [492]. Glahn [687]. Glidden [441]. GLIM4 [490]. Godfrey [806]. Goegebeur [425]. Golan [868]. Goldstein [691]. Gönen [759]. Good [281, 725, 431]. Gordon [757]. Gotway [439]. Governance [724]. Government [234, 499, 801]. Graduate [432, 231]. Graduates [799]. Graft [680]. Grant [160, 430, 584]. Graphical [670, 826]. Graphics [793, 563, 659]. Graphs [379]. Grave [766]. Gravity [672]. Gray [650]. Great [506, 804, 721, 18]. Greenfield [860]. Gregory [649]. Grieve [383]. Grigg [237]. Grimmett [712]. Grouped [285, 803]. Groups [669]. Growing [61]. Growth [597, 598, 825]. Gruijter [733]. Guidance [95]. Guide [497, 304, 813, 306, 348, 694, 192]. Guidelines [4]. Gut [432]. Guthrie [57].

H [792, 493, 492, 623, 624, 556, 497, 596, 713, 866, 555, 763, 617, 529, 729]. Hagan [654]. Hald [751]. Hand [557, 107]. Handbook [736, 586, 867, 526, 735, 367, 762, 202, 519, 833, 692, 203]. Handling [31]. Happier [452]. Härdle [518]. Harold [422, 455]. Hart [763]. Hauschke [760]. Having [148]. Hazard [181]. Hazardous [634]. Hazards [152, 285]. Heady [393]. Health [351, 577, 315, 572, 453, 749, 152, 597, 795, 775, 802, 479, 108, 845, 744, 236, 164, 627, 134, 42, 421, 676, 494]. Health-Care [315]. Health-Related [479, 676]. Healthy [675, 844]. Heavy [545]. Held [648, 623, 49]. Hemisphere [542]. Hendrik [853]. Herds [704]. Here [18]. Herzberg [557]. Hesketh [736]. Heterogeneity [508, 92]. Heterogeneous [94]. Heuristics [677]. Hewson [234]. Heyde [787]. Hibbert [456]. Hidden [545, 198]. Hierarchical [27, 673, 627, 418, 777]. High [723, 782, 704, 88, 503]. Higher [232, 231, 230, 162]. Highly [373]. Hilbe [734]. Hillier [761]. Hinde [490]. Hinkelmann [790]. Hirji [523, 524]. Historical [761, 543, 179]. Histories [144, 135]. History [761, 578]. Hitting [781]. Hjort [858]. Hoek [693]. Hoeting [492]. Holistic [547]. Hong [353, 78, 354, 293]. Honoré [613]. Honour [464]. Hoshmand [587]. Hospital [421]. Hothorn [586]. Hourly [482]. Hours [292].

House [537, 141]. Household

[96, 721, 541, 801, 10, 359, 473, 502, 479, 179, 391, 483, 632, 446]. Households [541, 800]. Housing [547]. Human [576, 218, 719, 748, 148, 805, 203]. Hunger [653]. Hunt [532]. Husmeier [525]. Hypotheses [416, 431, 589]. Hypothesis [363].

Ideas [381, 578, 549]. Identification [704, 549]. Identifier [548]. Identifying [627]. Identities [161]. Identity [175, 766]. If [40]. Ignorable [482, 388, 478]. Ignored [510]. Ignoring [849]. II [557]. Ill [108]. Ill-Health [108]. Illustrated [33]. Illustration [291]. Immigrant [722]. Immigrants [418]. Immune [773]. Immunization [180]. Immunodeficiency [719, 748, 148, 805]. Impact [92]. Imperfect [93]. Implementation [223]. Implications [510, 284]. Importance [150, 95]. Impotence [416]. Improve [824, 680]. Improved [68]. Improvement [222]. Improving [174]. Imputation [630, 480, 205, 211, 809, 210, 483, 124]. Imputed [291]. Inaccuracies [54]. Incidence [551, 450, 675, 20, 43]. Incidence-Based [675]. Inclusive [608]. Income [807, 284, 483]. Incomes [321]. Incomplete [82, 209, 372]. Inconsistent [551]. Increase [743]. Increasing [577]. Independent [118]. Index [35, 84, 129, 169, 277, 408, 534, 661, 768, 872, 3, 13, 338, 109, 443]. India [508]. Indicator [106]. Indicators [232, 234, 233, 281, 231, 230, 320]. Indices [537, 574]. Individual [56, 673, 668]. Individuals [807]. Indrayan [814]. Induced [22]. Industry [151, 516, 860]. Inequalities [577]. Inequality [386, 256, 745, 495]. Infant [386]. Infarction [419]. Inference [55, 167, 372, 776, 706, 433, 622, 647, 11, 311, 495, 657, 227, 562, 533, 690, 312, 465, 314].

Inferences [144]. Inflated [3, 13]. Influence [177]. Inform [851]. Informatics [525]. Information [36, 85, 130, 170, 278, 409, 442, 535, 662, 771, 875, 113, 150, 578, 743, 95, 600, 188]. Informative [414]. Initiatives [91]. Injustice [114]. Input [848]. Input-Output [848]. Inquiry [104]. Insights [576, 123, 477]. Institutional [283]. Instrument [287]. Insurance [518, 590, 310]. Integrals [530]. Integrated [592]. Integrating [122, 326]. Integration [578]. Intensity [747]. Interactive [659]. Intercept [800]. Internal [121]. International [603, 138]. Internet [564]. Interobserver [313]. Interpretation [642, 692]. Interpreting [23]. Intervals [287]. Interventions [805, 743]. Introduction [428, 834, 146, 463, 790, 715, 641, 730, 567, 430, 520, 339, 427, 621, 739, 763, 465, 870, 595, 836]. Introductory [562, 861, 590]. Invariance [588]. Inverse [479]. Inverse-Probability-Weighted [479]. Investigated [40]. Investigating [701, 112, 284, 420, 41]. Investigation [176, 43]. Invited [232, 237, 234]. Involving [390]. ipsos [72]. Iraq [747]. Irish [454, 848]. Isham [648]. issue [458]. Issues [121, 572, 576, 33, 845, 94, 236, 158, 601, 817]. Item [725, 34]. ITV [718].

J [526, 557, 428, 834, 430, 493, 380, 492, 757, 868, 527, 427, 466, 556, 618, 795, 685, 658, 566, 755, 558, 516, 519, 660, 655, 833, 642, 859, 734, 794, 437, 469, 793, 654, 589, 812, 863, 529, 619, 620, 711, 715, 465, 733, 816, 425, 495, 693, 730, 490, 522, 729]. Jack [456]. Jakob [261]. James [393]. Janke [465]. Java [591]. Jenkinson [654]. Jensen [554]. Jin [735]. JMP [738]. Job [65, 609, 420, 361]. Joblessness [177]. Jobs [137]. John [829, 44, 830, 422]. Join [803]. Joint [680, 47, 773, 777]. Judgements [654]. Judgements-Eliciting [654]. July [750, 387]. Jump [300]. Jurecková [558].

K.-A [647]. K.-T [521]. Kallen [791]. Kallenberg [588]. Kalman [585]. Kamp [733]. Karp [868]. Kempthorne [790]. Kemsley [614]. Kendall [786, 312]. Kenward [764]. Kerry [655]. King [433]. Kinnear [650]. Kleinman [494]. Knorr [49]. Knorr-Held [49]. Knowledge [600]. Knowles [761]. Koenker [559]. Kong [353, 78, 354, 293]. Konijn [853]. Kontoghiorghes [526]. Koopman [730]. Kotz [527]. Krause [560]. Krishnamoorthy [762]. Krishnan [652]. Kroonenberg [865]. Kulldorff [49]. Kushner [493].

L [530, 623, 868, 858, 467, 533, 867, 583, 622, 585, 714, 589, 711, 733, 648]. Laboratory [345]. Labour [538, 670, 122, 609, 323, 29, 94, 213, 633, 720, 10, 29]. Laird [556]. Lamberton [836]. Lapeyere [836]. Large [564, 140, 223]. Large-scale [564, 223]. Larger [576]. Largest [27]. Last [532]. Latent [544, 96, 678, 503, 680, 725, 179, 34, 632, 746, 669, 685]. Law [99, 672]. Law-like [99]. Lawson [494]. League [258, 851, 551]. Learning [782, 186, 832, 605, 855, 197, 811, 196]. Least [137]. Lee [619]. Leeuw [833]. Leg [512]. Legislation [163]. Lehmann [589]. Length [283]. Lens [414]. Lessler [380]. Lessons [104]. Leukaemia [54, 45]. Level [724, 454, 138]. Levy [653]. Li [521]. Life [742, 106, 139, 540, 675, 823, 360, 584]. Life-Tables [675]. like [99, 420]. Likelihood [220, 619]. Lim [735]. Limitations [851]. Limited [723, 679]. Limits [325]. Lin [590]. Lindsey [466]. Linear [233, 834, 290, 674, 449, 479, 621, 691, 619, 465, 816, 522, 828, 694, 370]. Linguistics [832]. Link [415, 613]. Linkage [476]. List [683, 708, 728, 753, 789]. Literacy [123]. Littell [561]. Local [234, 775, 141]. Locally [746]. Logical [649]. Logistic [362, 503, 137, 197]. Logit [633]. London [100]. Long [474, 1, 18, 477, 539, 542, 159, 283]. Long-Range [1]. Long-Term [474, 477, 539, 283]. Longitudinal [474, 556, 388, 89, 784, 773, 418, 726, 666, 862]. Look [785]. Loss [742]. Lot [808]. Lottery [850]. Louis [856]. Low [723, 826]. Lung [56]. Lupton [757]. lymphoma [45].

M [862, 557, 428, 520, 560, 380, 643, 868, 556, 435, 866, 528, 865, 759, 691, 684, 659, 583, 433, 642, 647, 734, 841, 529, 496, 711, 733, 764, 495, 754, 490, 644, 688]. M. [703]. MA [614]. MacDonald [713]. Macdougall [295]. Machine [186]. Madagascar [67]. Made [650, 651]. Madsen [792]. Magic [798]. Mahmoud [866]. Maindonald [793]. Mainline [12]. Major [258]. Majumdar [754]. Making [413, 323]. Malawi [327]. Male [125, 504]. Malnutrition [808]. Man [601]. Management [864, 593, 350]. Managers [643]. Mandel [830]. Manifesto [217]. Manipulation [839]. Manly [620]. Manual [338]. Map [795]. Maps [761]. Marginal [775, 699]. Marijuana [503]. Marital [610, 505]. Marked [200]. Marker [416]. Market [122, 609, 868, 323, 29, 94, 368, 159]. Market-Place [159]. Marketing [528]. Markov [623, 198, 544, 824, 165, 414, 632, 283]. Marks [612]. Marriage [17, 327, 785]. Married [175]. Marshall [237]. Martin [380]. Martinez [467, 867, 467, 867]. Martingales [530]. Match [144]. Matched [737]. Matching [723, 94, 785]. Mateu [687]. Mateu-Figueras [687]. Mathematica [649]. Mathematical [339]. Mathematics [758, 755]. MATLAB [467, 867]. Matter [14, 24, 37, 15, 25, 38, 59, 73, 86, 60, 74, 87, 102, 116, 131, 103, 117, 132, 142, 156, 171, 143, 157, 172, 184, 214, 238, 262, 204, 225, 252, 280, 296, 328, 364, 394, 317, 355, 382, 411, 423, 457, 487, 514, 444, 470, 498, 536, 552, 579, 615, 639, 568, 599, 625, 664, 695, 716, 740, 769, 696, 717, 741, 770, 796, 819, 842, 873, 797, 820, 843, 874, 175, 448]. Maximum [220]. MBE [422]. McCleery [763]. McCulloch [528, 441]. McDiarmid [712]. Me [420]. Measure [767, 538, 703, 162, 120]. Measurement [499, 781, 96, 678, 123, 218, 106, 109, 126, 711]. Measurements [825]. Measures [121, 359, 255, 784, 448, 622, 530, 313]. Measuring [847, 417, 745, 253, 325]. Medical [413, 181, 525, 655, 646, 496, 192, 814]. Medicine [871, 555]. Meert [795]. meeting [207]. Meijer [833]. Memory [542]. Men [175]. Meta [780, 849, 69, 779, 529]. Meta-Analysis [780, 849, 69, 779, 529]. Metastatic [726]. Methicillin [775]. Method [22, 77, 668, 504, 447]. Methodological [565, 845, 433]. Methodologies [501]. Methodology [351, 821]. Methods [221, 357, 339, 430, 468, 55, 212, 205, 123, 581, 463, 332, 597, 598, 19, 414, 369, 558, 94, 79, 760, 210, 164, 111, 812, 439, 620, 441, 489, 648, 644, 688, 45, 676, 856, 691, 617, 863, 380, 190]. Methylation [840]. Microanalyses [483]. Microarrays [840, 642]. Microdata [122, 291, 282]. Microlevel [391]. Middle [139]. Might [767, 703, 120]. Migrant [800]. Migration [121, 122, 628, 575, 135]. Millennia [542]. Milliken [561]. Mind [574]. Minimum [256]. Mining [193]. Minitab [818]. Misconceptions [53]. Missing [781, 3, 13, 480, 699, 179, 67, 773, 31, 213, 626, 764]. Missingness [388]. Misunderstandings [706]. Mixed [636, 233, 290, 449, 555, 633, 694, 561]. Mixture [546, 126, 827]. Mixtures [686, 23]. Mobile [604, 152]. Mobility [723, 577, 804, 807, 631, 669]. Model [415, 181, 233, 374, 544, 670, 678, 3, 13, 609, 634, 218, 606, 724, 680, 50, 858, 285, 449, 571, 802, 47, 22, 77, 356, 327, 164, 633, 826, 445, 418, 140, 522, 283, 613, 356, 858]. Model-Based [678, 571]. Modeling [525, 372, 521, 202, 765, 340]. Modelled [510]. Modelling [233, 850, 821, 1, 101, 453, 165, 27, 801, 546, 631, 825, 319, 566, 635, 475, 389, 699, 160, 450, 258, 542, 627, 632, 109, 92, 668, 667, 507, 300, 166, 611, 446, 774, 490, 803, 660]. Models [864, 636, 869, 357, 136, 362, 824, 290, 50, 685, 255, 802, 479, 289, 217, 822, 63, 555, 826, 34, 700, 746, 694, 126, 619, 711, 68, 2, 115, 465, 198, 666, 693, 8, 777, 335, 462, 370, 567, 779, 9, 803, 672, 828, 834, 866, 302, 561, 816, 531, 371, 348]. Modern [798, 711, 192]. modification [613]. Modular [140]. Molecular [468]. Molenberghs [764, 862]. Monahan [816]. Mondays [420]. Monitoring [237, 234, 229, 226, 702, 10, 583, 605, 236, 714, 48]. Monotone [388, 699]. Monte [620, 824, 414]. Montgomery [621, 860]. Mora [659]. Morgan [363]. Morphometrics [859]. Morrison [463]. Mortality [386, 64, 27, 389, 288, 823, 445, 679, 20]. Mother [476]. Mouth [724]. Movements [576, 822]. Moye [714]. Muenchen [837]. Mukhopadhyay [562]. Müller [647]. Multilevel [636, 65, 362, 821, 824, 801, 290, 825, 255, 160, 833, 327, 827, 507, 68, 140, 166, 611, 666, 446, 9, 98, 828, 302]. Multinomial [137, 633]. Multiparameter [413, 719]. Multiple [630, 234, 233, 480, 319, 809, 236, 826, 547, 777]. Multiple-Bias [319]. Multiple-sequence [547]. Multiply [291]. Multivariable [2, 115]. Multivariate [233, 209, 463, 825, 255, 331, 773, 692, 849, 715, 98, 343]. Multiway [865]. Murder [782]. Murray [105]. Murrell [563]. Murtagh [591]. Muth [618]. Myocardial [419]. Myth [798].

N [562, 464, 871, 735, 686, 556, 858, 709, 710, 733, 567, 838]. Nahin [794]. Nation [744]. National [850, 121, 96, 122, 133, 256, 89, 775, 475, 288, 705, 180, 573, 357, 164]. Natural [587]. Near [43]. Need [173, 163]. Needed [32]. Needs [795]. Negative [723, 734]. Neighbourhood [822, 114]. Neil [681]. Nelder [619]. Netherlands [137]. Networks [211, 571, 657]. Neurotic [477]. Neutral [512]. Never [42]. Newly [289]. Newton [672]. Niche [765]. Nielsen [468]. No [548]. Non [482, 801, 388, 475, 699, 479, 70, 67, 477, 471, 34, 607, 180, 478, 481]. Non-Contact [801]. Non-Ignorable [482, 388, 478]. Non-Linear [479]. Non-Monotone [388, 699]. Non-Random [67]. Non-Response [482, 475, 479, 477, 471, 34, 607, 180, 478, 481]. Non-White [70]. Nonlinear [304, 711, 863]. Nonparametric [863, 870, 595]. Nonuniform [303]. Normal [412]. Norman [829]. North [386, 387]. North-Eastern [387]. Northern [542]. notices [443]. Nuclear [125]. Null [416]. Number [511, 32, 124, 446, 478]. Numbers [825].

O [811, 790, 433, 654, 439, 561, 588]. Oakley [654]. OBE [295]. Obesity [778]. Obituaries [829, 294, 295, 456, 830, 455]. Obituary [393, 261, 681, 787, 786, 751, 853, 422, 614, 443]. Observational [319, 826, 342]. Observationalists [706]. Obtain [65]. Occupancy [8]. Occupational [137, 722]. Occupations [66]. October [387]. Offenders [176]. Official [66, 698]. Offspring [125]. Oliver [841]. Olofsson [531]. Olson [560]. Oncology [519]. One [259, 767, 511, 703, 141, 124, 120]. One-Day [259]. One-Number [511, 124]. O’Neill [795]. Onghena [731]. Operating [759]. Operation [33]. Operative [389]. Opinion [1]. Optimal [609, 674]. Optimum [709].

Order [495]. Ordered [290]. Ordinal [807]. Ordinary [137]. Ordinating [89]. Organization [33, 597]. Other [527]. Outcomes [259, 231, 177, 845, 773, 236, 505, 69, 9]. Outlier [209, 212]. Output [499, 848, 573]. Overcoming [53]. Overdispersion [362]. Overeducation [799]. Overtime [292]. Overview [235]. Owned [809]. Oxford [219].

P [383, 871, 518, 593, 839, 380, 686, 563, 581, 584, 795, 685, 650, 528, 865, 431, 731, 519, 645, 659, 642, 647, 794, 654, 582, 589, 649, 812, 861, 715, 495, 531, 462, 856, 657]. Packs [653]. Page [642]. Paid [292]. Pair [77]. Pal [735]. Palaeoclimate [472]. Panel [359, 665, 479, 292, 164, 667, 479, 391, 483]. Paper [145, 234, 228, 107, 105, 119]. Papers [232, 237, 57, 464, 557, 229, 44, 49, 52, 147]. Paradoxes [390]. Parallel [526, 56]. Parameter [308]. Parametric [450, 431]. Pardo [622]. Parental [177]. Parities [33]. Park [564]. Parliamentary [324]. Parolees [782]. Partial [180, 743]. Partially [179]. Participation [670, 112, 633]. Particular [385]. Partisanship [97]. Partitioning [362]. Partners [323]. Partnership [17, 177, 505]. Partnerships [505]. past [135]. Patents [158]. Pathways [453]. Patient [8]. Patients [726]. Pattern [546, 375]. Patterns [544, 137, 502, 179, 78, 679, 48]. Pawitan [619]. Pawlowsky [687]. Pawlowsky-Glahn [687]. Pay [482]. Peacock [655]. Peck [621]. Pedigrees [167]. Peer [160]. Penalized [220]. Pensioners [284]. Pensions [500]. People [177, 847, 795, 844, 148, 283]. Pepe [496]. Perceived [358]. Perception [667]. Perceptions [635, 151]. Perfect [743]. Performance [232, 237, 234, 233, 781, 226, 281, 231, 702, 230, 138, 627, 104, 229]. Period [43]. Periodic [46, 371]. Perloff [868]. Permutation [431]. Perspective [374, 565, 685, 711, 855, 838]. Perspectives [705, 372]. Petersen [612]. Petronis [840]. Petter [261]. Pfaff [592]. Pharmaceutical [618, 516]. Pharmaceuticals [158, 159]. Pharmaco [783]. Pharmaco-vigilance [783]. Pharmacokinetics [791]. Phase [72]. Phone [604]. Phones [152]. Physical [649]. Piantadosi [565]. Picek [558]. Piegorsch [437]. Pig [704]. Pig-Herds [704]. Pigeot [760]. Pioneer [806]. Place [757, 159]. Placing [761]. Plant [416, 125]. Plasma [322]. Plays [570]. Please [681]. PLUS [304, 560, 660, 517]. Point [800, 311, 803, 48]. Points [390, 6, 178, 79]. Poisson [3, 13]. Polarization [745]. Police [635]. Policies [90, 141]. Policy [577, 609, 643, 698]. Poliomyelitis [387]. Political [161, 1, 101, 97]. Poll [1]. Polling [718]. Polls [324]. Pollution [27, 445]. Polya [866]. Polynomial [308]. Polynomials [2, 115]. Pools [322]. Popularity [1, 101]. Population [782, 254, 167, 23, 210, 174, 144, 803]. Population-Based [803]. Populations [123, 200]. Porteous [593]. Possible [11, 148]. Post [412]. Post-Normal [412]. Potentially [779]. Poverty [123, 757]. Power [868, 33, 346]. Practical [869, 33, 304, 94, 794, 694, 447, 192]. Practice [860, 348, 687, 338]. Practitioner [306]. Pradesh [255]. Precipitation [777]. Predict [700]. Predicting [111, 632]. Prediction [846, 869, 680, 828, 658, 496]. Predictions [765]. Predictor [19]. Predictors [2, 115]. Preference [509]. Prehistory [363]. Preliminary [594]. Premarital [785]. Premier [551]. Prescott [555]. Presence [79]. Present [569]. Presenting [655]. Presser [380]. Prevalence [719, 451, 808, 748, 504]. Prevention [805]. Previous [505]. Price [537, 338, 658]. Prices [141]. Primary [67, 726]. Primer [316, 816, 813, 585].

Principles [588, 690]. Prior [414]. Priorities [705]. Prioritization [743]. Priors [779]. Prize [850]. Probabilistic [697, 812, 657, 525, 588]. Probabilities [675, 144, 654]. Probability [181, 850, 479, 794, 549, 71, 531, 432, 684, 582, 343, 652]. Probationers [782]. Problem [415, 613]. Problems [32, 126, 575, 794, 334]. Procedure [257]. Procedures [474, 605, 68, 755]. Process [377, 608, 477, 109]. Processes [466, 202, 531, 144, 531, 652, 300, 311]. Procrustes [334]. Producing [419]. Production [109]. Productivity [499, 138]. Products [826, 632]. Professionalization [384]. Professionals [316]. Professor [751]. Prognostic [2, 115]. Programme [93, 578]. Programmes [94]. Programming [835]. Programs [223]. progress [75]. Progressively [174]. Project [511, 205]. Projections [777]. promise [75]. Propensity [476]. Properties [32]. Property [446]. Proportion [719]. Proportional [285]. Proportions [548]. Proposal [655]. Proposals [76, 160]. Prospective [46, 236, 134, 676]. Prostate [136]. Prostitute [504]. Protective [388]. Proteomics [647]. Protocols [550]. Providers [627]. Psychological [358, 420]. Psychology [715]. Public [767, 226, 703, 643, 698, 494, 291, 151, 62, 325, 134, 448, 120]. Publication [655, 529]. Pulford [689]. Punishment [11]. Pupil [631, 611]. Pupils [822]. Purchasing [33]. Putting [795].

Qualified [162]. Qualitative [726]. Quality [871, 88, 106, 284, 162, 253, 808, 550, 345, 320, 461]. Quantifying [671, 381]. Quantile [704, 559]. Quantitative [194]. Quarter [287]. Question [356]. Questionnaires [380]. Questions [725]. quis [72]. Quitting [81, 168].

R [835, 518, 530, 525, 430, 520, 468, 686, 757, 527, 517, 463, 596, 650, 467, 304, 528, 521, 533, 867, 491, 559, 690, 516, 709, 555, 585, 763, 793, 654, 692, 459, 837, 841, 529, 561, 710, 711, 855, 693, 754, 756, 587, 765, 592, 428, 839, 586, 558, 591, 859, 861, 522, 567, 835, 563, 837, 370]. Rabe [736]. Rabe-Hesketh [736]. Radiation [125]. Radon [56]. Railways [12]. Rakow [654]. Random [136, 623, 285, 780, 67, 63, 110, 303, 582, 619, 166, 308, 652, 379, 754]. Random-Coefficient [63]. Random-Digit [110]. Random-Effects [136, 285, 780]. Randomization [7, 731, 620]. Randomized [6, 178, 173, 451, 91, 805]. Range [1]. Ranges [80]. Rank [209]. Ranked [361]. Ranking [634]. Rare [427]. Rasch [218]. Rate [389]. Raters [77]. Rates [353, 293, 354]. Rather [510]. Rating [864, 259]. Ratings [160]. Ratio [285]. Ratios [288]. Re [414, 780]. Re-Evaluation [780]. Re-Examining [414]. Reach [91]. Real [584]. Real-life [584]. Realistic [254]. Realities [288]. Reality [42]. Really [551, 141]. Realtime [353, 293, 354]. Recapture [504]. Receiver [759]. Reciprocal [670]. Reclassification [66]. Recognition [375]. Reconstruction [472]. Recorded [629, 11]. Records [76, 635, 29]. Recursive [493]. Reduce [577]. Reduced [256]. Reducing [501]. Referees [551]. Reference [33, 126, 80]. Refusal [474, 801]. Regional [288]. Registration [628]. Regression [93, 54, 704, 503, 137, 19, 304, 673, 621, 77, 855, 197, 827, 559, 734, 595, 441]. Regressions [50]. Regulatory [864, 605]. Reid [710]. Rekab [435]. Related [735, 673, 479, 642, 676]. Relation [43]. Relationship [176, 99, 11]. Relationships [670]. Relative [846]. Releasing [291]. Relevant [150]. Reliability [784, 135, 838]. Reliable [11]. Remembrances [135]. Repair [389]. Reparameterizations [824].

Repeated [254, 255, 9]. Reply [703]. Report [638, 419, 260, 392, 750, 852, 513]. Reported [96]. Reporting [629, 389]. Reports [775]. Reprocessing [125]. Requiring [167]. Research [173, 578, 203, 743, 5, 497, 729, 368]. Researcher [160]. Residence [823, 43]. Residential [56]. Resistant [775]. Resolving [390]. Resources [284, 611]. Respiratory [353, 293, 354]. Respondent [532]. Response [703, 482, 475, 479, 451, 477, 471, 34, 391, 607, 68, 180, 478, 481, 486, 686]. Responses [290]. Responsive [473]. Restrictions [495]. Results [357, 125, 98]. retest [135]. Retrospective [540, 135]. Review [864, 526, 811, 351, 523, 732, 224, 222, 221, 374, 562, 592, 656, 792, 862, 276, 314, 464, 557, 251, 869, 736, 835, 186, 250, 270, 376, 315, 377, 871, 266, 275, 428, 432, 518, 530, 593, 525, 834, 839, 402, 301, 339, 430, 520, 493, 560, 818, 264, 187, 310, 735, 380, 378, 379, 468, 492, 686, 405, 242, 336, 371, 761, 127, 623, 643, 757, 766, 868, 527, 341, 624, 563, 565, 316, 372, 427, 466, 517, 556, 581, 618, 352, 463, 554, 313, 271, 218, 332, 338, 381, 337, 407, 435, 461]. Review [497, 564, 584, 596, 597, 795, 685, 712, 832, 858, 650, 400, 366, 367, 713, 658, 566, 866, 467, 273, 267, 245, 330, 304, 331, 528, 865, 755, 521, 533, 586, 762, 790, 243, 399, 404, 867, 160, 240, 189, 298, 299, 491, 759, 813, 369, 431, 558, 559, 621, 731, 191, 193, 202, 217, 216, 246, 691, 690, 199, 200, 814, 190, 306, 516, 519, 660, 739, 860, 645, 684, 738, 845, 398, 494, 659, 687, 220, 591, 655, 840, 346, 583, 833, 406, 646, 433, 312, 201, 403, 83, 485, 760, 622, 642, 647, 709, 859]. Review [758, 58, 268, 333, 334, 373, 397, 340, 348, 734, 794, 555, 244, 585, 714, 763, 269, 274, 303, 307, 437, 469, 532, 793, 617, 654, 689, 692, 396, 302, 459, 582, 589, 649, 812, 863, 653, 694, 837, 841, 529, 265, 305, 439, 496, 561, 619, 620, 710, 861, 247, 344, 309, 342, 345, 711, 248, 195, 196, 347, 375, 300, 272, 715, 441, 465, 524, 594, 733, 764, 816, 855, 857, 241, 198, 311, 425, 368, 223, 134, 192, 641, 870, 194, 495, 595, 401, 308, 343, 489, 531, 588, 590, 652, 693, 754]. Review [756, 836, 182, 335, 203, 462, 370, 730, 490, 587, 648, 765, 522, 567, 644, 688, 737, 385, 729, 460, 5, 791, 817, 856, 249, 197, 188, 349, 350, 657, 838]. Reviews [683, 708, 728, 753, 789, 185, 215, 239, 263, 297, 329, 365, 395, 458, 424, 440, 488, 515, 553, 580, 616, 640, 682, 707, 727, 752, 788, 810, 831, 854, 429, 426, 219, 434, 598, 651, 436, 438, 815]. Revisited [110]. Ridge [686]. Rigby [757]. Right [356]. Risk [864, 846, 678, 593, 704, 572, 509, 152, 146, 76, 150, 149, 235, 673, 148, 549, 838, 147, 310, 309, 350]. Risk-Adjusted [235]. Risks [17, 22, 155]. Road [634, 288]. Roberts [525]. Robust [603, 480, 212, 287, 558]. Rogerson [49]. Role [161, 670, 318, 81, 168, 151, 505, 570]. Role-Plays [570]. Romano [589]. Ronen [460]. Rosen [433]. Rossi [528]. Rothgeb [380]. Rothstein [529]. Routine [104]. Rubin [737]. Rue [623]. Rules [187]. Ruppert [711]. Ruptured [389].

S [736, 525, 520, 380, 868, 527, 565, 584, 832, 658, 304, 755, 586, 813, 731, 860, 738, 655, 840, 646, 758, 469, 653, 441, 465, 590, 730, 567, 644, 688, 729, 817, 560, 517, 660]. S-PLUS [304, 560, 660, 517]. S.-C [840, 688, 729]. Saddlepoint [756]. Safety [234, 233, 322]. Saito [624]. Saleh [594]. Salmonella [704]. Salomon [853]. Salway [52]. Same [807]. Sample [543, 584, 510, 29, 710, 447, 729]. Samples [76, 80]. Sampling [72, 541, 808, 617, 737, 581]. Sanction [551]. SAS [813, 837, 759, 646, 709, 561]. Satisfaction [540, 420]. Scale [54, 784, 564, 223]. Scales [106, 34]. Scan [46, 676]. Schabenberger [439, 561]. Scheme [72]. Schemes [702, 459]. Schilling [530]. Schlotzhauer [738].

Scholarship [761]. Schölkopf [811]. School [181, 503, 290, 631, 822, 851, 141, 111, 611]. Science [497, 412, 812, 657]. Sciences [649, 733, 587, 351, 732, 218, 306]. Scientific [348]. Scientists [735, 684, 841]. Scores [149]. Scotland [30]. Screening [136, 20, 385, 743]. Search [609, 323, 331]. Secondary [10, 141, 805]. Sector [62, 26]. Sectoral [138]. Security [284, 451]. sed [72]. Seeing [659]. Segers [425]. Segmentation [257]. Segregation [722, 574]. Selected [557]. Selecting [21]. Selection [850, 858, 356, 416, 166, 361]. Selective [47]. Self [96, 197]. Self-learning [197]. Self-Reported [96]. Sellafield [125]. Semi [811]. Semi-Supervised [811]. Semiparametric [704, 863]. Semisupervised [832]. Sen [495]. Seniors [503]. Senn [817]. Sensitivity [54, 254, 476, 320, 348]. Sequence [111, 547]. Sequential [113, 783]. Series [592, 792, 678, 1, 371, 783, 190, 660, 542, 445, 863, 730, 307]. Serious [635]. Service [775, 118, 164]. Services [767, 781, 593, 226, 703, 325, 120, 421]. Session [638, 260, 392, 513]. Set [848]. Setting [784]. Severe [353, 188, 293, 354]. Severini [656]. Sex [148]. Sexuality [203]. Seymour [294]. Shaikh [435]. Shao [729]. Shape [495]. Shared [47]. Sheppard [57]. Shiboski [441]. Shifting [542]. Shifts [114]. Shoe [384]. Short [667]. Should [40]. Sibling [508]. Sick [749]. Sides [745]. Significance [416]. Silvapulle [495]. Simonoff [469]. Simple [824, 55, 651, 77, 476, 504, 650]. Simulation [63, 808, 311, 427, 202]. Simulation-Based [63]. Simultaneous [233, 611]. Singapore [353, 354, 293]. Singer [380]. Single [19]. Singpurwalla [838]. Sir [464, 557, 806, 295, 456]. Sites [634]. Six [801]. Sixth [359]. Size [543, 629, 510, 447, 729]. Skills [799]. Slovenia [604]. Small [572, 213, 633, 174, 710, 575, 282]. Small-Sample [710]. Smith [533]. Smoking [81, 168]. Snooping [636]. Snowball [800]. Social [161, 577, 386, 807, 90, 173, 284, 571, 578, 451, 30, 418, 91, 358, 175, 669]. Society [100]. Socio [749, 673]. Socio-Economic [749, 673]. Software [694]. Solutions [794, 581]. Some [122, 94, 138, 158, 385]. Source [166]. Sources [826]. south [56]. south-west [56]. Space [190, 335, 730, 48]. Spatial [701, 761, 439, 311, 387, 494, 340]. Spatio [648]. Spatio-Temporal [648]. Specifying [208]. Spector [839]. Spiegelhalter [105]. Sporting [112]. Sports [257]. SPSS [689, 837, 651, 692, 650]. Square [807]. Squares [137]. Stage [80]. Stakes [782]. Standard [292]. Standardization [417]. Standards [597, 598]. Staphylococcus [775]. Starter [653]. Starting [81, 168]. Stata [736]. State [508, 538, 190, 632, 730, 335]. Statistic [46]. Statistical [415, 630, 562, 314, 557, 806, 782, 736, 4, 618, 230, 463, 219, 533, 586, 762, 163, 369, 558, 690, 845, 840, 605, 79, 236, 61, 125, 550, 118, 302, 589, 694, 363, 496, 100, 196, 375, 158, 601, 465, 71, 134, 495, 570, 765, 51, 5, 412, 732, 377, 871, 518, 430, 187, 468, 336, 466, 517, 435, 596, 304, 200, 860, 346, 646, 622, 714, 812, 439, 309, 733, 855, 857, 311, 490, 648, 460, 817]. Statistician [133]. Statisticians [318, 554]. Statistics [314, 464, 66, 572, 818, 643, 618, 332, 417, 367, 798, 528, 698, 867, 578, 813, 202, 691, 519, 684, 738, 659, 655, 312, 763, 861, 549, 602, 192, 531, 26, 573, 526, 492, 758, 710, 428, 381, 584, 491, 516, 305, 345, 347, 425, 368]. Statistique [645]. Status [122, 749, 724]. Stay [18, 159, 283]. Steel [57]. Stefanski [711]. Stein [594, 44]. Stein-Type [594]. Steinijans [760]. Stenger [617]. Stereology [554]. Steward [44]. Stewardson [860]. Steyerberg [869]. Still [125]. Still-Births [125]. Stirzaker [531]. Stochastic [466, 202, 373, 531, 590, 836, 8, 493, 566, 531]. Stockwell [765]. Stone [703, 119]. Stoop [532]. Story [745]. Strategies [7, 868, 433]. Strategy [27, 89, 699, 653]. Stratified [80]. Streakiness [257]. Stroup [561]. Structural [685]. Structure [386, 359, 179, 387]. Structured [373]. Structures [290, 624]. Student [10]. Students [162, 71, 223]. Studies [604, 53, 673, 501, 79, 760, 826, 445, 126, 710, 601, 39, 51, 764, 221, 369, 342]. Study [415, 353, 122, 178, 576, 475, 289, 141, 477, 705, 808, 547, 291, 849, 68, 140, 114, 743, 293, 354]. Subject [443, 388, 7, 548]. Subjective [667]. Subjects [162]. Subsample [674]. Subsequent [823]. Substructure [23]. Success [162]. Successful [111]. Sudjianto [521]. Suicide [78]. Sullivan [496]. Sum [702]. Supervised [811]. Supply [166]. Support [527, 284, 358, 649]. Surfaces [686]. Surgery [164]. Surgical [702, 104]. Surrogate [390, 6, 178, 482, 79]. Surveillance [719, 46, 494, 748, 134, 676, 229]. Survey [209, 96, 380, 497, 89, 473, 284, 502, 725, 451, 67, 253, 391, 418, 507, 180, 361, 481, 357, 10, 359, 29, 479, 213, 607, 720, 617]. Surveying [800]. Surveys [603, 474, 604, 541, 254, 801, 473, 800, 584]. Survival [824, 367, 285, 289, 342, 803]. Sutton [529]. Symmetries [588]. Symptoms [726]. Syndrome [353, 773, 293, 354]. Syndromic [494]. Synoptic [415, 613]. Synthesis [413, 719, 748, 774]. Synthetic [291]. System [606, 124, 61, 564]. Systems [781, 566, 754, 648, 373].

T [864, 656, 380, 624, 463, 461, 713, 521, 586, 860, 642, 763, 654, 694, 856]. Tables [55, 807, 848, 675, 851, 227]. Taiwan [353, 354, 293]. Take [284]. Take-up [284].

Takezawa [595]. Taking [206]. Tanner [433]. Tapadar [593]. Targets [781]. Taroni [657]. Tavaré [520]. Taxane [165]. Taxes [81, 168]. Taylor [658]. Teaching [602]. Teams [259]. Technical [223]. Techniques [642, 320, 385]. Teeth [3, 13]. Tell [725, 420]. Temperature [777]. Temperatures [542]. Temporal [648]. Term [474, 477, 539, 283]. Terms [219]. Terrorist [629]. Terry [22]. Test [259, 732, 512, 594, 733, 135]. Testing [93, 380, 7, 363, 589]. Tests [431, 496, 731]. Teugels [425]. Text [197]. Th [733]. Their [386, 177]. Theoretic [608, 21]. Theoretical [33]. Théorique [645]. Theory [732, 623, 338, 578, 691, 687, 312, 617, 733, 425, 194, 754, 335, 188, 656, 341, 594]. There [548]. Therefore [844]. things [135]. Third [454]. Third-Level [454]. Thomas [829, 757, 766]. Thomson [806]. Thumb [187]. Ties [804]. Tillé [581]. Time [592, 65, 678, 824, 371, 680, 46, 660, 327, 542, 307, 632, 445, 863, 69, 730, 283, 48, 466, 792, 190]. Time-to-Event [327, 69]. Times [164]. Tinsley [465]. Tobacco [150, 81, 168, 151]. Tobias [709]. Todman [715]. Tools [518, 473, 304, 320, 299, 194]. Top [630, 321]. Toronto [353, 354, 293]. Total [497]. Track [491]. Tracking [201]. Traffic [234, 233]. Train [12]. Trajectories [637, 547]. Trajectory [669]. Tramontano [739]. Transactions [100]. Transformation [449, 2, 115]. Transformed [209]. Transition [65, 675]. Transitions [111]. Transnational [629]. Transparency [776]. Transport [154]. Treat [32]. Treatment [545, 460]. Treatments [21]. Tree [222]. Trends [386, 137, 542, 778, 548, 803]. trial [183]. Trials [113, 6, 315, 178, 565, 4, 173, 342, 805, 583, 714, 644, 688]. Tribute [712]. Triple [415, 613]. Triple-Link [415, 613]. Trust [847, 698]. Trusts [775]. Tsay [516]. Tsybakov [870]. Tumours [726]. Turner [500]. Tutorials [369]. Two [72, 122, 417, 745, 47, 809, 542, 80]. Two-Phase [72]. Two-Stage [80]. Type [509, 594].

Ubiquitous [286]. Ugly [281]. UK [181, 321, 850, 781, 66, 799, 256, 230, 801, 89, 775, 28, 500, 163, 675, 213, 110, 607, 75, 71, 602, 448, 575, 26, 97, 573]. Umpires [512]. Uncertain [654]. Uncertainty [23, 802, 671, 320, 188]. Under-Age [746]. Under-Recorded [11]. Under-Reporting [629]. Underlying [28]. Understanding [575]. Uneducated [749]. Unemployed [749, 323]. Unemployment [28, 540, 70, 481]. Unified [619]. Union [506, 749]. Unionization [292]. Unique [548]. Unit [572]. Units [236]. Univariate [255, 692]. Universities [602]. University [65, 231, 71]. Unobserved [508, 335]. Unsafe [148]. Unsuccessful [111]. Unusual [627]. Updated [689]. Updating [869]. Uptake [743]. Urban [386]. Urn [866]. USA [721, 604, 746, 387]. Use [82, 6, 96, 824, 503, 165, 255, 291, 676]. Users [837]. Using [234, 233, 846, 736, 428, 704, 136, 818, 178, 742, 576, 543, 482, 10, 724, 802, 479, 813, 46, 738, 851, 162, 646, 179, 327, 164, 111, 826, 34, 793, 637, 547, 694, 2, 115, 104, 570, 779, 478, 803, 586]. Utility [846, 725]. Uttar [255].

V [687, 760, 441, 652, 648]. Valero [659]. Valero-Mora [659]. Validation [869, 178, 451, 808]. Validity [223]. Value [69, 743]. Values [67]. Vancouver [504]. Vannucci [647]. Variable [19, 34, 166]. Variables [106, 582, 308]. Variance [362, 548, 189]. Variate [303]. Variations [701]. Varying [289]. Varying-Coefficient [289]. Vedel [554]. Verbeke [862]. Verification [306]. Versions [285, 689]. versus [508, 137, 288]. via [619]. Viability [110]. Victimization [176, 153]. Victims [176]. View [416, 206]. vigilance [783].

Vining [621]. Viral [322]. Virus [719, 748, 148, 805]. Visual [659]. Vittinghoff [441]. vol [790]. Volatility [658]. Vols [557]. Volume [36, 85, 130, 170, 278, 410, 409, 442, 535, 662, 771, 875, 367, 369, 202, 220, 35, 84, 129, 169, 279, 277, 408, 534, 663, 661, 768, 872, 312]. Voluntary [26]. Voting [755, 9].

W [869, 518, 430, 735, 564, 467, 867, 659, 642, 437, 793, 561, 756]. Wage [723, 256, 722]. Wages [292]. Waiting [164]. Wakefield [228, 44, 52]. Wales [671, 628, 20]. Wang [840, 660, 729]. Ware [556]. Warfare [42]. Wartenberg [44]. Was [42]. Water [826]. Waterman [520]. Watt [763]. Wave [359]. Ways [208]. Weak [287]. Weale [145]. Wealth [757]. Wealthy [844]. Weatherup [818]. Webster [841]. Week [420]. Weight [826]. Weighted [479]. Weighting [34]. Weisberg [497]. Welfare [91]. Welfare-to-Work [91]. Well [610, 844, 358, 420]. Well-Being [610, 358, 420]. Well-Educated [844]. Wellbeing [667]. Welsh [694, 43, 712]. Wermuth [464]. Weron [518]. West [694, 422, 56]. Wheeler [757]. Which [162, 356]. White [70]. Who [612, 598, 597]. Wicket [512]. Wilkinson [566]. Willan [596]. William [681, 455, 614]. Williams [49]. Willinger [564]. Wilson [145]. Windows [689]. Winkel [871]. Winners [700]. Winnings [850]. Withdrawal [71]. Within [849, 782, 781, 541, 7]. Within-Study [849]. within-Subject [7]. Wolfinger [561]. Women [670, 66, 122, 504, 175]. Wood [567]. Wooff [691]. Work [65, 66, 742, 323, 745, 292, 111, 637, 91]. Workbook [643]. Workers [723, 125]. Working [360]. World [381, 597, 844].

X [590].

Yadohisa [624]. Year [679]. Year-of-Birth [679]. Years [133, 326]. Yin [493]. York [114]. Young [177, 533, 659].

Z [460]. Zero [3, 13]. Zero-Inflated [3, 13]. Zhang [871]. Zien [811]. Zivot [660].

References

Byers:1997:MPP

[1] David Byers, James Davidson, and David Peel. Modelling political popularity: An analysis of long-range dependence in opinion poll series. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3): 471–490, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983265. See cor-

rection [101].

Sauerbrei:1999:BMP

[2] W. Sauerbrei and P. Royston. Building multivariable prognostic and diagnostic models: Transformation of the predictors by using fractional polynomials. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(1):71–94, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680468. See corri-

gendum [115].

Bohning:1999:ZIP

[3] Dankmar Bohning, Ekkehart Dietz, Peter Schlattmann, Lisette Mendonca, and Ursula Kirchner. The zero-inflated Poisson model and the decayed, missing and filled teeth index in dental epidemiology. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(2):195–209, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680578. See corri-

gendum [13].

Day:2000:ESG

[4] Simon D. Day and David J. Talbot. Editorial: Statistical guidelines for clinical trials. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(1):1–3, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680503.

Williamson:2000:SRR

[5] P. Williamson, J. L. Hutton, J. Bliss, J. Blunt, M. J. Campbell, and R. Nicholson. Statistical review by research ethics committees. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(1):5–13, ???? 2000. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2680504.

Begg:2000:USE

[6] Colin B. Begg and Denis H. Y. Leung. On the use of surrogate end points in randomized trials. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(1): 15–28, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680505.

Lavori:2000:DTC

[7] Philip W. Lavori and Ree Dawson. A design for testing clinical strategies: Biased adaptive within-subject randomization. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(1):29–38, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680506.

Taylor:2000:SMG

[8] G. J. Taylor, S. I. McClean, and P. H. Millard. Stochastic models of geriatric patient bed occupancy behaviour. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163 (1):39–48, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680507.

Yang:2000:MMR

[9] Min Yang, Harvey Goldstein, and Anthony Heath. Multilevel models for repeated binary outcomes: Attitudes and voting over the electoral cycle. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163 (1):49–62, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680508.

Egerton:2000:MCS

[10] Muriel Egerton. Monitoring contemporary student flows and characteristics: Secondary analyses using the Labour Force Survey and the General Household Survey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(1):63–80, ???? 2000. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2680509.

Pudney:2000:RBC

[11] Stephen Pudney, Derek Deadman, and David Pyle. The relationship between crime, punishment and economic conditions: Is reliable inference possible when crimes are under-recorded? Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(1): 81–97, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680510.

Evans:2000:FTA

[12] Andrew W. Evans. Fatal train accidents on Britain’s mainline railways. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163 (1):99–119, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680511.

Bohning:2000:CZI

[13] D. Böhning, E. Dietz, P. Schlattmann, L. Mendonça, and U. Kirchner. Corrigendum: The zero-inflated Poisson model and the decayed, missing and filled teeth index in dental epidemiology. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(1):121–122, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680512. See [3].

Anonymous:2000:BMa

[14] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(1):??, ???? 2000. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2680513.

Anonymous:2000:FMa

[15] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(1):??, ???? 2000. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2680502.

Murphy:2000:ECB

[16] Mike Murphy. Editorial: Cohabitation in Britain. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(2):123–126, ???? 2000. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2680493.

Berrington:2000:MCC

[17] Ann Berrington and Ian Diamond. Marriage or cohabitation: A competing risks analysis of first-partnership formation among the 1958 British birth cohort. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163 (2):127–151, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680494.

Ermisch:2000:CGB

[18] John Ermisch and Marco Francesconi. Cohabitation in Great Britain: Not for long, but here to stay. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(2): 153–171, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680495.

Frost:2000:CRD

[19] Chris Frost and Simon G. Thompson. Correcting for regression dilution bias: Comparison of methods for a single predictor variable. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(2):173–189, ???? 2000. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2680496.

Sasieni:2000:ACC

[20] Peter D. Sasieni and Joanna Adams. Analysis of cervical cancer mortality and incidence data from England and Wales: Evidence of a beneficial effect of screening. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(2):191–209, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680497.

Claxton:2000:STD

[21] Karl Claxton, Larry F. Lacey, and Stephen G. Walker. Selecting treatments: A decision theoretic approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163 (2):211–225, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680498.

Li:2000:EDC

[22] Lei Li and Karl Kim. Estimating driver crash risks based on the extended Bradley–Terry model: An induced exposure method. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(2):227–240, ????

2000. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2680499.

Fung:2000:IFD

[23] Wing K. Fung and Yue-Qing Hu. Interpreting forensic DNA mixtures: Allowing for uncertainty in population substructure and dependence. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(2): 241–254, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680500.

Anonymous:2000:BMb

[24] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(2):??, ???? 2000. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2680501.

Anonymous:2000:FMb

[25] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(2):??, ???? 2000. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2680492.

Tarling:2000:ESV

[26] Roger Tarling. Editorial: Statistics on the voluntary sector in the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(3): 255–261, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680516.

Dominici:2000:CEA

[27] Francesca Dominici, Jonathan M. Samet, and Scott L. Zeger. Combining evidence on air pollution and daily mortality from the 20 largest US cities: A hierarchical modelling strategy. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(3): 263–302, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680517.

Harvey:2000:EUC

[28] Andrew Harvey and Chia-Hui Chung. Estimating the underlying change in unemployment in the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(3): 303–339, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680518.

Heath:2000:EDL

[29] Anthony Heath, Dorren McMahon, and Jane Roberts. Ethnic differences in the labour market: A comparison of the sample of anonymized records and Labour Force Survey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(3): 341–361, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680519.

Paterson:2000:SCC

[30] Lindsay Paterson. The social class of Catholics in Scotland. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(3):

363–379, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680520.

Longford:2000:HMD

[31] Nicholas T. Longford, Margaret Ely, Rebecca Hardy, and Michael E. J. Wadsworth. Handling missing data in diaries of alcohol consumption. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(3): 381–402, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680521.

Hutton:2000:NNT

[32] J. L. Hutton. Number needed to treat: Properties and problems. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(3): 403–419, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680522.

Cuthbert:2000:TPI

[33] James R. Cuthbert. Theoretical and practical issues in purchasing power parities illustrated with reference to the 1993 organization for economic cooperation and development data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(3): 421–444, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680523.

Moustaki:2000:WIN

[34] Irini Moustaki and Martin Knott. Weighting for item non-response in attitude scales by using latent variable models with covariates. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(3):445–459, ???? 2000. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2680524.

Anonymous:2000:IAV

[35] Anonymous. Index of authors, volume 163, 2000. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(3):461–462, ???? 2000. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2000:VI

[36] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(3): 461–464, ???? 2000. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680514.

Anonymous:2000:BMc

[37] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(3):??, ???? 2000. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2680525.

Anonymous:2000:FMc

[38] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(3):??, ???? 2000. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2680515.

Wakefield:2001:EDC

[39] Jon Wakefield, Mike Quinn, and Gillian Raab. Editorial: Disease clusters and ecological studies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(1):1–2, ???? 2001. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2680527.

Elliott:2001:DCS

[40] Paul Elliott and Jon Wakefield. Disease clusters: Should they be investigated, and, if so, when and how? Journal of the Royal Statistical Society. Series A (Statistics in Society), 164 (1):3–12, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680528. See com-

ments [44].

Wartenberg:2001:IDC

[41] Daniel Wartenberg. Investigating disease clusters: Why, when and how? Journal of the Royal Statistical Society. Series A (Statistics in Society), 164 (1):13–22, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680529. See com-

ments [44].

Stein:2001:CNW

[42] C. E. Stein, S. Bennett, S. Crook, and F. Maddison. The cluster that never was: Germ warfare experiments and health authority reality in Dorset. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164 (1):23–27, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680530. See com-

ments [44].

Steward:2001:EII

[43] John Steward and Gareth John. An ecological investigation of the incidence of cancer in Welsh children for the period 1985–1994 in relation to residence near the coastline. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(1):29–43, ???? 2001. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2680531. See

comments [44].

Dolk:2001:CPE

[44] Helen Dolk. Comments on the papers by Elliott and Wakefield, Wartenberg, Stein et al. and Steward and John. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164 (1):45–47, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680532. See [40,

41, 42, 43].

Williams:2001:MEA

[45] Joan R. Williams, Freda E. Alexander, Ray A. Cartwright, and Richard J. Q. McNally. Methods for eliciting aetiological clues from geographically clustered cases of disease, with application to leukaemia–lymphoma data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164 (1):49–60, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680533. See com-

ments [49].

Kulldorff:2001:PTP

[46] Martin Kulldorff. Prospective time periodic geographical disease surveillance using a scan statistic. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(1): 61–72, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680534. See com-

ments [49].

Knorr-Held:2001:SCM

[47] Leonhard Knorr-Held and Nicola G. Best. A shared component model for detecting joint and selective clustering of two diseases. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(1):73–85, ???? 2001. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2680535. See

comments [49].

Rogerson:2001:MPP

[48] Peter A. Rogerson. Monitoring point patterns for the development of space– time clusters. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(1):87–96, ???? 2001. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2680536. See

comments [49].

Lawson:2001:CPW

[49] Andrew Lawson. Comments on the papers by Williams et al., Kulldorff, Knorr-Held and Best, and Rogerson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164 (1):97–99, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680537. See [45,

46, 47, 48].

Gelman:2001:MAM

[50] Andrew Gelman, David K. Park, Stephen Ansolabehere, Phillip N. Price, and Lorraine C. Minnite. Models, assumptions and model checking in ecological regressions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(1):101–118, ???? 2001. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/

2680538. See comments [52].

Wakefield:2001:SFE

[51] Jonathan Wakefield and Ruth Salway. A statistical framework for ecological and aggregate studies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(1):119–137, ???? 2001. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/

2680539. See comments [52].

Sheppard:2001:CPG

[52] Lianne Sheppard. Comments on the papers by Gelman et al. and Wakefield and Salway. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(1):139, ???? 2001. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2680540. See

[50, 51].

Guthrie:2001:OBM

[53] Katherine A. Guthrie and Lianne Sheppard. Overcoming biases and misconceptions in ecological studies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(1): 141–154, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680541. See com-

ments [57].

Best:2001:ERA

[54] Nicky Best, Samantha Cockings, James Bennett, Jon Wakefield, and Paul Elliott. Ecological regression analysis of environmental benzene exposure and childhood leukaemia: Sensitivity to data inaccuracies, geographical scale and ecological bias. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(1):155–174, ???? 2001. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2680542. See

comments [57].

Chambers:2001:SME

[55] R. L. Chambers and D. G. Steel. Simple methods for ecological inference in 2×2 tables. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(1):175–192, ???? 2001. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2680543. See

comments [57].

Darby:2001:PAI

[56] Sarah Darby, Harz Deo, Richard Doll, and Elise Whitley. A parallel analysis of individual and ecological data on residential radon and lung cancer in southwest England. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(1):193–203, ????

2001. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/

2680544. See comments [57].

Armstrong:2001:CPG

[57] Ben Armstrong. Comments on the papers by Guthrie and Sheppard, Best et al., Chambers and Steel and Darby et al. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(1):205–207, ???? 2001. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2680545. See

[53, 54, 55, 56].

Lumley:2001:BR

[58] Thomas Lumley, Margaret S. Pepe, Patrick J. Heagerty, Robert Crouchley, and Richard Davies. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(1): 209–212, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680546.

Anonymous:2001:BMa

[59] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(1):??, ???? 2001. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2680547.

Anonymous:2001:FMa

[60] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(1):??, ???? 2001. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2680526.

McKeever:2001:EGD

[61] Lucy McKeever and Mike Cuthbertson. Editorial: The growing demands on the European Statistical System. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(2): 213–215, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680549.

Smith:2001:CFP

[62] Peter C. Smith, Nigel Rice, and Roy Carr-Hill. Capitation funding in the public sector. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(2):217–257, ???? 2001. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2680550.

Longford:2001:SBD

[63] Nicholas T. Longford. Simulation-based diagnostics in random-coefficient models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164 (2):259–273, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680551.

Dellaportas:2001:BAM

[64] Petros Dellaportas, Adrian F. M. Smith, and Photis Stavropoulos. Bayesian analysis of mortality data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(2): 275–291, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680552.

Biggeri:2001:TUW

[65] L. Biggeri, M. Bini, and L. Grilli. The transition from university to work: A multilevel approach to the analysis of the time to obtain the first job. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(2): 293–305, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680553.

Blackwell:2001:WWU

[66] Louisa Blackwell. Women’s work in UK official statistics and the 1980 reclassification of occupations. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(2): 307–325, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680554.

Lindsey:2001:DCN

[67] J. K. Lindsey and P. J. Lindsey. Detecting covariates with non-random missing values in a survey of primary education in Madagascar. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(2):327–338, ???? 2001. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2680555.

Rodriguez:2001:IEP

[68] Germán Rodríguez and Noreen Goldman. Improved estimation procedures for multilevel models with binary response: A case-study. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(2): 339–355, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680556.

Tudur:2001:VAD

[69] Catrin Tudur, Paula R. Williamson, Saboor Khan, and Lesley Y. Best. The value of the aggregate data approach in meta-analysis with time-to-event outcomes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(2):357–370, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680557.

Leslie:2001:DFU

[70] Derek Leslie, Joanne Lindley, and Leighton Thomas. Decline and fall: Unemployment among Britain’s non-white ethnic communities 1960–1999. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(2): 371–387, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680558.

Smith:2001:DUS

[71] Jeremy P. Smith and Robin A. Naylor. Dropping out of university: A statistical analysis of the probability of withdrawal for UK university students. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164 (2):389–405, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680559.

Barnett:2001:TPS

[72] V. Barnett, J. Haworth, and T. M. F. Smith. A two-phase sampling scheme with applications to auditing or sed quis custodiet ipsos custodies? Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(2): 407–422, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680560.

Anonymous:2001:BMb

[73] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(2):??, ???? 2001. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2680561.

Anonymous:2001:FMb

[74] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(2):??, ???? 2001. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2680548.

Simpson:2001:EUC

[75] Ludi Simpson. Editorial: UK census 2001 — progress and promise. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(3): 423–425, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680564.

Dale:2001:PSA

[76] Angela Dale and Mark Elliot. Proposals for 2001 samples of anonymized records: An assessment of disclosure risk. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(3):427–447, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680565.

Lipsitz:2001:SME

[77] Stuart R. Lipsitz, John Williamson, Neil Klar, Joseph Ibrahim, and Michael Parzen. A simple method for estimating a regression model for κ between a pair of raters. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(3):449–465, ???? 2001. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2680566.

Lloyd:2001:CSP

[78] Chris J. Lloyd and Paul S. F. Yip. A comparison of suicide patterns in Australia and Hong Kong. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(3): 467–483, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680567.

Leung:2001:SMC

[79] Denis Heng-Yan Leung. Statistical methods for clinical studies in the presence of surrogate end points. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(3): 485–503, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680568.

Thompson:2001:ERR

[80] Mary Lou Thompson, Steven D. Edland, Laura E. Gibbons, and Susan M. McCurry. Estimating reference ranges from stratified two-stage samples. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(3): 505–516, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680569.

Forster:2001:RTT

[81] Martin Forster and Andrew M. Jones. The role of tobacco taxes in starting and quitting smoking: Duration analysis of British data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(3):517–547, ???? 2001. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2680570. See

corrigendum [168].

Ali:2001:AID

[82] Mohamed M. Ali, Tom Marshall, and Abdel G. Babiker. Analysis of incomplete durations with application to contraceptive use. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(3):549–563, ???? 2001. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2680571.

Longford:2001:BR

[83] Nicholas T. Longford, David Firth, Michael Noble, and George Smith. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(3):565–566, ???? 2001. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2680572.

Anonymous:2001:IIA

[84] Anonymous. Index: Index of authors, volume 164, 2001. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(3):567–569, ???? 2001. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2001:VI

[85] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(3): 567–569, ???? 2001. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680562.

Anonymous:2001:BMc

[86] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(3):??, ???? 2001. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2680573.

Anonymous:2001:FMc

[87] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 164(3):??, ???? 2001. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2680563.

Butcher:2002:EHQ

[88] Bob Butcher. Editorial: High quality evaluation in Britain. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(1):

1–3, February 2002. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559755.

Gershuny:2002:ECO

[89] J. Gershuny. Editorial: Co-ordinating longitudinal survey data in the UK: Towards a national strategy. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(1): 3–7, February 2002. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559756.

Dolton:2002:EES

[90] Peter Dolton. Evaluation of economic and social policies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(1):9–11, February 2002. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559757.

Riccio:2002:ERR

[91] James A. Riccio and Howard S. Bloom. Extending the reach of randomized social experiments: New directions in evaluations of American welfare-to-work and employment initiatives. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(1): 13–30, February 2002. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559758.

Plewis:2002:MIH

[92] Ian Plewis. Modelling impact heterogeneity. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(1):31–38, February 2002. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559759.

Battistin:2002:TPE

[93] Erich Battistin and Enrico Rettore. Testing for programme effects in a regression discontinuity design with imperfect compliance. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(1):39–57, February 2002. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559760.

Lechner:2002:SPI

[94] Michael Lechner. Some practical issues in the evaluation of heterogeneous labour market programmes by matching methods. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(1):59–82, February 2002. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559761.

White:2002:ECG

[95] Michael White and John Killeen. The effect of careers guidance for employed adults on continuing education: Assessing the importance of attitudinal information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(1):83–95, February 2002. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559762.

Biemer:2002:MEE

[96] Paul P. Biemer and Christopher Wiesen. Measurement error evaluation of selfreported drug use: A latent class analysis of the US national household survey on drug abuse. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(1):97–119, February 2002. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559763.

Tilley:2002:PGP

[97] James Tilley. Political generations and partisanship in the UK, 1964–1997. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(1): 121–135, February 2002. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559764.

Yang:2002:MMA

[98] Min Yang, Harvey Goldstein, William Browne, and Geoffrey Woodhouse. Multivariate multilevel analyses of examination results. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(1):137–153, February 2002. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559765.

Forcheh:2002:ELL

[99] N. Forcheh. Ehrenberg law-like relationship and anthropometry. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(1): 155–172, February 2002. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559766.

Rosenbaum:2002:TSS

[100] Sidney Rosenbaum. Transactions of the Statistical Society of London (1837). Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(1): 173–185, February 2002. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559767.

Byers:2002:MPP

[101] David Byers, James Davidson, and David Peel. Modelling political popularity: A correction. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(1):187–189, February 2002. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/

3559768. See [1].

Anonymous:2002:BMa

[102] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(1):??, ???? 2002. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559769.

Anonymous:2002:FMa

[103] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(1):??, ???? 2002. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559754.

Spiegelhalter:2002:CAS

[104] David J. Spiegelhalter, Paul Aylin, Nicola G. Best, Stephen J. W. Evans, and Gordon D. Murray. Commissioned analysis of surgical performance using routine data: Lessons from the Bristol Inquiry. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(2):191–221, June 2002. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559925. See

discussion [105].

Curnow:2002:DPS

[105] R. N. Curnow. Discussion on the paper by Spiegelhalter, Aylin, Best, Evans and Murray. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(2):221–231, June 2002. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [104].

Fayers:2002:CVI

[106] Peter M. Fayers and David J. Hand. Causal variables, indicator variables and measurement scales: An example from quality of life. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(2):233–253, June 2002. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/

3559926. See discussion [107].

Bartholomew:2002:DPF

[107] D. J. Bartholomew. Discussion on the paper by Fayers and Hand. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(2): 253–261, June 2002. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [106].

Lancaster:2002:DIH

[108] Gillian Lancaster and Mick Green. Deprivation, ill-health and the ecological fallacy. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(2):263–278, June 2002. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559927.

Patterson:2002:MDM

[109] K. D. Patterson. Modelling the data measurement process for the index of production. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(2):279–296, June 2002. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559928.

Nicolaas:2002:RDD

[110] Gerry Nicolaas and Peter Lynn.

Random-digit dialling in the UK: Viability revisited. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(2):297–316, June 2002. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559929.

McVicar:2002:PSU

[111] Duncan McVicar and Michael Anyadike-Danes. Predicting successful and unsuccessful transitions from school to work by using sequence methods. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(2): 317–334, June 2002. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559930.

Farrell:2002:IED

[112] Lisa Farrell and Michael A. Shields.

Investigating the economic and demographic determinants of sporting participation in England. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(2):335–348, June 2002. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559931.

Atkinson:2002:CDS

[113] Anthony C. Atkinson. The comparison of designs for sequential clinical trials with covariate information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(2): 349–373, June 2002. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559932.

Talih:2002:END

[114] Makram Talih and Ronald D. Fricker, Jr. Effects of neighbourhood demographic shifts on findings of environmental injustice: A New York City casestudy. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(2):375–397, June 2002. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559933.

Sauerbrei:2002:CBM

[115] W. Sauerbrei and P. Royston. Corrigendum: Building multivariable prognostic and diagnostic models: Transformation of the predictors by using fractional polynomials. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(2):399–400, June 2002. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/

3559934. See [2].

Anonymous:2002:BMb

[116] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(2):??, ???? 2002. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559935.

Anonymous:2002:FMb

[117] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(2):??, ???? 2002. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559924.

Phillpotts:2002:EIS

[118] Greg Phillpotts. Editorial: An independent statistical service? Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(3): 401–403, October 2002. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559695.

Smith:2002:DPS

[119] Peter C. Smith. Discussion on the paper by Stone. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(3):405–434, October 2002. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [120].

Stone:2002:HME

[120] M. Stone. How not to measure the efficiency of public services (and how one might). Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(3):405–434, October 2002. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559696. See

discussion [119], corrigendum [767], and response [703].

Bell:2002:CNC

[121] M. Bell, M. Blake, P. Boyle, O. Duke-Williams, P. Rees, J. Stillwell, and G. Hugo. Cross-national comparison of internal migration: Issues and measures. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(3): 435–464, October 2002. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559697.

Boyle:2002:CNS

[122] Paul Boyle, Thomas J. Cooke, Keith Halfacree, and Darren Smith. A crossnational study of the effects of family migration on women’s labour market status: Some difficulties with integrating microdata from two censuses. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(3): 465–480, October 2002. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559698.

Denny:2002:NMC

[123] Kevin Denny. New methods for comparing literacy across populations: Insights from the measurement of poverty. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(3): 481–493, October 2002. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559699.

Steele:2002:CDI

[124] Fiona Steele, James Brown, and Ray Chambers. A controlled donor imputation system for a one-number census. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(3): 495–522, October 2002. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559700.

Pearce:2002:SBA

[125] Mark S. Pearce, Heather O. Dickinson, Murray Aitkin, and Louise Parker. Stillbirths among the offspring of male radiation workers at the Sellafield Nuclear Reprocessing Plant: Detailed results and statistical aspects. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(3): 523–548, October 2002. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559701.

Richardson:2002:MMM

[126] Sylvia Richardson, Laurent Leblond, Isabelle Jaussent, and Peter J. Green. Mixture models in measurement error problems, with reference to epidemiological studies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(3):549–566, October 2002. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559702.

Cohen:2002:BR

[127] Bernard L. Cohen. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(3):

567–568, ???? 2002. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559703.

Cohen:2002:C

[128] Bernard L. Cohen. Correspondence. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(3): 567–568, October 2002. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2002:IAV

[129] Anonymous. Index of authors: volume 165, 2002. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(3):569–571, October 2002. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2002:VI

[130] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(3): 569–571, ???? 2002. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559693.

Anonymous:2002:BMc

[131] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(3):??, ???? 2002. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3559704.

Anonymous:2002:FMc

[132] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(3):??, ????

2002. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559694.

Cook:2003:ENS

[133] Len Cook. Editorial: The national statistician: 2 years on. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(1): 1–4, February 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559824.

Sonesson:2003:RDP

[134] Christian Sonesson and David Bock.

A review and discussion of prospective statistical surveillance in public health. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166 (1):5–21, February 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559825.

Smith:2003:RTP

[135] James P. Smith and Duncan Thomas.

Remembrances of things past: test– retest reliability of retrospective migration histories. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(1):23–49, February 2003. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559826.

Brant:2003:SPC

[136] Larry J. Brant, Shan L. Sheng, Christopher H. Morrell, Geert N. Verbeke, Emmanuel Lesaffre, and H. Ballentine Carter. Screening for prostate cancer by using random-effects models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(1): 51–62, February 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559827.

Dessens:2003:PTO

[137] Jos A. G. Dessens, Wim Jansen, Harry B. G. Ganzeboom, and Peter G. M. Van Der Heijden. Patterns and trends in occupational attainment of first jobs in the Netherlands, 1930–1995: Ordinary least squares regression versus conditional multinomial logistic regression. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166 (1):63–84, February 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559828.

Malley:2003:SNI

[138] Jim Malley, Anton Muscatelli, and Ulrich Woitek. Some new international comparisons of productivity performance at the sectoral level. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(1): 85–104, February 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559829.

Jonker:2003:ELE

[139] M. A. Jonker. Estimation of life expectancy in the middle ages. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(1): 105–117, February 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559830.

Simonite:2003:ELC

[140] V. Simonite and W. J. Browne. Estimation of a large cross-classified multilevel model to study academic achievement in a modular degree course. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(1): 119–133, February 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559831.

Leech:2003:CER

[141] Dennis Leech and Erick Campos. Is comprehensive education really free?: A case-study of the effects of secondary school admissions policies on house prices in one local area. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(1): 135–154, February 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559832.

Anonymous:2003:BMa

[142] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(1):??, ???? 2003. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559833.

Anonymous:2003:FMa

[143] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(1):??, ???? 2003. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559823.

Wilson:2003:IDD

[144] Ian J. Wilson, Michael E. Weale, and David J. Balding. Inferences from DNA data: Population histories, evolutionary processes and forensic match probabilities. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(2):155–188, June 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559656. See dis-

cussion [145].

Anonymous:2003:DPW

[145] Anonymous. Discussion on the paper by Wilson, Weale and Balding. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166 (2):188–201, June 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [144].

Cox:2003:CRI

[146] D. R. Cox and S. C. Darby. The communication of risk: [introduction]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166 (2):203–204, June 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559657.

Smith:2003:FPC

[147] Adrian Smith. Foreword to the papers on ‘The Communication of Risk’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166 (2):205–206, June 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559658.

Richens:2003:HIV

[148] John Richens, John Imrie, and Helen Weiss. Human immunodeficiency virus risk: Is it possible to dissuade people from having unsafe sex? Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(2): 207–215, June 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559659.

Graham:2003:CRC

[149] Ian M. Graham and Elizabeth Clavel.

Communicating risk — coronary risk scores. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(2):217–223, June 2003. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559660.

Darby:2003:TIR

[150] Sarah C. Darby. Tobacco — the importance of relevant information on risk. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166 (2):225–231, June 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559661.

Simpson:2003:TPP

[151] David Simpson and Sue Lee. Tobacco — public perceptions and the role of the industry. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(2):233–239, June 2003. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559662.

Cox:2003:CRH

[152] D. R. Cox. Communication of risk: Health hazards from mobile phones. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166 (2):241–246, June 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559663.

Wiles:2003:CVE

[153] Paul Wiles, Jon Simmons, and Ken Pease. Crime victimization: Its extent and communication. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(2): 247–252, June 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559664.

Evans:2003:AFT

[154] Andrew W. Evans. Accidental fatalities in transport. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(2):253–260, June 2003. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/3559665.

Rosenbaum:2003:CRA

[155] Michael S. Rosenbaum and Martin G. Culshaw. Communicating the risks arising from geohazards. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(2):261–270, June 2003. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3559666.

Anonymous:2003:BMb

[156] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(2):??, ???? 2003. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559667.

Anonymous:2003:FMb

[157] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(2):??, ???? 2003. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559655.

Senn:2003:EPP

[158] Stephen Senn and Nicoletta Rosati. Editorial: Pharmaceuticals, patents and competition — some statistical issues. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(3): 271–274, October 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559742.

Urquhart:2003:EHL

[159] John Urquhart. Editorial: How long do pharmaceuticals stay in the marketplace? Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(3):274–277, October 2003. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559743.

Jayasinghe:2003:MCC

[160] Upali W. Jayasinghe, Herbert W.

Marsh, and Nigel Bond. A multilevel cross-classified modelling approach to peer review of grant proposals: The effects of assessor and researcher attributes on assessor ratings. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(3): 279–300, October 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559744.

Andersen:2003:SIP

[161] Robert Andersen and Anthony Heath. Social identities and political cleavages: The role of political context. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(3): 301–327, October 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559745.

Leslie:2003:USM

[162] Derek Leslie. Using success to measure quality in British higher education: Which subjects attract the bestqualified students? Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(3):329–347, October 2003. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559746.

Holt:2003:NNS

[163] D. Holt. The need for new statistical legislation for the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(3): 349–367, October 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559747.

Martin:2003:UPM

[164] Stephen Martin and Peter C. Smith.

Using panel methods to model waiting times for National Health Service surgery. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(3):369–387, October 2003. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559748.

Cooper:2003:BAM

[165] Nicola J. Cooper, Keith R. Abrams, Alex J. Sutton, David Turner, and Paul C. Lambert. A Bayesian approach to Markov modelling in costeffectiveness analyses: Application to taxane use in advanced breast cancer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(3): 389–405, October 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559749.

Steele:2003:SES

[166] Fiona Steele. Selection effects of source of contraceptive supply in an analysis of discontinuation of contraception: Multilevel modelling when random effects are correlated with an explanatory variable. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(3): 407–423, October 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559750.

Corradi:2003:EDE

[167] Fabio Corradi, Giampietro Lago, and Federico M. Stefanini. The evaluation of DNA evidence in pedigrees requiring population inference. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(3): 425–440, October 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559751.

Forster:2003:CRT

[168] Martin Forster and Andrew M. Jones. Corrigendum: The role of tobacco taxes in starting and quitting smoking: Duration analysis of British data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(3): 441–442, October 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559752. See [81].

Anonymous:2003:IAV

[169] Anonymous. Index of authors, volume 166, 2003. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(3):443–444, October 2003. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2003:VI

[170] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(3): 443–444, ???? 2003. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559740.

Anonymous:2003:BMc

[171] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(3):??, ???? 2003. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559753.

Anonymous:2003:FMc

[172] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 166(3):??, ???? 2003. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559741.

Fitz-Gibbon:2004:ENR

[173] Carol Fitz-Gibbon. Editorial: The need for randomized trials in social research. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167 (1):1–4, February 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559796.

Rees:2004:FPI

[174] Philip Rees, Paul Norman, and Dominic Brown. A framework for progressively improving small area population estimates. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1):5–36, February 2004. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559797.

Sobel:2004:MMM

[175] Michael E. Sobel, Nan Dirk De Graaf, Anthony Heath, and Ying Zou. Men matter more: The social class identity of married British women, 1985–1991. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167 (1):37–52, February 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559798.

Deadman:2004:OVC

[176] Derek Deadman and Ziggy MacDonald. Offenders as victims of crime?: An investigation into the relationship between criminal behaviour and victimization. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1):53–67, February 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559799.

Ermisch:2004:PPJ

[177] John Ermisch, Marco Francesconi, and David J. Pevalin. Parental partnership and joblessness in childhood and their influence on young people’s outcomes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1): 69–101, February 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559800.

Burzykowski:2004:VSE

[178] Tomasz Burzykowski, Geert Molenberghs, and Marc Buyse. The validation of surrogate end points by using data from randomized clinical trials: A case-study in advanced colorectal cancer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167 (1):103–124, February 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559801.

Liao:2004:EHS

[179] Tim Futing Liao. Estimating household structure in Ancient China by using historical data: A latent class analysis of partially missing patterns. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1): 125–139, February 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559802.

Smith:2004:EAP

[180] Philip J. Smith, David C. Hoaglin, J. N. K. Rao, Michael P. Battaglia, and Danni Daniels. Evaluation of adjustments for partial non-response bias in the US national immunization survey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1): 141–156, February 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559803.

Arulampalam:2004:HMP

[181] Wiji Arulampalam, Robin A. Naylor, and Jeremy P. Smith. A hazard model of the probability of medical school dropout in the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1):157–178, February 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559804.

Taves:2004:BR

[182] Donald R. Taves and Anthony C. Atkinson. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1):179–181, ???? 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559805.

Taves:2004:CFA

[183] Donald R. Taves. Correspondence: Faulty assumptions in Atkinson’s criteria for clinical trial design. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1): 179–181, February 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2004:BMa

[184] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1):183, ???? 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3559822.

Anonymous:2004:BRa

[185] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1): 183, February 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Best:2004:BRB

[186] Mark A. Best. Book review: Bioinformatics: the Machine Learning Approach, 2nd edn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1):184, February 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3559806.

Burrell:2004:BRB

[187] Quentin L. Burrell. Book review: Statistical Rules of Thumb. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1):

184–185, February 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559807.

Wright:2004:BRB

[188] Lesley F. Wright. Book review: Information Gap Decision Theory: Decisions under Severe Uncertainty. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1): 185–186, February 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559808.

Jones:2004:BRB

[189] M. C. Jones. Book review: Components of Variance. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1):186–187, February 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559809.

Kumar:2004:BRB

[190] Kuldeep Kumar. Book review: Time Series Analysis by State Space Methods. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1): 187–188, February 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559810.

Kemp:2004:BRBa

[191] Freda Kemp. Book review: A Course in Financial Calculus. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1):188–189, February 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3559811.

Sprent:2004:BRB

[192] P. Sprent. Book review: Modern Medical Statistics: a Practical Guide. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1): 189–190, February 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559812.

Kemp:2004:BRBb

[193] Freda Kemp. Book review: Data Mining. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1):190–191, February 2004. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/3559813.

Sreenivasan:2004:BRB

[194] Ravi Sreenivasan. Book review: Applied Quantitative Finance: Theory and Computational Tools. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1):191–192, February 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3559814.

Sapatinas:2004:BR

[195] Theofanis Sapatinas. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1):192, ???? 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559815.

Sapatinas:2004:BRB

[196] Theofanis Sapatinas. Book review:

The Elements of Statistical Learning. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167 (1):192, February 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Wooff:2004:BRB

[197] David Wooff. Book review: Logistic Regression: a Self-learning Text, 2nd edn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1): 192–194, February 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559816.

Sisson:2004:BRB

[198] Scott Sisson. Book review: Hidden Markov Models for Bioinformatics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1): 194–195, February 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559817.

King:2004:BR

[199] Ruth King. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1):195, ???? 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559818.

King:2004:BRB

[200] Ruth King. Book review: Statistical Analysis of Data from Marked Bird Populations. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1):195, February 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lingaiah:2004:BRB

[201] Dharmendra Lingaiah. Book review: The Estimation and Tracking of Frequency. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1):195–196, February 2004. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559819.

Kemp:2004:BRBc

[202] Freda Kemp. Book review: Handbook of Statistics, Volume 21, Stochastic Processes: Modeling and Simulation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1): 196–197, February 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559820.

Tighe:2004:BRB

[203] James Tighe. Book review: Handbook for Conducting Research on Human Sexuality. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1):197–198, February 2004. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559821.

Anonymous:2004:FMa

[204] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(1):??, ???? 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559795.

Charlton:2004:EEA

[205] John Charlton. Editorial: Evaluating automatic edit and imputation methods, and the EUREDIT project. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(2): 199–207, May 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559858.

Redfern:2004:AVC

[206] Philip Redfern. An alternative view of the 2001 census and future census taking. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(2):209–228, May 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559859. See dis-

cussion [207].

Anonymous:2004:DMC

[207] Anonymous. Discussion on the meeting on ‘The 2001 census and beyond’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167 (2):229–248, May 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [206].

Petrakos:2004:NWS

[208] George Petrakos, Claudio Conversano, Gregory Farmakis, Francesco Mola, Roberta Siciliano, and Photis Stavropoulos. New ways of specifying data edits. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(2):249–274, May 2004. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/3559860.

Beguin:2004:MOD

[209] Cédric Béguin and Beat Hulliger. Multivariate outlier detection in incomplete survey data: The epidemic algorithm and transformed rank correlations. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(2):275–294, May 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559861.

Manzari:2004:CEI

[210] Antonia Manzari. Combining editing and imputation methods: An experimental application on population census data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(2):295–307, May 2004. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/3559862.

DiZio:2004:BNI

[211] Marco Di Zio, Mauro Scanu, Lucia Coppola, Orietta Luzi, and Alessandra Ponti. Bayesian networks for imputation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(2):309–322, May 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559863.

Chambers:2004:RAM

[212] Ray Chambers, Adão Hentges, and Xinqiang Zhao. Robust automatic methods for outlier and error detection. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167 (2):323–339, May 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559864.

Longford:2004:MDS

[213] Nicholas T. Longford. Missing data and small area estimation in the UK Labour Force Survey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(2):341–373, May 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559865.

Anonymous:2004:BMb

[214] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(2):375, ???? 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559873.

Anonymous:2004:BRb

[215] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(2):375, May 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kemp:2004:BRa

[216] Freda Kemp. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(2):376, ???? 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559866.

Kemp:2004:BRBd

[217] Freda Kemp. Book review: Environmental Foresight and Models: a Manifesto. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(2):376, May 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Espejo:2004:BRBa

[218] Mariano Ruiz Espejo. Book review: Applying the Rasch Model: Fundamental Measurement in the Human Sciences. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(2):377, May 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559867;   http:
//www.jstor.org/stable/3559868.

Espejo:2004:BRBb

[219] Mariano Ruiz Espejo. Book reviews: The Oxford Dictionary of Statistical Terms. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(2):377, May 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Larsen:2004:BRB

[220] Pia Veldt Larsen. Book review: Maximum Penalized Likelihood Estimation, Volume I, Density Estimation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(2):378, May 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559869.

Auton:2004:BRB

[221] Tim Auton. Book review: Applied Functional Data Analysis: Methods and Case Studies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(2):378–379, May 2004. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559870.

Arnold:2004:BRB

[222] Gillian M. Arnold. Book review: Experimental Design and Analysis for Tree Improvement. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(2):379, May 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Sofroniou:2004:BRB

[223] Nick Sofroniou. Book review: Largescale Assessment Programs for All Students: Validity, Technical Adequacy, and Implementation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(2):379, May 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559871.

Arnold:2004:BR

[224] Gillian M. Arnold. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(2):380, ???? 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559872.

Anonymous:2004:FMb

[225] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(2):??, ???? 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559857.

Bird:2004:EPM

[226] Sheila M. Bird. Editorial: Performance monitoring in the public services. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3): 381–383, August 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559771.

Wakefield:2004:EIT

[227] Jon Wakefield. Ecological inference for 2 × 2 tables (with discussion). Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3): 385–425, August 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559772. See dis-

cussion [228].

Anonymous:2004:DPW

[228] Anonymous. Discussion on the paper by Wakefield. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3):425–445, August 2004. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [227].

Best:2004:FPP

[229] Nicky Best and Simon Day. Foreword to the papers on ‘Performance Monitoring and Surveillance’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3): 447–448, August 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559773.

Draper:2004:SAP

[230] David Draper and Mark Gittoes. Statistical analysis of performance indicators in UK higher education. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3): 449–474, August 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559774. See com-

ments [232].

Bratti:2004:HEO

[231] Massimiliano Bratti, Abigail McKnight, Robin Naylor, and Jeremy Smith. Higher education outcomes, graduate employment and university performance indicators. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3):475–496, August 2004. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559775. See

comments [232].

Anonymous:2004:HEP

[232] Anonymous. Higher education performance indicators: Invited comments on the papers by Draper and Gittoes and Bratti et al. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3):497–499, August 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/

3559776. See [230, 231].

Bailey:2004:SMM

[233] Trevor C. Bailey and Paul J. Hewson. Simultaneous modelling of multiple traffic safety performance indicators by using a multivariate generalized linear mixed model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3):501–517, August 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559777. See com-

ments [234].

Anonymous:2004:MLG

[234] Anonymous. Monitoring local government traffic safety performance by using multiple indicators: Invited comments on the paper by Bailey and Hewson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3):519–521, August 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559778. See [233].

Grigg:2004:ORA

[235] O. Grigg and V. Farewell. An overview of risk-adjusted charts. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3): 523–539, August 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559779. See com-

ments [237].

Marshall:2004:SIP

[236] Clare Marshall, Nicky Best, Alex Bottle, and Paul Aylin. Statistical issues in the prospective monitoring of health outcomes across multiple units. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3): 541–559, August 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559780. See com-

ments [237].

Anonymous:2004:MCP

[237] Anonymous. Monitoring clinical performance: Invited comments on the papers by Grigg and Farewell and Marshall et al. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3):561–563, August 2004. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559781. See

[235, 236].

Anonymous:2004:BMc

[238] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3):564, ???? 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559794.

Anonymous:2004:BRc

[239] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167 (3):564, August 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Jones:2004:BR

[240] M. C. Jones. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3): 565–566, August 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559782.

Sisson:2004:BR

[241] Scott Sisson. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3): 566–567, August 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559783.

Chatfield:2004:BR

[242] Chris Chatfield. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3): 567–568, August 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559784.

Huson:2004:BR

[243] Les Huson. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167 (3):568, August 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559785.

Moreno:2004:BR

[244] Victor Moreno. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3): 568–569, August 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559786.

Hewson:2004:BR

[245] Paul Hewson. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167 (3):569, August 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559787.

Kemp:2004:BRc

[246] David Kemp. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3): 569–570, August 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559788.

Roddam:2004:BR

[247] Andrew W. Roddam. Book review.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3): 570–571, August 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559789.

Rushton:2004:BR

[248] Lesley Rushton. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3): 571–572, August 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559790.

Wooff:2004:BR

[249] David Wooff. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167 (3):572, August 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559791.

Best:2004:BRa

[250] Mark A. Best. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3): 572–573, August 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559792.

Bartkowiak:2004:BR

[251] Anna Bartkowiak. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3):

573–574, August 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559793.

Anonymous:2004:FMc

[252] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(3):??, ???? 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559770.

Lynn:2004:EMC

[253] Peter Lynn. Editorial: Measuring and communicating survey quality. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4): 575–578, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559876.

Copas:2004:SEC

[254] Andrew J. Copas, Vern T. Farewell, Catherine H. Mercer, and Guiqing Yao. The sensitivity of estimates of the change in population behaviour to realistic changes in bias in repeated surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4): 579–595, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559877.

Griffiths:2004:CUM

[255] Paula L. Griffiths, James J. Brown, and Peter W. F. Smith. A comparison of univariate and multivariate multilevel models for repeated measures of use of antenatal care in Uttar Pradesh. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4): 597–611, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559878.

Dickens:2004:NMW

[256] Richard Dickens and Alan Manning.

Has the national minimum wage reduced UK wage inequality? Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4): 613–626, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559879.

Yang:2004:BBS

[257] Tae Young Yang. Bayesian binary segmentation procedure for detecting streakiness in sports. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4):627– 637, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559880.

Koop:2004:MED

[258] Gary Koop. Modelling the evolution of distributions: An application to major league baseball. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4):639–655, November 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559881.

Allsopp:2004:RTA

[259] P. E. Allsopp and Stephen R. Clarke.

Rating teams and analysing outcomes in one-day and test cricket. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4): 657–667, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559882.

Grieve:2004:RCS

[260] Andy Grieve. Report of the Council for the session 2003–2004. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4): 669–756, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559883.

Bjerkholt:2004:OPJ

[261] Olav Bjerkholt and Svein Longva. Obituary: Petter Jakob Bjerve. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4):757– 758, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559884.

Anonymous:2004:BMd

[262] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4):759, ???? 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559898.

Anonymous:2004:BRd

[263] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4): 759, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Burrell:2004:BR

[264] Quentin L. Burrell. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4): 760, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559885.

Robinson:2004:BR

[265] Derek Robinson. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4): 760, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559886.

Bingham:2004:BRa

[266] N. H. Bingham. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4): 761–762, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559887.

Hershberger:2004:BR

[267] Scott L. Hershberger. Book review.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 167 (4):762, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559888.

Lyn:2004:BR

[268] Jennifer Lyn. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4): 763, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559889.

Norat:2004:BRa

[269] Mohamed Afzal Norat. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4): 763–764, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559890.

Best:2004:BRc

[270] Mark A. Best. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4): 764, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559891.

Espejo:2004:BR

[271] Mariano Ruiz Espejo. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167 (4):764, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559892.

Scott:2004:BR

[272] Marian Scott. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4): 764–765, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559893.

Harbord:2004:BR

[273] Roger Harbord. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4):

765–766, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559894.

Norat:2004:BRb

[274] Mohamed Afzal Norat. Book review.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 167 (4):766, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559895.

Bingham:2004:BRb

[275] N. H. Bingham. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4): 767–768, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559896.

Barber:2004:BR

[276] Stuart Barber. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4): 768, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559897.

Anonymous:2004:IAV

[277] Anonymous. Index of authors: volume 167, 2004. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4):769–773, November 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2004:VI

[278] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4): 769–773, ???? 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559874.

Anonymous:2004:CV

[279] Anonymous. Contents of volume 167, 2004. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4):??, November 2004. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2004:FMd

[280] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 167(4):??, ???? 2004. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559875.

Bird:2005:PIG

[281] Sheila M. Bird, Sir David Cox, Vern T. Farewell, Harvey Goldstein, Tim Holt, and Peter C. Smith. Performance indicators: Good, bad, and ugly. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1): 1–27, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559706.

Tranmer:2005:CSA

[282] Mark Tranmer, Andrew Pickles, Ed Fieldhouse, Mark Elliot, Angela Dale, Mark Brown, David Martin, David Steel, and Chris Gardiner. The case for small area microdata. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1):29–49, January 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559707.

Xie:2005:CTM

[283] H. Xie, T. J. Chaussalet, and P. H.

Millard. A continuous time Markov model for the length of stay of elderly people in institutional long-term care. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168 (1):51–61, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559708.

Hancock:2005:QSS

[284] Ruth Hancock and Geraldine Barker.

The quality of social security benefit data in the British family resources survey: Implications for investigating income support take-up by pensioners. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168 (1):63–82, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559709.

Grilli:2005:REP

[285] Leonardo Grilli. The random-effects proportional hazards model with grouped survival data: A comparison between the grouped continuous and continuation ratio versions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1):83–94, January 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559710.

Wood:2005:UA

[286] Graham R. Wood and David J. Saville. The ubiquitous angle. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1): 95–107, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559711.

Imbens:2005:RAC

[287] Guido W. Imbens and Paul R. Rosenbaum. Robust, accurate confidence intervals with a weak instrument: Quarter of birth and education. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1): 109–126, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559712.

Lassarre:2005:ERM

[288] Sylvain Lassarre and Isabelle Thomas. Exploring road mortality ratios in Europe: National versus regional realities. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1): 127–144, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559713.

Kauermann:2005:SNF

[289] Göran Kauermann, Gerhard Tutz, and Josef Brüderl. The survival of newly founded firms: A case-study into varying-coefficient models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1): 145–158, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559714.

Fielding:2005:GLM

[290] Antony Fielding and Min Yang. Generalized linear mixed models for ordered responses in complex multilevel structures: Effects beneath the school or college in education. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1):159–183, January 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559715.

Reiter:2005:RMI

[291] Jerome P. Reiter. Releasing multiply imputed, synthetic public use microdata: An illustration and empirical study. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1):185–205, January 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559716.

Kalwij:2005:PDA

[292] Adriaan S. Kalwij and Mary Gregory.

A panel data analysis of the effects of wages, standard hours and unionization on paid overtime work in Britain. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1): 207–231, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559717.

Yip:2005:CSR

[293] Paul S. F. Yip, K. F. Lam, Eric H. Y. Lau, Pui-Hing Chau, Kenneth W. Tsang, and Anne Chao. A comparison study of realtime fatality rates: Severe acute respiratory syndrome in Hong Kong, Singapore, Taiwan, Toronto and Beijing, China. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1):233–243, January 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/

3559718. See corrigenda [353, 354].

Berry:2005:OSG

[294] Donald A. Berry. Obituaries: Seymour Geisser, 1929–2004. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1):245–246, January 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3559719.

Croham:2005:OSD

[295] Lord Croham. Obituaries: Sir Donald Macdougall OBE, 1912–2004. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1): 246–247, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559720.

Anonymous:2005:BMa

[296] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1):249, ???? 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3559739.

Anonymous:2005:BRa

[297] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168 (1):249, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Jones:2005:BR

[298] M. C. Jones. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1):250, ???? 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559721.

Jones:2005:BRB

[299] M. C. Jones. Book review: Tools for Constructing Chronologies; Crossing Disciplinary Boundaries. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168 (1):250, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Schilling:2005:BRB

[300] René L. Schilling. Book review: Financial Modelling with Jump Processes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1): 250–251, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559722.

Brix:2005:BRB

[301] Anders Brix. Book review: Bayesian Data Analysis, 2nd edn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1): 251–252, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559723.

Ravi:2005:BRB

[302] Sreenivasan Ravi. Book review: Multilevel Statistical Models, 3rd edn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1): 252–253, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559724.

Norat:2005:BRBa

[303] Mohamed Afzal Norat. Book review: Automatic Nonuniform Random Variate Generation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1):253–254, January 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559725.

Hewson:2005:BRBa

[304] Paul Hewson. Book review: Statistical Tools for Nonlinear Regression: a Practical Guide with S-PLUS and R Examples, 2nd edn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1):254, January 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559726.

Robinson:2005:BRB

[305] Derek Robinson. Book review: Statistics for Epidemiology. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1):254–255, January 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559727.

Kumar:2005:BRB

[306] Kuldeep Kumar. Book review: Forecast Verification: a Practitioner’s Guide in Atmospheric Sciences. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1):255, January 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic). URL http://www.jstor.org/
stable/3559728.

Norat:2005:BRBb

[307] Mohamed Afzal Norat. Book review:

Diagnostic Checks in Time Series. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168 (1):256, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559730.

Stoyanov:2005:BRBa

[308] Jordan Stoyanov. Book review: Polynomial Parameter Estimations of Close to Gaussian Random Variables. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168 (1):256, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559729.

Roddam:2005:BRBa

[309] Andrew W. Roddam. Book review:

Statistical Estimation of Epidemiological Risk. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1):256–257, January 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559731.

Burrell:2005:BRB

[310] Quentin L. Burrell. Book review:

Risk Analysis in Finance and Insurance. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1):

257–258, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559732.

Sisson:2005:BRB

[311] Scott A. Sisson. Book review: Statistical Inference and Simulation for Spatial Point Processes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1):258–259, January 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559733.

Lindley:2005:BRB

[312] Dennis Lindley. Book review: Kendall’s Advanced Theory of Statistics, volume 2B, Bayesian Inference, 2nd edn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1): 259–260, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559734.

Ecob:2005:BRB

[313] Russell Ecob. Book review: Measures of Interobserver Agreement. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1): 260–261, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559735.

Barber:2005:BRB

[314] Stuart Barber. Book review: All of Statistics: a Concise Course in Statistical Inference. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1):261, January 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559737.

Best:2005:BRBa

[315] Mark A. Best. Book review: Bayesian Approaches to Clinical Trials and Health-Care Evaluation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1):261, January 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic). URL http://www.jstor.org/
stable/3559736.

Dialsingh:2005:BRBa

[316] Isaac Dialsingh. Book review: Biostatistics and Epidemiology: a Primer for Biomedical Professionals. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168 (1):262, January 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559738.

Anonymous:2005:FMa

[317] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(1):??, ???? 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559705.

Brookmeyer:2005:EBR

[318] Ron Brookmeyer. Editorial: Biosecurity and the role of statisticians. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2): 263–266, March 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559962.

Greenland:2005:MBM

[319] Sander Greenland. Multiple-bias modelling for analysis of observational data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168 (2):267–306, March 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559963.

Saisana:2005:USA

[320] M. Saisana, A. Saltelli, and S. Tarantola. Uncertainty and sensitivity analysis techniques as tools for the quality assessment of composite indicators. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168 (2):307–323, March 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559964.

Atkinson:2005:TIU

[321] A. B. Atkinson. Top incomes in the UK over the 20th Century. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2): 325–343, March 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559965.

Vansteelandt:2005:VSP

[322] S. Vansteelandt, E. Goetghebeur, I. Thomas, E. Mathys, and F. Van Loock. On the viral safety of plasma pools and plasma derivatives. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2): 345–363, March 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559966.

Dorsett:2005:UCL

[323] Richard Dorsett. Unemployed couples:

The labour market effects of making both partners search for work. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2): 365–385, March 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559967.

Udina:2005:EPC

[324] Frederic Udina and Pedro Delicado.

Estimating parliamentary composition through electoral polls. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):387–399, March 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic). URL http://www.jstor.org/
stable/3559968.

Smith:2005:MEP

[325] Peter C. Smith and Andrew Street.

Measuring the efficiency of public services: The limits of analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2): 401–417, March 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559969.

Gregory:2005:BBG

[326] Ian N. Gregory and Paul S. Ell. Breaking the boundaries: Geographical approaches to integrating 200 years of the census. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):419–437, March 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559970.

Manda:2005:AFM

[327] Samuel Manda and Renate Meyer. Age at first marriage in Malawi: A Bayesian multilevel analysis using a discrete timeto-event model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):439–455, March 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3559971.

Anonymous:2005:BMb

[328] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):457, ???? 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3559990.

Anonymous:2005:BRb

[329] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168 (2):457, March 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hewson:2005:BR

[330] Paul Hewson. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):458, ???? 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3559972.

Hewson:2005:BRBb

[331] Paul Hewson. Book review: Exploring Multivariate Data with the Forward Search. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):458, March 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Espejo:2005:BRBa

[332] Mariano Ruiz Espejo. Book review: Environmental Statistics: Methods and Applications. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):458–459, March 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559973.

Maybank:2005:BR

[333] S. J. Maybank. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):459, ???? 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559974.

Maybank:2005:BRBa

[334] S. J. Maybank. Book review: Procrustes Problems. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):459, March 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Thomas:2005:BRB

[335] Len Thomas. Book review: State Space and Unobserved Component Models: Theory and Applications. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2): 459–460, March 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559975.

Chatfield:2005:BRBa

[336] Chris Chatfield. Book review: Statistical Analysis and Data Display. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2): 460–461, March 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559976.

Espejo:2005:BRa

[337] Mariano Ruiz Espejo. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):461, ???? 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559977.

Espejo:2005:BRBb

[338] Mariano Ruiz Espejo. Book review: Consumer Price Index Manual: Theory and Practice. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):461, March 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Broom:2005:BRB

[339] Mark Broom. Book review: Introduction to Mathematical Methods in Bioinformatics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):461–462, March 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/3559978.

McCulloch:2005:BRBa

[340] Andrew McCulloch. Book review: Spatial Evolutionary Modeling. Journal of the Royal Statistical Society.

Series A (Statistics in Society), 168 (2):462, March 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559979.

Craig:2005:BRB

[341] Peter Craig. Book review: Bayesian Field Theory. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):462–463, March 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559980.

Roddam:2005:BRBb

[342] Andrew W. Roddam. Book review:

Analysing Survival Data from Clinical Trials and Observational Studies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):463, March 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559981.

Stoyanov:2005:BRBb

[343] Jordan Stoyanov. Book review: Multivariate Probability. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):463–464, March 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559982.

Roddam:2005:BR

[344] Andrew W. Roddam. Book review.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):464, ???? 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559983.

Roddam:2005:BRBc

[345] Andrew W. Roddam. Book review: Statistics for the Quality Control Chemistry Laboratory. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):464, March 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

LeFevre:2005:BRB

[346] Andrea Le Fevre. Book review: Statistical Power Analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):465, March 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3559984.

Sapatinas:2005:BRBa

[347] Theofanis Sapatinas. Book review: Statistics and Finance. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):465–466, March 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3559985.

McCulloch:2005:BRBb

[348] Andrew McCulloch. Book review: Sensitivity Analysis in Practice: a Guide to Assessing Scientific Models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168 (2):466, March 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559986.

Wright:2005:BR

[349] Lesley F. Wright. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):466, ???? 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559987.

Wright:2005:BRB

[350] Lesley F. Wright. Book review: Risk and Financial Management. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168 (2):466, March 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2005:BRBa

[351] Anonymous. Book review: Biostatistics: a Methodology for the Health Sciences. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):467, March 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Driver:2005:BR

[352] Ron Driver. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):467, ???? 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559988.

Anonymous:2005:CCS

[353] Anonymous. Corrigendum: A comparison study of realtime fatality rates: Severe acute respiratory syndrome in Hong Kong, Singapore, Taiwan, Toronto and Beijing, China. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):468, ???? 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559989. See

[293].

Yip:2005:CCS

[354] Paul S. F. Yip, K. F. Lam, Eric H. Y. Lau, Pui-Hing Chau, Kenneth W. Tsang, and Anne Chao. Corrigendum: A comparison study of realtime fatality rates: severe acute respiratory syndrome in Hong Kong, Singapore, Taiwan, Toronto and Beijing, China. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):468, March 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [293].

Anonymous:2005:FMb

[355] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(2):??, ???? 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559961.

Longford:2005:EMS

[356] N. T. Longford. Editorial: Model selection and efficiency: Is ‘which model … ?’ the right question? Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3): 469–472, July 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559835.

Blundell:2005:EEE

[357] Richard Blundell, Lorraine Dearden, and Barbara Sianesi. Evaluating the effect of education on earnings: Models, methods and results from the National Child Development Survey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3): 473–512, July 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559836.

Shields:2005:EES

[358] Michael A. Shields and Stephen Wheatley Price. Exploring the economic and social determinants of psychological well-being and perceived social support in England. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):513–537, July 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559837.

Francesconi:2005:ECF

[359] Marco Francesconi. An evaluation of the childhood family structure measures from the sixth wave of the British Household Panel Survey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):539–566, July 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559838.

Nurminen:2005:WLE

[360] Markku M. Nurminen, Christopher R.

Heathcote, Brett A. Davis, and Borek D. Puza. Working life expectancies: The case of Finland 1980–2006. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3): 567–581, July 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559839.

Yu:2005:FAR

[361] Philip L. H. Yu, K. F. Lam, and S. M. Lo. Factor analysis for ranked data with application to a job selection attitude survey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):583–597, July 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/3559840.

Browne:2005:VPM

[362] W. J. Browne, S. V. Subramanian, K. Jones, and H. Goldstein. Variance partitioning in multilevel logistic models that exhibit overdispersion. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168 (3):599–613, July 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559841.

Rice:2005:MPS

[363] Adrian Rice and Eugene Seneta. De Morgan in the prehistory of statistical hypothesis testing. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):615–627, July 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3559842.

Anonymous:2005:BMc

[364] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):629, ???? 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559856.

Anonymous:2005:BRc

[365] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):629, July 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Griffin:2005:BR

[366] Jamie Griffin. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):630, ???? 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559843.

Griffin:2005:BRB

[367] Jamie Griffin. Book review: Handbook of Statistics, Volume 23, Advances in Survival Analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):630, July 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Smart:2005:BRB

[368] John Smart. Book review: Statistics in Market Research. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):630–631, July 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559844.

Karlsson:2005:BRB

[369] Andreas Karlsson. Book review: Tutorials in Biostatistics, Volume 1, Statistical Methods in Clinical Studies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):

631–632, July 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559845.

Tusell:2005:BRB

[370] Fernando Tusell. Book review: Linear Models with R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):632, July 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559846.

Chatfield:2005:BRBb

[371] Chris Chatfield. Book review: Periodic Time Series Models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3): 632–633, July 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559847.

Dialsingh:2005:BRBb

[372] Isaac Dialsingh. Book review: Applied Bayesian Modeling and Causal Inference from Incomplete Data Perspectives. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):633–634, July 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559848.

Maybank:2005:BRBb

[373] S. J. Maybank. Book review: Highly Structured Stochastic Systems. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):634, July 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559849.

Bailey:2005:BRB

[374] R. A. Bailey. Book review: Designing Experiments and Analyzing Data: a Model Comparison Perspective, 2nd edn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):634–635, July 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559850.

Sapatinas:2005:BRBb

[375] Theofanis Sapatinas. Book review: Discriminant Analysis and Statistical Pattern Recognition. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):635–636, July 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559851.

Best:2005:BR

[376] Mark A. Best. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):636, ???? 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559852.

Best:2005:BRBb

[377] Mark A. Best. Book review: Statistical Process Control, 5th edn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):636, July 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Cannings:2005:BR

[378] Chris Cannings. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):636, ???? 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3559853.

Cannings:2005:BRB

[379] Chris Cannings. Book review: Random Geometric Graphs. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):636, July 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Campanelli:2005:BRB

[380] Pamela Campanelli. Book review: Methods for Testing and Evaluating Survey Questionnaires, by S. Presser; J. M. Rothgeb; M. P. Couper; J. T. Lessler; E. Martin; J. Martin; E. Singer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):637, July 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3559854.

Espejo:2005:BRBc

[381] Mariano Ruiz Espejo. Book review: Quantifying the World: UN Ideas and Statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):637–638, July 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/3559855.

Anonymous:2005:FMc

[382] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(3):??, ???? 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559834.

Anonymous:2005:DAP

[383] Anonymous. Dr Andrew P. Grieve.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 168 (4):i, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Grieve:2005:PSC

[384] Andrew P. Grieve. The professionalization of the ‘shoe clerk’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4): 639–656, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559938.

Warwick:2005:RCS

[385] Jane Warwick and Stephen W. Duffy.

A review of cancer screening evaluation techniques, with some particular examples in breast cancer screening. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4): 657–677, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559939.

Congdon:2005:TII

[386] P. Congdon and H. Southall. Trends in inequality in infant mortality in the North of England, 1921–1973, and their association with urban and social structure. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4):679–700, November 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559940.

Trevelyan:2005:SSE

[387] Barry Trevelyan, Matthew Smallman-Raynor, and Andrew D. Cliff. The spatial structure of epidemic emergence: Geographical aspects of poliomyelitis in North-Eastern USA, July–October 1916. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4):701–722, November 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559941.

Fitzmaurice:2005:PEL

[388] Garrett M. Fitzmaurice, Stuart R. Lipsitz, Geert Molenberghs, and Joseph G. Ibrahim. A protective estimator for longitudinal binary data subject to non-ignorable non-monotone missingness. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4):723–735, November 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559942.

Jackson:2005:MRB

[389] Dan Jackson, John Copas, and Alex J. Sutton. Modelling reporting bias: The operative mortality rate for ruptured abdominal aortic aneurysm repair. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4): 737–752, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559943.

Baker:2005:RPI

[390] Stuart G. Baker, Grant Izmirlian, and Victor Kipnis. Resolving paradoxes involving surrogate end points. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4): 753–762, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559944.

Nicoletti:2005:SRS

[391] Cheti Nicoletti and Franco Peracchi.

Survey response and survey characteristics: Microlevel evidence from the European Community Household Panel. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4): 763–781, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559945.

Grieve:2005:RCS

[392] Andy Grieve. Report of the Council for the session 2004–2005. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4): 783–863, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559946.

Armitage:2005:OJA

[393] Peter Armitage. Obituary: James Austin Heady, 1917–2004. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4):865– 867, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559947.

Anonymous:2005:BMd

[394] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4):869, ???? 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3559960.

Anonymous:2005:BRd

[395] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4): 869, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ravi:2005:BR

[396] Sreenivasan Ravi. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4): 870, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559948.

McCrorie:2005:BR

[397] J. Roderick McCrorie. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4): 870–871, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559949.

Lang:2005:BR

[398] Iain Lang. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4): 871, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559950.

Hutchison:2005:BRa

[399] Dougal Hutchison. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4): 871–872, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559951.

Goldie:2005:BR

[400] Charles M. Goldie. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4): 872, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559952.

Stoyanov:2005:BR

[401] Jordan Stoyanov. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4): 872–873, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559953.

Brix:2005:BR

[402] Anders Brix. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4): 873–874, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559954.

Little:2005:BR

[403] Mark P. Little. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4): 874–875, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559955.

Hutchison:2005:BRb

[404] Dougal Hutchison. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4): 875, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559956.

Chappell:2005:BR

[405] Francesca Chappell. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4): 875–876, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559957.

Lewin:2005:BR

[406] Alex Lewin and Anne-Mette Hein. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4):876–877, November 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/3559958.

Espejo:2005:BRc

[407] Mariano Ruiz Espejo. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4): 877–878, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559959.

Anonymous:2005:IAV

[408] Anonymous. Index of authors: volume 168, 2005. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4):879–884, November 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2005:VI

[409] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4): 879–884, ???? 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559936.

Anonymous:2005:VCC

[410] Anonymous. Volume contents: Contents of volume 168, 2005. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4):885– 886, November 2005. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2005:FMd

[411] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168(4):??, ???? 2005. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559937.

Zidek:2006:EPN

[412] James V. Zidek. Editorial: (postnormal) statistical science. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1): 1–4, January 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559901.

Ades:2006:MES

[413] A. E. Ades and A. J. Sutton. Multiparameter evidence synthesis in epidemiology and medical decision-making: Current approaches. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):5–35, January 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559902.

Hahn:2006:REI

[414] Eugene D. Hahn. Re-examining informative prior elicitation through the lens of Markov chain Monte Carlo methods. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169 (1):37–48, January 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559903.

Abakuks:2006:SST

[415] Andris Abakuks. A statistical study of the triple-link model in the synoptic problem. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):49–60, January 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559904.

Moerkerke:2006:SIT

[416] B. Moerkerke, E. Goetghebeur, J. De Riek, and I. Roldán-Ruiz. Significance and impotence: Towards a balanced view of the null and the alternative hypotheses in marker selection for plant breeding. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):61–79, January 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559905.

Fay:2006:MAB

[417] Michael P. Fay and Ji-Hyun Lee. Measuring agreement between two statistics with applications to age standardization. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):81–96, January 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559906.

Pettitt:2006:BHM

[418] A. N. Pettitt, T. T. Tran, M. A. Haynes, and J. L. Hay. A Bayesian hierarchical model for categorical longitudinal data from a social survey of immigrants. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1): 97–114, January 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559907.

Austin:2006:CCD

[419] Peter C. Austin and Jack V. Tu. Comparing clinical data with administrative data for producing acute myocardial infarction report cards. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1): 115–126, January 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559908.

Taylor:2006:TMW

[420] Mark P. Taylor. Tell me why I don’t like Mondays: Investigating day of the week effects on job satisfaction and psychological well-being. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):127–142, January 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559909.

Stone:2006:HFH

[421] Mervyn Stone and Jane Galbraith. How not to fund hospital and community health services in England. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1): 143–164, January 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559910.

Whitehead:2006:OJH

[422] Frank E. Whitehead. Obituary: John Harold West, MBE. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):165–166, January 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559911.

Anonymous:2006:BMa

[423] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):167, ???? 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559923.

Anonymous:2006:BRa

[424] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169 (1):167, January 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Sisson:2006:BRB

[425] Scott Sisson. Book review: Statistics of Extremes: Theory and Applications, by J. Beirlant; Y. Goegebeur; J. Segers; J. Teugels. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):168–169, January 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559912.

Dialsingh:2006:BR

[426] Isaac Dialsingh. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169 (1):169, January 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dialsingh:2006:BRBa

[427] Isaac Dialsingh. Book review: Introduction to Rare Event Simulation, by J. A. Bucklew. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):169, ???? 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559913.

Bingham:2006:BRBa

[428] N. H. Bingham. Book review: Statistics: An Introduction Using R, by M. J. Crawley. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):169–170, January 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559914.

Broom:2006:BR

[429] Mark Broom. Book reviews. Journal of the Royal Statistical Society.

Series A (Statistics in Society), 169 (1):170, January 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Broom:2006:BRBa

[430] Mark Broom. Book review: Statistical Methods in Bioinformatics: An Introduction, by W. J. Ewens; G. R. Grant. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):170, ???? 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559915.

Karlsson:2006:BRB

[431] Andreas Karlsson. Book review: Permutation, Parametric, and Bootstrap Tests of Hypotheses, by P. Good. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169 (1):171, January 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559916.

Bingham:2006:BRBb

[432] N. H. Bingham. Book review: Probability: A Graduate Course, by A. Gut. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1): 171–172, January 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559917.

Liao:2006:BRB

[433] Tim Futing Liao. Book review: Ecological Inference: New Methodological Strategies, by G. King; O. Rosen; M. A. Tanner. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):172–173, January 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559918.

Espejo:2006:BR

[434] Mariano Ruiz Espejo and Housila P.

Singh. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):173, January 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Espejo:2006:BRBa

[435] Mariano Ruiz Espejo and Housila P.

Singh. Book review: Statistical Design of Experiments with Engineering Applications, by K. Rekab; M. Shaikh. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):173, ???? 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559920.

Nunn:2006:BR

[436] Susan Nunn. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169 (1):173, January 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Nunn:2006:BRBa

[437] Susan Nunn. Book review: Analysing Environmental Data, by W. Piegorsch; A. J. Bailer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):173, ???? 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559919.

Robinson:2006:BR

[438] Andrew Robinson. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169 (1):174, January 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Robinson:2006:BRBa

[439] Andrew Robinson. Book review: Statistical Methods for Spatial Data Analysis, by O. Schabenberger; C. A. Gotway. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):174, ???? 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559921.

Anonymous:2006:BRb

[440] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1): 174–175, January 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shalabh:2006:BRBa

[441] Shalabh. Book review: Regression Methods in Biostatistics, by E. Vittinghoff; D. V. Glidden; S. C. Shiboski; C. E. Mc-Culloch. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):174–175, ???? 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/3559922.

Anonymous:2006:VIa

[442] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):

177–178, ???? 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559899.

Farebrother:2006:SIO

[443] Richard William Farebrother and Peter Neal. Subject index of obituary notices, 1994–2005: Subject index of obituary notices, 1994–2005. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):177–178, January 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2006:FMa

[444] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(1):??, ???? 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559900.

Peng:2006:MCT

[445] Roger D. Peng, Francesca Dominici, and Thomas A. Louis. Model choice in time series studies of air pollution and mortality. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2):179–203, March 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559669.

Tseloni:2006:MMN

[446] Andromachi Tseloni. Multilevel modelling of the number of property crimes: Household and area effects. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2): 205–233, March 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559670.

Sahu:2006:BMS

[447] S. K. Sahu and T. M. F. Smith. A Bayesian method of sample size determination with practical applications. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2): 235–253, March 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559671.

Soroka:2006:PEU

[448] Stuart N. Soroka, Christopher Wlezien, and Iain McLean. Public expenditure in the UK: How measures matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2): 255–271, March 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559672.

Gurka:2006:EBC

[449] Matthew J. Gurka, Lloyd J. Edwards, Keith E. Muller, and Lawrence L. Kupper. Extending the Box–Cox transformation to the linear mixed model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2): 273–288, March 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559673.

Jeong:2006:NPF

[450] Jong-Hyeon Jeong. A new parametric family for modelling cumulative incidence functions: Application to breast cancer data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2):289–303, March 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559674.

Lensvelt-Mulders:2006:VCA

[451] Gerty J. L. M. Lensvelt-Mulders, Peter G. M. Van Der Heijden, Olav Laudy, and Ger Van Gils. A validation of a computer-assisted randomized response survey to estimate the prevalence of fraud in social security. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2): 305–318, March 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559675.

Gardner:2006:DDC

[452] Jonathan Gardner and Andrew J. Oswald. Do divorcing couples become happier by breaking up? Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2): 319–336, March 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559676.

Chandola:2006:PBE

[453] Tarani Chandola, Paul Clarke, J. N.

Morris, and David Blane. Pathways between education and health: A causal modelling approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2): 337–359, March 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559677.

Gormley:2006:AIT

[454] Isobel Claire Gormley and Thomas Brendan Murphy. Analysis of Irish thirdlevel college applications data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2): 361–379, March 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559678.

Swires-Hennessy:2006:OHW

[455] Ed Swires-Hennessy. Obituaries: Harold William George Gearing, BSc, FIS, FCIS, FBCS, 1912–2005. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2): 381–382, March 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3559679.

Lynch:2006:OSJ

[456] Robin Lynch. Obituaries: Sir Jack Hibbert, 1932–2005. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2):382–383, March 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3559680.

Anonymous:2006:BMb

[457] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2):385, ???? 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3559692.

Anonymous:2006:BRR

[458] Anonymous. Book reviews: Reviews in this issue. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2):385, March 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ravi:2006:BRB

[459] Sreenivasan Ravi. Book review: Association Schemes: Designed Experiments, Algebra and Combinatorics, by R. A. Bailey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2):386, March 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559682.

Whitlock:2006:BRB

[460] Mark Whitlock. Book review: Statistical Treatment of Analytical Data, by Z. B. Alfassi; Z. Boger; Y. Ronen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2):386, March 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559681.

Espejo:2006:BRBb

[461] Mariano Ruiz Espejo. Book review:

Quality by Experimental Design, by T. B. Barker. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2):386–387, March 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559683.

Tong:2006:BRB

[462] So Moon Tong. Book review: Bayesian Models for Categorical Data, by P. Congdon. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2):387, March 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559684.

Driver:2006:BRB

[463] Ron Driver. Book review: An Introduction to Multivariate Data Analysis, by T. R. Cox; Multivariate Statistical Methods by D. F. Morrison. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2): 387–389, March 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559685.

Bartholomew:2006:BRB

[464] David Bartholomew. Book review: Celebrating Statistics: Papers in Honour of Sir David Cox on His 80th Birthday, by A. C. Davison; Y. Dodge; N. Wermuth. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2):389, March 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559686.

Shalabh:2006:BRBb

[465] Shalabh. Book review: Introduction to Linear Models and Statistical Inference, by S. J. Janke; F. C. Tinsley. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2): 389–390, March 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559687.

Dialsingh:2006:BRBb

[466] Isaac Dialsingh. Book review: Statistical Analysis of Stochastic Processes in Time, by J. K. Lindsey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169 (2):390, March 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559688.

Hannachi:2006:BRB

[467] A. Hannachi. Book review: Exploratory Data Analysis with MATLAB, by W. L. Martinez; A. R. Martinez. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2): 390–391, March 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559689.

Cannings:2006:BRB

[468] Chris Cannings. Book review: Statistical Methods in Molecular Evolution, by R. Nielsen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2):391, March 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3559690.

Nunn:2006:BRBb

[469] Susan Nunn. Book review: Analysing Categorical Data, by J. S. Simonoff. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169 (2):391–392, March 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3559691.

Anonymous:2006:FMb

[470] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(2):??, ???? 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3559668.

Lynn:2006:EAN

[471] Peter Lynn. Editorial: Attrition and non-response. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):393–394, July 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3877427.

Haslett:2006:BPR

[472] J. Haslett, M. Whiley, S. Bhattacharya, M. Salter-Townshend, Simon P. Wilson, J. R. M. Allen, B. Huntley, and F. J. G. Mitchell. Bayesian palaeoclimate reconstruction. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):395–438, July 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3877428.

Groves:2006:RDH

[473] Robert M. Groves and Steven G. Heeringa. Responsive design for household surveys: Tools for actively controlling survey errors and costs. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3): 439–457, July 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877429.

Burton:2006:LTE

[474] Jonathan Burton, Heather Laurie, and Peter Lynn. The long-term effectiveness of refusal conversion procedures on longitudinal surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):459–478, July 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3877430.

Hawkes:2006:MNR

[475] Denise Hawkes and Ian Plewis. Modelling non-response in the national child development study. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):479–491, July 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3877431.

Nitsch:2006:LBE

[476] D. Nitsch, B. L. DeStavola, S. M. B.

Morton, and D. A. Leon. Linkage bias in estimating the association between childhood exposures and propensity to become a mother: An example of simple sensitivity analyses. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3): 493–505, July 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877432.

Longford:2006:ALT

[477] Nicholas T. Longford, P. Tyrer, U. A. M.

Nur, and H. Seivewright. Analysis of a long-term study of neurotic disorder, with insights into the process of nonresponse. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):507–523, July 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3877433.

Wood:2006:UNF

[478] Angela M. Wood, Ian R. White, and Matthew Hotopf. Using number of failed contact attempts to adjust for non-ignorable non-response. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3): 525–542, July 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877434.

Jones:2006:HRN

[479] Andrew M. Jones, Xander Koolman, and Nigel Rice. Health-related nonresponse in the British Household Panel Survey and European Community Household Panel: Using inverseprobability-weighted estimators in nonlinear models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):543–569, July 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3877435.

Carpenter:2006:CMI

[480] James R. Carpenter, Michael G. Kenward, and Stijn Vansteelandt. A comparison of multiple imputation and doubly robust estimation for analyses with missing data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):571–584, July 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3877436.

vandenBerg:2006:SNR

[481] Gerard J. van den Berg, Maarten Lindeboom, and Peter J. Dolton. Survey non-response and the duration of unemployment. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):585–604, July 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3877437.

Durrant:2006:UDA

[482] Gabriele B. Durrant and Chris Skinner.

Using data augmentation to correct for non-ignorable non-response when surrogate data are available: An application to the distribution of hourly pay. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169 (3):605–623, July 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877438.

Nicoletti:2006:EII

[483] Cheti Nicoletti and Franco Peracchi.

The effects of income imputation on microanalyses: Evidence from the European Community Household Panel. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169 (3):625–646, July 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877439.

Longford:2006:C

[484] N. T. Longford. Correspondence. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):

647–648, July 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Longford:2006:BR

[485] N. T. Longford, Dan Jackson, John Copas, and Alex J. Sutton. Book review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169 (3):647–649, ???? 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3877440.

Jackson:2006:CAR

[486] Dan Jackson, John Copas, and Alex J. Sutton. Correspondence: Authors’ response. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):648–649, July 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2006:BMc

[487] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):651, ???? 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3877450.

Anonymous:2006:BRc

[488] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):651, July 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Stoyanov:2006:BRBa

[489] Jordan Stoyanov. Book review: Combinatorial Methods in Discrete Distributions, by C. A. Charalambides. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):652, July 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3877442.

Ugarte:2006:BRB

[490] M. Dolores Ugarte. Book review: Statistical Modelling in GLIM4, 2nd edn, by M. Aitkin, B. Francis and J. Hinde. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):652, July 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877441.

Jones:2006:BRB

[491] M. C. Jones. Book review: Statistics for Fission Track Analysis, by R. F. Galbraith. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):652–653, July 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3877443.

Castagliola:2006:BRB

[492] Philippe Castagliola. Book review:

Computational Statistics, by G. H. Givens and J. A. Hoeting. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3): 653–654, July 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877444.

Brooms:2006:BRB

[493] A. C. Brooms. Book review: Stochastic Approximation and Recursive Algorithms with Applications, 2nd edn by H.

J. Kushner and G. G. Yin. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):654, July 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3877445.

Lang:2006:BRB

[494] Iain Lang. Book review: Spatial and Syndromic Surveillance for Public Health, by A. B. Lawson and K. Kleinman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):655, July 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877446.

Stoimenova:2006:BRB

[495] Eugenia Stoimenova. Book review: Constrained Statistical Inference: Inequality, Order and Shape Restrictions, by M. J. Silvapulle and P. K. Sen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169 (3):655–656, July 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877447.

Robinson:2006:BRBb

[496] Derek Robinson. Book review: The Statistical Evaluation of Medical Tests for Classification and Prediction, by M. Sullivan Pepe. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):656, July 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877448.

Espejo:2006:BRBc

[497] Mariano Ruiz Espejo and Miguel Delgado Pineda. Book review: The Total Survey Error Approach: A Guide to the New Science of Survey Research, by H. F. Weisberg. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):656–657, July 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3877449.

Anonymous:2006:FMc

[498] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(3):??, ???? 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3877426.

Atkinson:2006:EMG

[499] A. B. Atkinson. Editorial: Measurement of government output and productivity. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4): 659–662, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877393.

Hills:2006:BTD

[500] John Hills. From Beveridge to Turner:

Demography, distribution and the future of pensions in the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4): 663–679, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877394.

Lancaster:2006:RBE

[501] Gillian A. Lancaster, Mick Green, and Steven Lane. Reducing bias in ecological studies: An evaluation of different methodologies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):681–700, October 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3877395.

Jenkins:2006:PCE

[502] Stephen P. Jenkins, Lorenzo Cappellari, Peter Lynn, Annette Jäckle, and Emanuela Sala. Patterns of consent: Evidence from a general household survey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4): 701–722, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877396.

Chung:2006:LCL

[503] Hwan Chung, Brian P. Flaherty, and Joseph L. Schafer. Latent class logistic regression: Application to marijuana use and attitudes among high school seniors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):723–743, October 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877397.

Roberts:2006:EPM

[504] John M. Roberts, Jr. and Devon D. Brewer. Estimating the prevalence of male clients of prostitute women in Vancouver with a simple capture-recapture method. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):745–756, October 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3877398.

Steele:2006:FOC

[505] Fiona Steele, Constantinos Kallis, and Heather Joshi. The formation and outcomes of cohabiting and marital partnerships in early adulthood: The role of previous partnership experience. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4): 757–779, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877399.

Aassve:2006:EFU

[506] Arnstein Aassve, Simon Burgess, Carol Propper, and Matt Dickson. Employment, family union and childbearing decisions in Great Britain. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4): 781–804, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877400.

Rabe-Hesketh:2006:MMC

[507] Sophia Rabe-Hesketh and Anders Skrondal. Multilevel modelling of complex survey data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):805–827, October 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3877401.

Arulampalam:2006:SDC

[508] Wiji Arulampalam and Sonia Bhalotra.

Sibling death clustering in India: State dependence versus unobserved heterogeneity. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):829–848, October 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/3877402.

Brown:2006:RPE

[509] Sarah Brown, Lisa Farrell, Mark N. Harris, and John G. Sessions. Risk preference and employment contract type. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4): 849–863, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3877403.

Gustafson:2006:SSI

[510] Paul Gustafson. Sample size implications when biases are modelled rather than ignored. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):865–881, October 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3877404.

Brown:2006:DON

[511] James Brown, Owen Abbott, and Ian Diamond. Dependence in the 2001 one-number census project. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4): 883–902, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3877405.

Ringrose:2006:NUL

[512] Trevor J. Ringrose. Neutral umpires and leg before wicket decisions in test cricket.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4): 903–911, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877406.

Holt:2006:RCS

[513] Tim Holt. Report of the Council for the session 2005–2006. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):913–1002, October 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3877407.

Anonymous:2006:BMd

[514] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):1003, ???? 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3877425.

Anonymous:2006:BRd

[515] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4): 1003, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kumar:2006:BRBa

[516] Kuldeep Kumar. Book review: Statistics in the Pharmaceutical Industry, 3rd edn, by C. R. Buncher and J. Y. Tsay (eds). Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):1004, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877408.

Dialsingh:2006:BRBc

[517] Isaac Dialsingh. Book review: Statistical Analysis of Financial Data in S-Plus, by R. A. Carmona. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4): 1004–1005, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3877409.

Bingham:2006:BRBc

[518] N. H. Bingham. Book review: Statistical Tools for Finance and Insurance, by P. Cizek, W. Härdle and R. Weron (eds). Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):1005, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3877410.

Kumar:2006:BRBb

[519] Kuldeep Kumar. Book review: Handbook of Statistics in Clinical Oncology, 2nd edn, by J. Crowley and D. P. Ankerst (eds). Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):1006, October 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3877411.

Broom:2006:BRBb

[520] Mark Broom. Book review: Computational Genome Analysis: an Introduction, by R. C. Deonier, S. Tavaré and M. S. Waterman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):1006–1007, October 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
3877412.

Hochkirchen:2006:BRBa

[521] Thomas Hochkirchen. Book review: Design and Modeling for Computer Experiments, by K.-T. Fang, R. Li and A. Sudjianto. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):1007–1008, October 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3877413.

Verbeke:2006:BRB

[522] T. Verbeke. Book review: Extending the Linear Model with R, by J. J. Faraway. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):1008, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877414.

Anonymous:2006:BRBa

[523] Anonymous. Book review: Exact Analysis of Discrete Data, by K. F. Hirji. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169 (4):1009, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shalabh:2006:BRBc

[524] Shalabh. Book review: Exact Analysis of Discrete Data, by K. F. Hirji. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169 (4):1009, ???? 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877415.

Bochkina:2006:BRB

[525] Natalia Bochkina. Book review: Probabilistic Modeling in Bioinformatics and Medical Informatics, by D. Husmeier, R. Dybowski and S. Roberts (eds). Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4): 1009–1010, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877416.

Adams:2006:BRB

[526] Niall Adams. Book review: Handbook of Parallel Computing and Statistics, by E. J. Kontoghiorghes (ed.). Journal of the Royal Statistical Society. Series A (Statistics in Society), 169 (4):1010, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877417.

Cortina-Borja:2006:BRB

[527] Mario Cortina-Borja. Book review: Beyond Beta: Other Continuous Families of Distributions with Bounded Support and Applications, by S. Kotz; J. R. van Dorp. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):1010–1011, October 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3877418.

Hewson:2006:BRB

[528] Paul Hewson. Book review: Bayesian Statistics and Marketing, by P. E. Rossi; G. M. Allenby; R. McCulloch. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4): 1011, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877419.

Rhodes:2006:BRB

[529] Sarah Rhodes. Book review: Publication Bias in Meta-analysis, by H. R. Rothstein, A. J. Sutton and M. Borenstein (eds). Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):1012, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877420.

Bingham:2006:BRBd

[530] N. H. Bingham. Book review: Measures, Integrals and Martingales, by R. L. Schilling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):1012–1013, October 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/3877421.

Stoyanov:2006:BRBb

[531] Jordan Stoyanov. Book review: Stochastic Processes and Models, by D. Stirzaker; Probability, Statistics, and Stochastic Processes, by P. Olofsson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4): 1013–1014, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/3877422.

Nunn:2006:BRBc

[532] Susan Nunn. Book review: The Hunt for the Last Respondent, by I. Stoop. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169 (4):1014, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3877423.

Hochkirchen:2006:BRBb

[533] Thomas Hochkirchen. Book review: Essentials of Statistical Inference, by G. A. Young and R. L. Smith. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4): 1014–1015, October 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3877424.

Anonymous:2006:IAV

[534] Anonymous. Index of authors: volume 169, 2006. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):1017–1022, October 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2006:VIb

[535] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4): 1017–1022, ???? 2006. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/3877391.

Anonymous:2006:FMd

[536] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):??, ???? 2006. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        3877392.

Briscoe:2007:EHP

[537] Simon Briscoe. Editorial: House price indices. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1):1–3, January 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/4623130.

Battistin:2007:CBA

[538] Erich Battistin, Enrico Rettore, and Ugo Trivellato. Choosing between alternative classification criteria to measure the labour force state. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1): 5–27, January 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623131.

Loughin:2007:ALT

[539] Thomas M. Loughin, Mollie Poehlman Roediger, George A. Milliken, and John P. Schmidt. On the analysis of long-term experiments. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1): 29–42, January 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623132.

Jurges:2007:ULS

[540] Hendrik Jürges. Unemployment, life satisfaction and retrospective error. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1): 43–61, January 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623133.

Clark:2007:SWH

[541] Robert G. Clark and David G. Steel. Sampling within households in household surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1):63–82, January 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        4623134.

Mills:2007:TSM

[542] Terence C. Mills. Time series modelling of two millennia of Northern Hemisphere temperatures: Long memory or shifting trends? Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1):83–94, January 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/4623135.

DeSantis:2007:UHD

[543] Fulvio De Santis. Using historical data for Bayesian sample size determination. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1): 95–113, January 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623136.

Bartolucci:2007:LMM

[544] Francesco Bartolucci, Fulvia Pennoni, and Brian Francis. A latent Markov model for detecting patterns of criminal activity. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1):115–132, January 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/4623137.

Lee:2007:FBT

[545] Myoung jae Lee, Unto Häkkinen, and Gunnar Rosenqvist. Finding the best treatment under heavy censoring and hidden bias. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1):133–147, January 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
4623138.

Fahey:2007:CGM

[546] Michael T. Fahey, Christopher W.

Thane, Gemma D. Bramwell, and W. Andy Coward. Conditional Gaussian mixture modelling for dietary pattern analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1):149–166, January 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/4623139.

Pollock:2007:HTS

[547] Gary Pollock. Holistic trajectories: A study of combined employment, housing and family careers by using multiplesequence analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1):167–183, January 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
4623140.

Mountford:2007:EVE

[548] William K. Mountford, Stuart R. Lipsitz, Garrett M. Fitzmaurice, Rickey E. Carter, Jeremy B. Soule, John A. Colwell, and Daniel T. Lackland. Estimating the variance of estimated trends in proportions when there is no unique subject identifier. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1):185–193, January 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
4623141.

Skinner:2007:PIA

[549] C. J. Skinner. The probability of identification: Applying ideas from forensic statistics to disclosure risk assessment. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1): 195–212, January 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623142.

Perez-Quezada:2007:ESP

[550] Jorge F. Perez-Quezada, Nicanor Z. Saliendra, William E. Emmerich, and Emilio A. Laca. Evaluation of statistical protocols for quality control of ecosystem carbon dioxide fluxes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1): 213–230, January 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623143.

Dawson:2007:FRR

[551] Peter Dawson, Stephen Dobson, John Goddard, and John Wilson. Are football referees really biased and inconsistent?: Evidence on the incidence of disciplinary sanction in the English Premier League. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1): 231–250, January 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623144.

Anonymous:2007:BMa

[552] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1):251, ???? 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
4623159.

Anonymous:2007:BRa

[553] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170 (1):251, January 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dryden:2007:BRB

[554] Ian Dryden. Book review: Stereology for Statisticians, by A.Baddeley and E. B. Vedel Jensen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1):252, January 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/4623145.

Militino:2007:BRB

[555] Ana F. Militino. Book review: Applied Mixed Models in Medicine, 2nd edn, by H. Brown and R. Prescott. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1): 252–253, January 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623146.

Dialsingh:2007:BRBa

[556] Isaac Dialsingh. Book review: Applied Longitudinal Analysis, by G. M. Fitzmaurice, N. M. Laird and J. H. Ware.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 170 (1):253, January 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623147.

Bartholomew:2007:BRB

[557] David Bartholomew. Book review: Selected Statistical Papers of Sir David Cox, Vols I and II, by D. J. Hand; A. M. Herzberg. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1):253–255, January 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/4623148.

Karlsson:2007:BRBa

[558] Andreas Karlsson. Book review: Robust Statistical Methods with R, by J. Jurečková and J. Picek. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1): 255–256, January 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623149.

Karlsson:2007:BRBb

[559] Andreas Karlsson. Book review: Quantile Regression, by R. Koenker. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170 (1):256, January 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623150.

Brooms:2007:BRB

[560] A. C. Brooms. Book review: The Basics of S-PLUS, by A. Krause; M. Olson. Journal of the Royal Statistical Society.

Series A (Statistics in Society), 170(1): 256–257, January 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623151.

Robinson:2007:BRBa

[561] Derek Robinson. Book review: SAS for Mixed Models, by R. C. Littell; G. A. Milliken; W. W. Stroup; R. D. Wolfinger; O. Schabenberger. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1): 257–258, January 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623152.

Bakouch:2007:BRBa

[562] Hassan S. Bakouch. Book review: Introductory Statistical Inference, by N. Mukhopadhyay. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1):258–259, January 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
4623153.

Davis:2007:BRBa

[563] J. Wade Davis. Book review: R Graphics, by P. Murrell. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1):259–260, January 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
4623154.

Espejo:2007:BRBa

[564] Mariano Ruiz Espejo. Book review:

The Internet as a Large-scale Complex System, by K. Park and W. Willinger.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 170 (1):260, January 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623155.

Davis:2007:BRBb

[565] J. Wade Davis. Book review: Clinical Trials: a Methodological Perspective, 2nd edn, by S. Piantadosi. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1): 260–261, January 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623156.

Haigh:2007:BRB

[566] John Haigh. Book review: Stochastic Modelling for Systems Biology, by D. J. Wilkinson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1):261, January 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/4623157.

Verbeke:2007:BRB

[567] T. Verbeke. Book review: Generalized Additive Models: An Introduction with R, by S. N. Wood. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170 (1):262, January 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623158.

Anonymous:2007:FMa

[568] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1):??, ????

2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
4623129.

Martin:2007:ECP

[569] D. J. Martin. Editorial: Census present and future. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):263–266, March 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/4623161.

Taplin:2007:ESE

[570] Ross Taplin. Enhancing statistical education by using role-plays of consultations. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):267–300, March 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623162.

Handcock:2007:MBC

[571] Mark S. Handcock, Adrian E. Raftery, and Jeremy M. Tantrum. Model-based clustering for social networks. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2): 301–354, March 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623163.

Briggs:2007:CDI

[572] David Briggs, Daniela Fecht, and Kees De Hoogh. Census data issues for epidemiology and health risk assessment: Experiences from the small area health statistics unit. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):355–378, March 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        4623164.

Vickers:2007:CUN

[573] Dan Vickers and Phil Rees. Creating the UK national statistics 2001 output area classification. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):379–403, March 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        4623165.

Simpson:2007:GME

[574] Ludi Simpson. Ghettos of the mind: The empirical behaviour of indices of segregation and diversity. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2): 405–424, March 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623166.

Stillwell:2007:UUC

[575] John Stillwell and Oliver Duke-Williams. Understanding the 2001 UK census migration and commuting data: The effect of small cell adjustment and problems of comparison with 1991. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2): 425–445, March 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623167.

Champion:2007:UCS

[576] Tony Champion and Mike Coombes. Using the 2001 census to study human capital movements affecting Britain’s larger cities: Insights and issues. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2): 447–467, March 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623168.

Bartley:2007:ISM

[577] Mel Bartley and Ian Plewis. Increasing social mobility: An effective policy to reduce health inequalities. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2): 469–481, March 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623169.

Judson:2007:IIC

[578] D. H. Judson. Information integration for constructing social statistics: History, theory and ideas towards a research programme. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):483–501, March 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/4623170.

Anonymous:2007:BMb

[579] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):503, ???? 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
4623188.

Anonymous:2007:BRb

[580] Anonymous. Book reviews. Journal of the Royal Statistical Society.

Series A (Statistics in Society), 170 (2):503, March 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dialsingh:2007:BRBb

[581] Isaac Dialsingh. Book review: Sampling Methods: Exercises and Solutions, by P. Ardilly and Y. Tillé. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170 (2):504, March 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623171.

Ravi:2007:BRBa

[582] Sreenivasan Ravi. Book review: Probability and Random Variables, by G. P. Beaumont. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):504, March 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/4623172.

LeFevre:2007:BRB

[583] Andrea Le Fevre. Book review: Data Monitoring in Clinical Trials, by D. L. DeMets, C. D. Furberg and L. M. Friedman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):504–505, March 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623173.

Espejo:2007:BRBb

[584] Mariano Ruiz Espejo. Book review: Statistics for Real-life Sample Surveys, by S. Dorofeev and P. Grant. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):

505–506, March 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623174.

Morgan:2007:BRB

[585] Byron J. T. Morgan. Book review: A Kalman Filter Primer, by R. L. Eubank. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):506, March 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623175.

Hochkirchen:2007:BRB

[586] Thomas Hochkirchen. Book review: A Handbook of Statistical Analysis using R, by B. S. Everitt and T. Hothorn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170 (2):506–507, March 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623176.

Ugarte:2007:BRBa

[587] M. Dolores Ugarte. Book review: Design of Experiments for Agriculture and the Natural Sciences, 2nd edn, by A. R. Hoshmand. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):507, March 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/4623177.

Stoyanov:2007:BRBa

[588] Jordan Stoyanov. Book review: Probabilistic Symmetries and Invariance Principles, by O. Kallenberg. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):

507–508, March 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623178.

Ravi:2007:BRBb

[589] Sreenivasan Ravi. Book review: Testing Statistical Hypotheses, 3rd edn, by E. L. Lehmann and J. P. Romano. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):508, March 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623179.

Stoyanov:2007:BRBb

[590] Jordan Stoyanov. Book review: Introductory Stochastic Analysis for Finance and Insurance, by X. S. Lin. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2): 508–509, March 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623180.

Lazic:2007:BRBa

[591] S. E. Lazic. Book review: Correspondence Analysis and Data Coding with Java and R, by F. Murtagh. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170 (2):509, March 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623181.

Bakouch:2007:BRBb

[592] Hassan S. Bakouch. Book review: Analysis of Integrated and Cointegrated Time Series with R, by B. Pfaff. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2): 509–510, March 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623182.

Bingham:2007:BRB

[593] N. H. Bingham. Book review: Economic Capital and Financial Risk Management for Financial Services Firms and Conglomerates, by B. Porteous and P. Tapadar. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):510, March 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/4623183.

Shalabh:2007:BRB

[594] Shalabh. Book review: Theory of Preliminary Test and Stein-Type Estimation with Applications, by A. K. E. Saleh. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):510–511, March 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623184.

Stoimenova:2007:BRB

[595] Eugenia Stoimenova. Book review: Introduction to Nonparametric Regression, by K. Takezawa. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):511, March 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
4623185.

Espejo:2007:BRBc

[596] Mariano Ruiz Espejo. Book review:

Statistical Analysis of Cost-effectiveness Data, by A. R. Willan and A. H. Briggs. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170 (2):511–512, March 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623186.

Espejo:2007:BRBd

[597] Mariano Ruiz Espejo. Book review: WHO Child Growth Standards: Methods and Development, by World Health Organization. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):512, ???? 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/4623187.

Espejo:2007:BRW

[598] Mariano Ruiz Espejo. Book reviews: WHO child growth standards: Methods and development. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):512, March 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2007:FMb

[599] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):??, ???? 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
4623160.

Zeger:2007:EKI

[600] Scott L. Zeger. Editorial: Knowledge from information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):513–516, July 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
4623190.

Senn:2007:SIF

[601] Professor Stephen Senn, Dr Dipti Amin, Professor Rosemary A. Bailey, Professor Sheila M. Bird, Dr Barbara Bogacka, Mr Peter Colman, Dr Andrew Garrett, Professor Andrew Grieve, and Professor Sir Peter Lachmann. Statistical issues in first-in-man studies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170 (3):517–579, July 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623191.

Smith:2007:TSU

[602] T. M. F. Smith and L. Staetsky. The teaching of statistics in UK universities. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):581–622, July 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623192.

Brown:2007:ISE

[603] Giorgina Brown, John Micklewright, Sylke V. Schnepf, and Robert Waldmann. International surveys of educational achievement: How robust are the findings? Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):623–646, July 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/4623193.

Callegaro:2007:FDC

[604] Mario Callegaro, Charlotte Steeh, Trent D. Buskirk, Vasja Vehovar, Vesa Kuusela, and Linda Piekarski. Fitting disposition codes to mobile phone surveys: Experiences from studies in Finland, Slovenia and the USA. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170 (3):647–670, July 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623194.

Lennert-Cody:2007:SLP

[605] Cleridy E. Lennert-Cody and Richard A. Berk. Statistical learning procedures for monitoring regulatory compliance: An application to fisheries data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3): 671–689, July 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623195.

Ferguson:2007:MCC

[606] C. A. Ferguson, A. W. Bowman, E. M. Scott, and L. Carvalho. Model comparison for a complex ecological system. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170 (3):691–711, July 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623196.

Pont:2007:CNR

[607] Mark Pont. Coverage and non-response errors in the UK New Earnings Survey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):713–733, July 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623197.

Bokov:2007:TED

[608] Vladimir B. Bokov. Theoretic and empirical data-inclusive process characterization. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):735–758, July 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/4623198.

Boys:2007:AOJ

[609] Richard J. Boys, Peter J. Dolton, and Neil R. Walker. Adapting the optimal job search model for active labour market policy. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):759–779, July 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/4623199.

Aassve:2007:MDE

[610] Arnstein Aassve, Gianni Betti, Stefano Mazzuco, and Letizia Mencarini. Marital disruption and economic wellbeing: A comparative analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3): 781–799, July 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623200.

Steele:2007:ESR

[611] Fiona Steele, Anna Vignoles, and Andrew Jenkins. The effect of school resources on pupil attainment: A multilevel simultaneous equation modelling approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):801–824, July 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/4623201.

Goudie:2007:WCM

[612] I. B. J. Goudie and M. Goudie. Who captures the marks for the Petersen estimator? Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):825–839, July 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/4623202.

Abakuks:2007:MHT

[613] Andris Abakuks. A modification of honoré’s triple-link model in the synoptic problem. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):841–850, July 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/4623203.

Whitehead:2007:OWF

[614] F. E. Whitehead. Obituary: William Frederick Fenn Kemsley, MA, 1914– 2006. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):851–852, July 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623204.

Anonymous:2007:BMc

[615] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):853, ???? 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        4623213.

Anonymous:2007:BRc

[616] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):853, July 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Penny:2007:BRB

[617] Richard Penny. Book review: Survey Sampling: Theory and Methods, 2nd edn, by A. Chaudhuri; H. Stenger. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):854, July 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623205.

Dialsingh:2007:BRBc

[618] Isaac Dialsingh. Book review: Basic Statistics and Pharmaceutical Statistical Applications, 2nd edn, by J. E. De Muth. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):854–855, July 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623206.

Robinson:2007:BRBb

[619] Andrew Robinson. Book review: Generalized Linear Models with Random Effects: Unified Analysis via H-Likelihood, by Y. Lee; J. A. Nelder; Y. Pawitan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170 (3):855–856, July 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623207.

Robinson:2007:BRBc

[620] Andrew Robinson. Book review: Randomization, Bootstrap and Monte Carlo Methods in Biology, by B. F. J. Manly. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):856, July 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623208.

Karlsson:2007:BRBc

[621] Andreas Karlsson. Book review: Introduction to Linear Regression Analysis, by D. C. Montgomery; E. A. Peck; G. G. Vining. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):856–857, July 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/4623209.

Lu:2007:BRBa

[622] Z. Q. John Lu. Book review: Statistical Inference Based on Divergence Measures, by L. Pardo. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):857–858, July 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
4623210.

Congdon:2007:BRB

[623] Peter Congdon. Book review: Gaussian Markov Random Fields: Theory and Applications, by H. Rue; L. Held. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):858, July 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623211.

DAubigny:2007:BRB

[624] Gérard D’Aubigny. Book review: Data Analysis of Asymmetric Structures, by T. Saito; H. Yadohisa. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):858–860, July 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic). URL http://www.jstor.org/
stable/4623212.

Anonymous:2007:FMc

[625] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):??, ???? 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
4623189.

Molenberghs:2007:EWD

[626] Geert Molenberghs. Editorial: What to do with missing data? Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4): 861–863, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623216.

Ohlssen:2007:HMF

[627] David I. Ohlssen, Linda D. Sharples, and David J. Spiegelhalter. A hierarchical modelling framework for identifying unusual performance in health care providers. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):865–890, October 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/4623217.

Raymer:2007:CCR

[628] James Raymer, Guy Abel, and Peter W. F. Smith. Combining census and registration data to estimate detailed elderly migration flows in England and Wales. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):891–908, October 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623218.

Drakos:2007:SUR

[629] Konstantinos Drakos. The size of underreporting bias in recorded transnational terrorist activity. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):909–921, October 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
4623219.

An:2007:MIA

[630] Di An and Roderick J. A. Little. Multiple imputation: An alternative to top coding for statistical disclosure control. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4): 923–940, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623220.

Goldstein:2007:MEP

[631] Harvey Goldstein, Simon Burgess, and Brendon McConnell. Modelling the effect of pupil mobility on school differences in educational achievement. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4): 941–954, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623221.

Paas:2007:DTD

[632] Leonard J. Paas, Jeroen K. Vermunt, and Tammo H. A. Bijmolt. Discrete time, discrete state latent Markov modelling for assessing and predicting household acquisitions of financial products. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4): 955–974, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623222.

Molina:2007:SAE

[633] Isabel Molina, Ayoub Saei, and M. José Lombardía. Small area estimates of labour force participation under a multinomial logit mixed model. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4): 975–1000, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623223.

Brijs:2007:BMR

[634] Tom Brijs, Dimitris Karlis, Filip Van den Bossche, and Geert Wets. A Bayesian model for ranking hazardous road sites. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):1001–1017, October 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
4623224.

Haining:2007:CPP

[635] Robert Haining and Jane Law. Combining police perceptions with police records of serious crime areas: A modelling approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):1019–1034, October 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
4623225.

Afshartous:2007:ADS

[636] David Afshartous and Michael Wolf. Avoiding ‘data snooping’ in multilevel and mixed effects models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):1035– 1059, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623226.

Piccarreta:2007:CWF

[637] Raffaella Piccarreta and Francesco C. Billari. Clustering work and family trajectories, by using a divisive algorithm. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4): 1061–1078, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623227.

Anonymous:2007:RCS

[638] Anonymous. Report of the Council for the session 2006–2007. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4): 1079–1175, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623228.

Anonymous:2007:BMd

[639] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):1177, ???? 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
4623248.

Anonymous:2007:BRd

[640] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4): 1177, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Sprent:2007:BRB

[641] P. Sprent. Book review: An Introduction to Categorical Data Analysis, by A. Agresti. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):1178, October 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/4623229.

Lu:2007:BRBb

[642] Z. Q. John Lu. Book review: DNA Microarrays and Related Genomics Techniques: Design, Analysis, and Interpretation of Experiments, by D. B. Allison; G. P. Page; T. M. Beasley; J. W. Edwards. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):1178–1179, October 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/4623230.

Corbyn:2007:BRB

[643] Judith Corbyn. Book review: Essential Statistics for Public Managers and Policy Analysts; Exercising Essential Statistics Workbook, by E. M. Berman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170 (4):1179, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623231.

Wade:2007:BRB

[644] Angela Wade. Book review: Adaptive Design Methods in Clinical Trials, by S. C. Chow; M. Chang. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4): 1179–1180, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623232.

Lallmahomed:2007:BRB

[645] Naguib Lallmahomed. Book review: Statistique Théorique et Appliquée, by P. Dagnelie. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):1180, October 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/4623233.

Lewis:2007:BRB

[646] Steff Lewis. Book review: Statistical Analysis of Medical Data Using SAS, by G. Der; B. S. Everitt. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4): 1180–1181, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623234.

Lu:2007:BRBc

[647] Z. Q. John Lu. Book review: Bayesian Inference for Gene Expression and Proteomics, by K.-A. Do; P. Müller; M. Vannucci. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):1181, October 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/4623235.

Ugarte:2007:BRBb

[648] M. Dolores Ugarte. Book review: Statistical Methods for Spatio-Temporal Systems, by B. Finkenstädt; L. Held; V. Isham. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):1182, October 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/4623236.

Ravi:2007:BRBc

[649] Sreenivasan Ravi. Book review:

Bayesian Logical Data Analysis for the Physical Sciences: A Comparative Approach with Mathematica^{(}R) Support, by P. C. Gregory. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):1182–1183, October 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
4623237.

Gokcek:2007:BRB

[650] Güldem Gökçek. Book review: SPSS 14 Made Simple, by P. R. Kinnear; C. D. Gray. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):1183, ???? 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623238.

Gokcek:2007:BRS

[651] Güldem Gökçek. Book reviews: SPSS 14 made simple. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):1183, October 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Stoyanov:2007:BRBc

[652] Jordan Stoyanov. Book review: Probability and Random Processes, by V. Krishnan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):1183–1184, October 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/4623239.

Reed:2007:BRB

[653] Emma Reed. Book review: Starter Packs: A Strategy to Fight Hunger in Developing Countries?, by S. Levy. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170 (4):1184, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623240.

Peters:2007:BRB

[654] Gareth Peters. Book review: Uncertain Judgements-Eliciting Experts’ Probabilities, by A. O. Hagan; C. E. Buck; A. Daneshkhah; J. R. Eiser; P. Garthwaite; D. Jenkinson; J. Oakley; T. Rakow. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4): 1184–1185, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623241.

Lazic:2007:BRBb

[655] S. E. Lazic. Book review: Presenting Medical Statistics from Proposal to Publication, by J. Peacock; S. Kerry. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4): 1185–1186, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623242.

Bakouch:2007:BRBc

[656] Hassan S. Bakouch. Book review: Elements of Distribution Theory, by T. A. Severini. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):1186–1187, October 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/4623243.

Wright:2007:BRB

[657] L. F. Wright. Book review: Bayesian Networks and Probabilistic Inference in Forensic Science, by F. Taroni; C. Aitken; P. Garbolino; A. Biedermann. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170 (4):1187, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623244.

Gyles:2007:BRB

[658] Anthony F. Gyles. Book review: Asset Price Dynamics, Volatility, and Prediction, by S. J. Taylor. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):1187– 1189, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/4623245.

Lang:2007:BRB

[659] Iain A. Lang. Book review: Visual Statistics: Seeing Data with Dynamic Interactive Graphics, by F. W. Young; P. M. Valero-Mora; M. Friendly. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4): 1189, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623246.

Kumar:2007:BRB

[660] Kuldeep Kumar. Book review: Modelling Financial Time Series with S-PLUS, by E. Zivot; J. Wang. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4): 1189–1190, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623247.

Anonymous:2007:IAV

[661] Anonymous. Index of authors: volume 170, 2007. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):1191–1196, October 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2007:VI

[662] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4): 1191–1196, ???? 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623214.

Anonymous:2007:CVC

[663] Anonymous. Contents of volume 170, 2007: Contents of volume 170, 2007. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4): 1197–1198, October 2007. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2007:FMd

[664] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):??, ???? 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
4623215.

Hinde:2008:EAP

[665] Andrew Hinde. Editorial: The analysis of panel data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1):1–3, January 2008. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/30130727.

Steele:2008:MML

[666] Fiona Steele. Multilevel models for longitudinal data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1):5–19, January 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
30130728.

Pudney:2008:DPM

[667] Stephen Pudney. The dynamics of perception: Modelling subjective wellbeing in a short panel. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1):21–40, January 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
30130729.

Pohl:2008:MME

[668] Steffi Pohl, Rolf Steyer, and Katrin Kraus. Modelling method effects as individual causal effects. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1): 41–63, January 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130730.

Sturgis:2008:ESM

[669] Patrick Sturgis and Louise Sullivan. Exploring social mobility with latent trajectory groups. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1):65–88, January 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
30130731.

Berrington:2008:GCM

[670] Ann Berrington, Yongjian Hu, Peter W. F. Smith, and Patrick Sturgis. A graphical chain model for reciprocal relationships between women’s gender role attitudes and labour force participation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1): 89–108, January 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130732.

Kennedy:2008:QUB

[671] Marc Kennedy, Clive Anderson, Anthony O’Hagan, Mark Lomas, Ian Woodward, John Paul Gosling, and Andreas Heinemeyer. Quantifying uncertainty in the biospheric carbon flux for England and Wales. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1):109–135, January 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
30130733.

deRooij:2008:ACN

[672] Mark de Rooij. The analysis of change, Newton’s Law of Gravity and association models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1):137–157, January 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
30130734.

Jackson:2008:HRR

[673] Christopher Jackson, Nicky Best, and Sylvia Richardson. Hierarchical related regression for combining aggregate and individual data in studies of socioeconomic disease risk factors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1): 159–178, January 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130735.

Glynn:2008:ALE

[674] Adam N. Glynn, Jon Wakefield, Mark S.

Handcock, and Thomas S. Richardson. Alleviating linear ecological bias and optimal design with subsample data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1): 179–202, January 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130736.

Khoman:2008:IBE

[675] Ehsan Khoman, James Mitchell, and Martin Weale. Incidence-based estimates of life expectancy of the healthy for the UK: Coherence between transition probabilities and aggregate lifetables. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1):203–222, January 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30130737.

Woodall:2008:UEP

[676] William H. Woodall, J. Brooke Marshall, Michael D. Joner, Jr., Shannon E. Fraker, and Abdel-Salam G. Abdel-Salam. On the use and evaluation of prospective scan methods for health-related surveillance. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1): 223–237, January 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30130738.

Kynn:2008:HBB

[677] Mary Kynn. The ‘heuristics and biases’ bias in expert elicitation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1): 239–264, January 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30130739.

Bijleveld:2008:MBM

[678] Frits Bijleveld, Jacques Commandeur, Phillip Gould, and Siem Jan Koopman. Model-based measurement of latent risk in time series with applications. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1): 265–277, January 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130740.

Richards:2008:DYB

[679] S. J. Richards. Detecting year-of-birth mortality patterns with limited data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1): 279–298, January 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130741.

Garre:2008:JLC

[680] Francisca Galindo Garre, Aeilko H.

Zwinderman, Ronald B. Geskus, and Yvo W. J. Sijpkens. A joint latent class changepoint model to improve the prediction of time to graft failure. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1): 299–308, January 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130742.

Galbraith:2008:ONW

[681] J. I. Galbraith. Obituary: Neil William Please, 1919–2005. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1):309, January 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
30130743.

Anonymous:2008:BRa

[682] Anonymous. Book reviews. Journal of the Royal Statistical Society.

Series A (Statistics in Society), 171 (1):311, January 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2008:LRa

[683] Anonymous. [list of reviews]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1):311, ???? 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        30130744.

Lallmahomed:2008:BRBa

[684] Naguib Lallmahomed. Book review: Probability and Statistics for Computer Scientists, by M. Baron. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1): 312, January 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30130745.

Ginestet:2008:BRBa

[685] Cedric E. Ginestet. Book review: Latent Curve Models: A Structural Equation Perspective, by K. A. Bollen; P. J. Curran. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1):312–313, January 2008. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/30130746.

Castagliola:2008:BRB

[686] Philippe Castagliola. Book review: Response Surfaces, Mixtures, and Ridge Analyses, by G. E. P. Box; N. R. Draper. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171 (1):313, January 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130747.

Lark:2008:BRB

[687] R. M. Lark. Book review: Compositional Data Analysis in the Geosciences: From Theory to Practice, by A. Buccianti; G. Mateu-Figueras; V. Pawlowsky-Glahn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1): 313–314, January 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130748.

Wade:2008:BRBa

[688] Angela Wade. Book review: Adaptive Design Methods in Clinical Trials, by S.-C. Chow; M. Chang. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171 (1):314, January 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130749.

Putcha:2008:BRBa

[689] Venkata Putcha. Book review: A Crash Course in SPSS for Windows: Updated for Versions 10, 11, 12 and 13, by A. Colman; B. Pulford. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1): 314–315, January 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130750.

Kimber:2008:BRB

[690] Alan Kimber. Book review: Principles of Statistical Inference, by D. R. Cox. Journal of the Royal Statistical Society.

Series A (Statistics in Society), 171(1): 315–316, January 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30130751.

Kibria:2008:BRB

[691] B. M. Golam Kibria. Book review: Bayes Linear Statistics: Theory and Methods, by M. Goldstein; D. Wooff. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171 (1):316, January 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30130752.

Putcha:2008:BRBb

[692] Venkata Putcha. Book review: Handbook of Univariate and Multivariate Data Analysis and Interpretation with SPSS, by R. Ho. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1):317, January 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        30130753.

Stoyanov:2008:BRBa

[693] Jordan Stoyanov. Book review: Binomial Models in Finance, by J. van der Hoek; R. J. Elliott. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1):317–318, January 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        30130754.

Reese:2008:BRB

[694] R. Allan Reese. Book review: Linear Mixed Models: A Practical Guide Using Statistical Software, by B. T. West; K. B. Welsh; A. T. Gałecki. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171 (1):318, January 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130755.

Anonymous:2008:BMa

[695] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1):??, ???? 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
30130756.

Anonymous:2008:FMa

[696] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(1):??, ???? 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
30130726.

Gneiting:2008:EPF

[697] Tilmann Gneiting. Editorial: Probabilistic forecasting. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2):319–321, April 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
30130759.

Holt:2008:OSP

[698] D. Tim Holt. Official statistics, public policy and public trust. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2): 323–346, April 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30130760.

Jansen:2008:FMM

[699] Ivy Jansen and Geert Molenberghs. A flexible marginal modelling strategy for non-monotone missing data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2): 347–373, April 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30130761.

Pardoe:2008:ADC

[700] Iain Pardoe and Dean K. Simonton. Applying discrete choice models to predict Academy Award winners. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2): 375–394, April 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30130762.

Choo:2008:NAI

[701] Louise Choo and Stephen G. Walker. A new approach to investigating spatial variations of disease. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2): 395–405, April 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30130763.

Chang:2008:CSS

[702] Tee-Chin Chang. Cumulative sum schemes for surgical performance monitoring. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2):407–432, April 2008. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/30130764.

Cooper:2008:RMS

[703] W. W. Cooper, Subhash C. Ray, and M. Stone. A response to M. Stone: ‘How Not to Measure the Efficiency of Public Services (And How One Might)’ [with reply]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2):433–448, April 2008. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/30130765. See

[120].

Bollaerts:2008:ISH

[704] Kaatje Bollaerts, Marc Aerts, Stefaan Ribbens, Yves Van der Stede, Ides Boone, and Koen Mintiens. Identification of salmonella high risk pig-herds in Belgium, by using semiparametric quantile regression. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2):449–464, April 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
30130766.

Michael:2008:DPD

[705] Robert T. Michael and Colm A.

O’Muircheartaigh. Design priorities and disciplinary perspectives: The case of the US national children’s study. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2): 465–480, April 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130767.

Imai:2008:MBE

[706] Kosuke Imai, Gary King, and Elizabeth A. Stuart. Misunderstandings between experimentalists and observationalists about causal inference. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2): 481–502, April 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30130768.

Anonymous:2008:BRb

[707] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2):503, April 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2008:LRb

[708] Anonymous. [list of reviews]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2):503, ???? 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        30130769.

Lu:2008:BRB

[709] Z. Q. John Lu. Book review: Optimum Experimental Designs, with SAS, by A. C. Atkinson; A. N. Donev; R. D. Tobias. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2):504, April 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30130770.

Robinson:2008:BRB

[710] Andrew Robinson. Book review: Applied Asymptotics: Case Studies in Small-Sample Statistics, by A. R. Brazzale; A. C. Davison; N. Reid. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2): 504–505, April 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130771.

Roddam:2008:BRB

[711] Andrew W. Roddam. Book review:

Measurement Error in Nonlinear Models: A Modern Perspective, by R. J. Carroll; D. Ruppert; L. A. Stefanski; C. M. Crainiceanu. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2):505–506, April 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
30130772.

Ginestet:2008:BRBb

[712] Cedric E. Ginestet. Book review: Combinatorics, Complexity, and Chance: A Tribute to Dominic Welsh, by G. Grimmett; C. McDiarmid. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2):506, April 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
30130773.

Gurka:2008:BRB

[713] Matthew J. Gurka. Book review: Basic Concepts in Biostatistics and Epidemiology, by T. H. MacDonald. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2): 506–507, April 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130774.

Muthami:2008:BRB

[714] Lawrence Muthami. Book review: Statistical Monitoring of Clinical Trials, by L. A. Moye. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2):507, April 2008. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/30130775.

Seuc:2008:BRB

[715] Armando H. Seuc. Book review: Approaching Multivariate Analysis: An Introduction for Psychology, by J. Todman; P. Dugard. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2):507–508, April 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        30130776.

Anonymous:2008:BMb

[716] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2):??, ???? 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        30130777.

Anonymous:2008:FMb

[717] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2):??, ???? 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        30130758.

Curtice:2008:EPC

[718] John Curtice and David Firth. Exit polling in a cold climate: The BBC–ITV experience in Britain in 2005 [with discussion]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3):509–539, June 2008. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/30135085.

Goubar:2008:EHI

[719] A. Goubar, A. E. Ades, D. De Angelis, C. A. McGarrigle, C. H. Mercer, P. A. Tookey, K. Fenton, and O. N. Gill. Estimates of human immunodeficiency virus prevalence and proportion diagnosed based on Bayesian multiparameter synthesis of surveillance data [with discussion]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3):541–580, June 2008. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/30135086.

vandenBrakel:2008:DBA

[720] Jan A. van den Brakel. Design-based analysis of embedded experiments with applications in the Dutch Labour Force Survey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3):581–613, June 2008. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/30135087.

Brown:2008:HDF

[721] Sarah Brown and Karl Taylor. Household debt and financial assets: Evidence from Germany, Great Britain and the USA. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3):615–643, June 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30135088.

Elliott:2008:IWD

[722] Robert J. R. Elliott and Joanne K. Lindley. Immigrant wage differentials, ethnicity and occupational segregation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171 (3):645–671, June 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30135089.

Andrews:2008:HWW

[723] M. J. Andrews, L. Gill, T. Schank, and R. Upward. High wage workers and low wage firms: Negative assortative matching or limited mobility bias? Journal of the Royal Statistical Society. Series A (Statistics in Society), 171 (3):673–697, June 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30135090.

Garabed:2008:EAB

[724] R. B. Garabed, W. O. Johnson, J. Gill, A. M. Perez, and M. C. Thurmond. Exploration of associations between governance and economics and country level foot-and-mouth disease status by using Bayesian model averaging. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3): 699–722, June 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30135091.

Kreuter:2008:GIB

[725] Frauke Kreuter, Ting Yan, and Roger Tourangeau. Good item or bad — can latent class analysis tell?: The utility of latent class analysis for the evaluation of survey questions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3): 723–738, June 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30135092.

Rijmen:2008:QLA

[726] Frank Rijmen, Edward H. Ip, Stephen Rapp, and Edward G. Shaw. Qualitative longitudinal analysis of symptoms in patients with primary and metastatic brain tumours. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3):739–753, June 2008. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/30135093.

Anonymous:2008:BRc

[727] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3):755, June 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2008:LRc

[728] Anonymous. [list of reviews]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3):755, ???? 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
30135094.

White:2008:BRB

[729] Ian White. Book review: Sample Size Calculations in Clinical Research, by S.- C. Chow; J. Shao; H. Wang. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3):756, June 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
30135095.

Tusell:2008:BRB

[730] Fernando Tusell. Book review: An Introduction to State Space Time Series Analysis, by J. J. F. Commandeur; S. J. Koopman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3):756–757, June 2008. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/30135096.

Karlsson:2008:BRB

[731] Andreas Karlsson. Book review: Randomization Tests, by E. S. Edgington; P. Onghena. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3):757–758, June 2008. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/30135097.

Anonymous:2008:BRBa

[732] Anonymous. Book review: Statistical Test Theory for the Behavioral Sciences. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3):758, June 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shalabh:2008:BRBa

[733] Shalabh. Book review: Statistical Test Theory for the Behavioral Sciences, by D. N. M. de Gruijter; L. J. Th. van der Kamp. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3):758, ???? 2008. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/30135098.

McElduff:2008:BRB

[734] Fiona McElduff. Book review: Negative Binomial Regression, by J. M. Hilbe. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171 (3):758–759, June 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30135099.

Burrell:2008:BRB

[735] Quentin L. Burrell. Book review: Handbook of Exponential and Related Distributions for Engineers and Scientists, by N. Pal; C. Jin; W. K. Lim. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3): 759–760, June 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30135100.

Berry:2008:BRB

[736] Diane Berry. Book review: A Handbook of Statistical Analyses Using Stata, by S. Rabe-Hesketh; B. S. Everitt. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3):760, June 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
30135101.

Wade:2008:BRBb

[737] Angie Wade. Book review: Matched Sampling for Causal Effects, by D. B. Rubin. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3):760–761, June 2008. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/30135102.

Lallmahomed:2008:BRBb

[738] Naguib Lallmahomed. Book review: Elementary Statistics Using JMP, by S. Schlotzhauer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3):761, June 2008. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/30135103.

Kumar:2008:BRB

[739] Kuldeep Kumar. Book review: Introduction to Bioinformatics, by A. Tramontano. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3):761–762, June 2008. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/30135104.

Anonymous:2008:BMc

[740] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3):??, ???? 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        30135105.

Anonymous:2008:FMc

[741] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(3):??, ???? 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        30135084.

Butt:2008:CCL

[742] Zoltan Butt, Steven Haberman, Richard Verrall, and Victoria Wass. Calculating compensation for loss of future earnings: Estimating and using work life expectancy [with discussion]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4): 763–805, October 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130781.

Welton:2008:RPB

[743] N. J. Welton, A. E. Ades, D. M. Caldwell, and T. J. Peters. Research prioritization based on expected value of partial perfect information: A case-study on interventions to increase uptake of breast cancer screening [with discussion]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4): 807–841, October 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130782.

LeGrand:2008:GEC

[744] Julian Le Grand. The giants of excess:

A challenge to the nation’s health. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4): 843–856, October 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130783.

Gregg:2008:TSE

[745] Paul Gregg and Jonathan Wadsworth.

Two sides to every story: Measuring polarization and inequality in the distribution of work. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4):857–875, October 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        30130784.

Reboussin:2008:LDL

[746] Beth A. Reboussin, Edward H. Ip, and Mark Wolfson. Locally dependent latent class models with covariates: An application to under-age drinking in the USA. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4):877–897, October 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30130785.

Nason:2008:EIC

[747] Guy P. Nason and Daniel Bailey. Estimating the intensity of conflict in Iraq. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4): 899–914, October 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30130786.

Presanis:2008:CEB

[748] A. M. Presanis, D. De Angelis, D. J. Spiegelhalter, S. Seaman, A. Goubar, and A. E. Ades. Conflicting evidence in a Bayesian synthesis of surveillance data to estimate human immunodeficiency virus prevalence. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4):915–937, October 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        30130787.

Cooper:2008:UUS

[749] D. Cooper, W. D. McCausland, and I. Theodossiou. Unemployed, uneducated and sick: The effects of socio-economic status on health duration in the European Union. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4): 939–952, October 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130788.

Grimes:2008:RCJ

[750] Tom Grimes, David Hand, and Sayer Vincent. Report of the Council for July– December 2007. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4):953–1028, October 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
30130789.

Lauritzen:2008:OPA

[751] Steffen Lauritzen. Obituary: Professor Anders Hald, 1913–2007. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4): 1029–1030, October 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130790.

Anonymous:2008:BRd

[752] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4): 1031, October 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2008:LRd

[753] Anonymous. [list of reviews]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4): 1031, October 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130791.

Stoyanov:2008:BRBb

[754] Jordan Stoyanov. Book review: Random Dynamical Systems: Theory and Applications, by R. Bhattacharya; M. Majumdar. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4):1032, October 2008. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/30130792.

Hill:2008:BRB

[755] I. D. Hill. Book review: Mathematics and Democracy: Designing Better Voting and Fair-Division Procedures, by S. J. Brams. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4):1032–1033, October 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
30130793.

Stoyanov:2008:BRBc

[756] Jordan Stoyanov. Book review: Saddlepoint Approximations with Applications, by R. W. Butler. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4):1033–1034, October 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
30130794.

Corbyn:2008:BRBa

[757] Judith Corbyn. Book review: Poverty, Wealth and Place in Britain, 1968–2005, by D. Dorling; J. Rigby; B. Wheeler; D. Ballas; B. Thomas; E. Fahmy; D. Gordon; R. Lupton. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4):1034–1035, October 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
30130795.

Lula:2008:BRB

[758] Usman I. Lula. Book review: Companion to Advanced Mathematics and Statistics, by S. Dudzic. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4):1035– 1036, October 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130796.

Karagrigoriou:2008:BRB

[759] Alex Karagrigoriou. Book review:

Analysing Receiver Operating Characteristic Curves with SAS, by M. Gönen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171 (4):1036, October 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130797.

Longford:2008:BRB

[760] N. T. Longford. Book review: Bioequivalence Studies in Drug Development: Methods and Applications, by D. Hauschke; V. Steinijans; I. Pigeot. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171 (4):1037, October 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130798.

Clark:2008:BRB

[761] Stephen Clark. Book review: Placing History: How Maps, Spatial Data, and GIS Are Changing Historical Scholarship, by A. K. Knowles; A. Hillier. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4): 1037–1038, October 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30130799.

Hochkirchen:2008:BRB

[762] Thomas Hochkirchen. Book review: Handbook of Statistical Distributions with Applications, by K. Krishnamoorthy. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4):1038–1039, October 2008. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/30130800.

Naguib:2008:BRB

[763] Lallmahomed Naguib. Book review: Introduction to Statistics for Biology, 3rd edn, by R. H. McCleery; T. A. Watt; T. Hart. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4):1039, October 2008. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/30130801.

Shalabh:2008:BRBb

[764] Shalabh. Book review: Missing Data in Clinical Studies, by G. Molenberghs; M. G. Kenward. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4):1039–1040, October 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        30130802.

Vaughan:2008:BRB

[765] Ian Vaughan. Book review: Niche Modeling: Predictions from Statistical Distributions, by D. R. B. Stockwell. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4): 1040–1041, October 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130803.

Corbyn:2008:BRBb

[766] Judith Corbyn. Book review: Identity in Britain: A Cradle-to-Grave Atlas, by B. Thomas; D. Dorling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4): 1041–1042, October 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30130804.

Anonymous:2008:CHM

[767] Anonymous. Corrigendum: How not to measure the efficiency of public services (and how one might). Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4):1043, October 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic). URL http://www.jstor.org/
stable/30130805. See [120].

Anonymous:2008:IAV

[768] Anonymous. Index of authors: volume 171, 2008. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4):1045–1050, October 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2008:BMd

[769] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4):??, October 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        30130806.

Anonymous:2008:FMd

[770] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4):??, October 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        30130780.

Anonymous:2008:VI

[771] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171 (4):??, October 2008. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30130779.

Fielding:2009:EBF

[772] Antony Fielding and Simon Day. Editorial: Built on firm foundations. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (1):1–2, January 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30136737.

Lipsitz:2009:JGE

[773] Stuart R. Lipsitz, Garrett M. Fitzmaurice, Joseph G. Ibrahim, Debajyoti Sinha, Michael Parzen, and Steven Lipshultz. Joint generalized estimating equations for multivariate longitudinal binary outcomes with missing data: An application to acquired immune deficiency syndrome data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(1): 3–20, January 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30136738.

Turner:2009:BME

[774] Rebecca M. Turner, David J. Spiegelhalter, Gordon C. S. Smith, and Simon G. Thompson. Bias modelling in evidence synthesis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(1):21–47, January 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/30136739.

Grigg:2009:LMC

[775] O. A. Grigg, D. J. Spiegelhalter, and H. E. Jones. Local and marginal control charts applied to methicillin resistant Staphylococcus aureus bacteraemia reports in UK Acute National Health Service Trusts. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(1):49–66, January 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/30136740.

Greiner:2009:EIB

[776] D. James Greiner and Kevin M. Quinn.

R × C ecological inference: Bounds, correlations, flexibility and transparency of assumptions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(1):67–81, January 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        30136741.

Tebaldi:2009:JPT

[777] Claudia Tebaldi and Bruno Sansó. Joint projections of temperature and precipitation change from multiple climate models: A hierarchical Bayesian approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(1):83–106, January 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/30136742.

Mills:2009:FOT

[778] Terence C. Mills. Forecasting obesity trends in England. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(1):107–117, January 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        30136743.

Welton:2009:MPB

[779] N. J. Welton, A. E. Ades, J. B. Carlin, D. G. Altman, and J. A. C. Sterne. Models for potentially biased evidence in meta-analysis using empirically based priors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(1):119–136, January 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30136744.

Higgins:2009:RER

[780] Julian P. T. Higgins, Simon G. Thompson, and David J. Spiegelhalter. A re-evaluation of random-effects metaanalysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(1):137–159, January 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/30136745.

Bevan:2009:HMT

[781] Gwyn Bevan and Richard Hamblin. Hitting and missing targets, by ambulance services for emergency calls: Effects of different systems of performance measurement within the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(1): 161–190, January 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30136746.

Berk:2009:FMW

[782] Richard Berk, Lawrence Sherman, Geoffrey Barnes, Ellen Kurtz, and Lindsay Ahlman. Forecasting murder within a population of probationers and parolees: A high stakes application of statistical learning. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(1):191–211, January 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/30136747.

Hocine:2009:SCS

[783] Mounia N. Hocine, Patrick Musonda, Nick J. Andrews, and C. Paddy Farrington. Sequential case series analysis for pharmaco-vigilance. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(1): 213–236, January 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30136748.

Laenen:2009:FME

[784] Annouschka Laenen, Ariel Alonso, Geert Molenberghs, and Tony Vangeneugden. A family of measures to evaluate scale reliability in a longitudinal setting. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(1): 237–253, January 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30136749.

Mazzuco:2009:ALE

[785] Stefano Mazzuco. Another look into the effect of premarital cohabitation on duration of marriage: An approach based on matching. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(1):255–273, January 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        30136750.

Grimmett:2009:ODG

[786] Geoffrey Grimmett. Obituary: David George Kendall, 1918–2007. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(1): 275–278, January 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30136751.

Gani:2009:OCC

[787] Joe Gani. Obituary: C. C. Heyde, 1939– 2008. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(1):278–280, January 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30136752.

Anonymous:2009:BRa

[788] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (1):281, January 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2009:LR

[789] Anonymous. List of reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (1):281, January 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30136753.

Hochkirchen:2009:BRB

[790] Thomas Hochkirchen and Aachen. Book review: Design and Analysis of Experiments, vol. I, Introduction to Experimental Designs, by K. Hinkelmann; O. Kempthorne. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(1):282, January 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/30136754.

Woods:2009:BRB

[791] David Woods. Book review: Computational Pharmacokinetics, by A. Kallen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(1): 282–283, January 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30136755.

Bakouch:2009:BRB

[792] Hassan S. Bakouch. Book review:

Time Series Analysis, by H. Madsen.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (1):283, January 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30136756.

Patel:2009:BRB

[793] Deven D. Patel. Book review: Data Analysis and Graphics Using R: An Example-Based Approach, by J. Maindonald; W. J. Braun. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(1):283– 284, January 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30136757.

McElduff:2009:BRB

[794] Fiona McElduff. Book review: Digital Dice: Computational Solutions to Practical Probability Problems, by P. J. Nahin. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(1):284, January 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30136758.

Espejo:2009:BRB

[795] Mariano Ruiz Espejo. Book review: Putting People and Health Needs on the Map, by K. O’Neill; J. P. Meert. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (1):285, January 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30136759.

Anonymous:2009:BMa

[796] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (1):??, January 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30161796.

Anonymous:2009:FMa

[797] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (1):??, January 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/30136736.

Hand:2009:MSM

[798] David J. Hand. Modern statistics:

The myth and the magic. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(2): 287–306, April 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622500.

Chevalier:2009:OSU

[799] Arnaud Chevalier and Joanne Lindley.

Overeducation and the skills of UK graduates. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(2):307–337, April 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622501.

McKenzie:2009:SMH

[800] David J. McKenzie and Johan Mistiaen. Surveying migrant households:

A comparison of census-based, snowball and intercept point surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(2): 339–360, April 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622502.

Durrant:2009:MMR

[801] Gabriele B. Durrant and Fiona Steele. Multilevel modelling of refusal and noncontact in household surveys: Evidence from six UK government surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(2): 361–381, April 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622503.

Jackson:2009:AUH

[802] Christopher H. Jackson, Simon G. Thompson, and Linda D. Sharples. Accounting for uncertainty in health economic decision models, by using model averaging. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(2):383–404, April 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/20622504.

Yu:2009:MPB

[803] Binbing Yu, Lan Huang, Ram C. Tiwari, Eric J. Feuer, and Karen A. Johnson. Modelling population-based cancer survival trends by using join point models for grouped survival data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(2): 405–425, April 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622505.

Belot:2009:FTG

[804] Michèle Belot and John Ermisch.

Friendship ties and geographical mobility: Evidence from Great Britain. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (2):427–442, April 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622506.

Rosenblum:2009:ADE

[805] Michael Rosenblum, Nicholas P. Jewell, Mark van der Laan, Stephen Shiboski, Ariane van der Straten, and Nancy Padian. Analysing direct effects in randomized trials with secondary interventions: An application to human immunodeficiency virus prevention trials. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (2):443–465, April 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622507.

Bartholomew:2009:SGT

[806] David J. Bartholomew, Ian J. Deary, and Martin Lawn. Sir Godfrey Thomson: A statistical pioneer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(2): 467–482, April 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622508.

Cox:2009:SOC

[807] D. R. Cox, Michelle Jackson, and Shiwei Lu. On square ordinal contingency tables: A comparison of social class and income mobility for the same individuals. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (2):483–493, April 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/20622509.

Olives:2009:CDA

[808] Casey Olives, Marcello Pagano, Megan Deitchler, Bethany L. Hedt, Kari Egge, and Joseph J. Valadez. Cluster designs to assess the prevalence of acute malnutrition by lot quality assurance sampling: A validation study by computer simulation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(2):495–510, April 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/20622510.

Kohnen:2009:MIC

[809] Christine N. Kohnen and Jerome P. Reiter. Multiple imputation for combining confidential data owned by two agencies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(2):511–528, April 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/20622511.

Anonymous:2009:BRb

[810] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(2):529, April 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Adams:2009:BRB

[811] Niall Adams. Book review: Semi-Supervised Learning, by O. Chappelle; B. Schölkopf; A. Zien. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(2):530, April 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
20622513.

Ravi:2009:BRBa

[812] Sreenivasan Ravi. Book review: Statistical and Probabilistic Methods in Actuarial Science, by P. J. Boland. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(2):530, April 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
20622512.

Karagrigoriou:2009:BRB

[813] Alex Karagrigoriou. Book review: Basic Statistics Using SAS Enterprise Guide, a Primer, by G. Der; B. S. Everitt. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (2):530–531, April 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622514.

Koutoumanou:2009:BRB

[814] Eirini Koutoumanou. Book review:

Medical Biostatistics, by A. Indrayan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (2):531–532, April 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622515.

Shalabh:2009:BR

[815] Shalabh. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(2):532, April 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shalabh:2009:BRBa

[816] Shalabh. Book review: A Primer on Linear Models, by J. F. Monahan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(2):532, April 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/20622516.

Woodward:2009:BRBa

[817] Phil Woodward. Book review: Statistical Issues in Drug Development, by S. Senn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(2):532–533, April 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/20622517.

Brooms:2009:BRB

[818] A. C. Brooms. Book review: Experimental Statistics Using Minitab, by C. Weatherup. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(2):533–534, April 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/20622518.

Anonymous:2009:BMb

[819] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(2):??, April 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
20622519.

Anonymous:2009:FMb

[820] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(2):??, April 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
20622499.

Browne:2009:ERA

[821] William J. Browne and Fiona Steele.

Editorial: Recent advances in multilevel modelling methodology and applications. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(3):535–536, June 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/20622522.

Leckie:2009:CSN

[822] George Leckie. The complexity of school and neighbourhood effects and movements of pupils on school differences in models of educational achievement. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (3):537–554, June 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622523.

Leyland:2009:EAR

[823] Alastair H. Leyland and Øyvind Næss.

The effect of area of residence over the life course on subsequent mortality. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (3):555–578, June 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622524.

Browne:2009:USR

[824] William J. Browne, Fiona Steele, Mousa Golalizadeh, and Martin J. Green. The use of simple reparameterizations to improve the efficiency of Markov chain Monte Carlo estimation for multilevel models with applications to discrete time survival models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(3): 579–598, June 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622525.

Goldstein:2009:MMM

[825] Harvey Goldstein and Daphne Kounali. Multilevel multivariate modelling of childhood growth, numbers of growth measurements and adult characteristics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (3):599–613, June 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622526.

Molitor:2009:UBG

[826] Nuoo-Ting Molitor, Nicky Best, Chris Jackson, and Sylvia Richardson. Using Bayesian graphical models to model biases in observational studies and to combine multiple sources of data: Application to low birth weight and water disinfection by-products. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(3): 615–637, June 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622527.

Muthen:2009:MRM

[827] Bengt Muthén and Tihomir Asparouhov. Multilevel regression mixture analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(3):639–657, June 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/20622528.

Skrondal:2009:PMG

[828] Anders Skrondal and Sophia Rabe-Hesketh. Prediction in multilevel generalized linear models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(3):659–687, June 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic). URL http://www.jstor.org/
stable/20622529.

Armitage:2009:ONT

[829] P. Armitage and J. F. Bithell. Obituaries: Norman Thomas John Bailey, 1923–2007. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(3):689–690, June 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/20622530.

Possolo:2009:OJM

[830] Antonio Possolo. Obituaries: John Mandel, 1914–2007. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(3):691–692, June 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
20622531.

Anonymous:2009:BRc

[831] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(3):693, June 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Ginestet:2009:BRBa

[832] Cedric Ginestet. Book review: Semisupervised Learning for Computational Linguistics, by S. Abney. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(3):694, June 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        20622532.

Leckie:2009:BRB

[833] George Leckie. Book review: Handbook of Multilevel Analysis, by J. De Leeuw; E. Meijer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(3):694–695, June 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/20622533.

Borja:2009:BRBa

[834] Mario Cortina Borja. Book review: An Introduction to Generalized Linear Models, by A. J. Dobson; A. G. Barnett. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (3):695–696, June 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/20622534.

Berry:2009:BRB

[835] Diane Berry. Book review: R Programming for Bioinformatics, by R. Gentleman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(3):696, June 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622535.

Stoyanov:2009:BRB

[836] Jordan Stoyanov. Book review: Introduction to Stochastic Calculus Applied to Finance, by D. Lamberton; B. Lapeyere. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(3):697, June 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622536.

Reese:2009:BRBa

[837] R. Allan Reese. Book review: R for SAS and SPSS Users, by R. A. Muenchen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (3):697–698, June 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622537.

Wright:2009:BRB

[838] Lesley F. Wright. Book review: Reliability and Risk — a Bayesian Perspective, by N. D. Singpurwalla. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(3):698, June 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
20622538.

Borja:2009:BRBb

[839] Mario Cortina Borja. Book review:

Data Manipulation in R, by P. Spector. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(3):699, June 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/20622539.

Lazic:2009:BRB

[840] S. E. Lazic. Book review: DNA Methylation Microarrays: Experimental Design and Statistical Analysis, by S.-C. Wang; A. Petronis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(3):699–700, June 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/20622540.

Reese:2009:BRBb

[841] R. Allan Reese. Book review: Geostatistics for Environmental Scientists, by R. Webster; M. A. Oliver. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(3):700, June 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        20622541.

Anonymous:2009:BMc

[842] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(3):??, June 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        20622542.

Anonymous:2009:FMc

[843] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(3):??, June 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        20622521.

Lutz:2009:ETW

[844] Wolfgang Lutz. Editorial: Towards a world of 2–6 billion well-educated and therefore healthy and wealthy people. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4): 701–705, October 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622546.

Lancaster:2009:SIA

[845] Gillian A. Lancaster. Statistical issues in the assessment of health outcomes in children: A methodological review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4): 707–727, October 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622547.

Baker:2009:URU

[846] Stuart G. Baker, Nancy R. Cook, Andrew Vickers, and Barnett S. Kramer. Using relative utility curves to evaluate risk prediction. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):729–748, October 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
20622548.

Ermisch:2009:MPT

[847] John Ermisch, Diego Gambetta, Heather Laurie, Thomas Siedler, and S. C. Noah Uhrig. Measuring people’s trust. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4): 749–769, October 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/20622549.

Keogh:2009:CAC

[848] Gerard Keogh and Patrick Quill. The construction and analysis of a consistent set of input-output tables for the Irish economy. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):771–788, October 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/20622550.

Riley:2009:MMA

[849] Richard D. Riley. Multivariate metaanalysis: The effect of ignoring withinstudy correlation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):789–811, October 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        20622551.

Baker:2009:MPD

[850] R. D. Baker and I. G. McHale. Modelling the probability distribution of prize winnings in the UK national lottery: Consequences of conscious selection. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):813–834, October 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/20622552.

Leckie:2009:LUS

[851] George Leckie and Harvey Goldstein. The limitations of using school league tables to inform school choice. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):

835–851, October 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622553.

Hand:2009:RC

[852] David Hand, Martin Dougherty, Sarah Simpson, Sheila Bird, Andrew Garratt, Colin Aitken, Paul Gentry, Adrian Bowman, Nicola Emmerson, Andy Garrett, Gerald Goodall, Susan Starkings, Debra Hurcomb, Robin Henderson, Martin Owen, Rob Henderson, Keith Dugmore, and Jenny Church. Report of the Council for 2008. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):853–931, October 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
20622554.

Nathan:2009:OHS

[853] Gad Nathan. Obituary: Hendrik Salomon Konijn, 1918–2008. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):933, October 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
20622555.

Anonymous:2009:BRd

[854] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):934, October 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shalabh:2009:BRBb

[855] Shalabh. Book review: Statistical Learning from a Regression Perspective, by R. A. Berk. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):935, October 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/20622556.

Woodward:2009:BRBb

[856] Phil Woodward. Book review: Bayesian Methods for Data Analysis, by B. P. Carlin; T. A. Louis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):935–936, October 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        20622557.

Shalabh:2009:BRBc

[857] Shalabh. Book review: Statistical Design, by G. Casella. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):936, October 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        20622558.

Ginestet:2009:BRBb

[858] Cedric E. Ginestet. Book review: Model Selection and Model Averaging, by G. Claeskens; N. L. Hjort. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):937, October 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        20622559.

Lu:2009:BRB

[859] Z. Q. John Lu. Book review: Morphometrics with R, by J. Claude. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):

937–938, October 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622560.

Kumar:2009:BRB

[860] Kuldeep Kumar. Book review: Statistical Practice in Business and Industry, by S. Coleman; T. Greenfield; D. Stewardson; D. D. Montgomery. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4): 938–939, October 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622561.

Robinson:2009:BRB

[861] Derek Robinson. Book review: Introductory Statistics with R, by P. Dalgaard. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):939, October 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622562.

Baltagi:2009:BRB

[862] Badi H. Baltagi. Book review: Longitudinal Data Analysis, by G. Fitzmaurice; M. Davidian; G. Verbeke; G. Molenberghs. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):939–940, October 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/20622563.

Ravi:2009:BRBb

[863] Sreenivasan Ravi. Book review: Nonlinear Time Series: Semiparametric and Nonparametric Methods, by J. Gao. Journal of the Royal Statistical Society.

Series A (Statistics in Society), 172(4): 940–941, October 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622564.

Aalabaf-Sabaghi:2009:BRB

[864] Morteza Aalabaf-Sabaghi. Book review: Credit Risk Management: Basic Concepts, Financial Risk Components, Rating Analysis, Models, Economic and Regulatory Capital, by T. van Gestel; B. Baesens. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):941, October 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/20622565.

Hewson:2009:BRB

[865] Paul Hewson. Book review: Applied Multiway Data Analysis, by P. M. Kroonenberg. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):941–942, October 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/20622566.

Haigh:2009:BRB

[866] John Haigh. Book review: Polya Urn Models, by H. M. Mahmoud. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):942, October 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
20622567.

Jain:2009:BRB

[867] Varinder Jain. Book review: Computational Statistics Handbook with MAT-LAB, by W. L. Martinez; A. R. Martinez. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):942–943, October 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622568.

Corbyn:2009:BRB

[868] Judith Corbyn. Book review: Estimating Market Power and Strategies, by J. M. Perloff; L. S. Karp; A. Golan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4): 943–944, October 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622569.

Bedogni:2009:BRB

[869] Giorgio Bedogni. Book review: Clinical Prediction Models — a Practical Approach to Development, Validation and Updating, by E. W. Steyerberg. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (4):944, October 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/20622570.

Sprent:2009:BRB

[870] P. Sprent. Book review: Introduction to Nonparametric Estimation, by A. B. Tsybakov. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):944–945, October 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/20622571.

Best:2009:BRB

[871] Mark A. Best. Book review: Statistical Development of Quality in Medicine, by P. Winkel; N. F. Zhang. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4): 945–946, October 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/20622572.

Anonymous:2009:IAV

[872] Anonymous. Index of authors: volume 172, 2009. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):947–951, October 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:2009:BMd

[873] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):??, October 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        20622573.

Anonymous:2009:FMd

[874] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(4):??, October 2009. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        20622545.

Anonymous:2009:VI

[875] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (4):??, October 2009. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/20622544.